A Complete Bibliography of Publications in Computers in Physics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

11 April 2024 Version 1.03

Title word cross-reference

2 [332]. $20 [127]. 3 [136, 155, 205, 231, 332, 411, 428, 486, 559, 836, 938, 976]. $30 [227]. 32 [114]. 3d^{N} [923]. $45 [357]. $54.95 [1173]. $59.95 [656]. $69.95 [656]. 9 [1523]. _{2} [921, 1086]. _{3} [877]. F (t) · r [1106]. g_{3/2} [1312]. g_{5/2} [1312]. j [1523]. N [343, 456, 1311]. n +^{12} C [314]. U(1) [156]. W_{k,μ}(z) [1021]. X [700].

-Bit [114]. -Body [343, 456, 1311]. -D [136, 155, 205, 231, 332, 411, 428, 486, 559, 836].

0-12-174590-2 [656]. 0-387-94274-2 [1173]. 0-86720-225-4 [656].

1 [118, 119]. 1-2-3 [118, 119]. 10 [51]. 101 [367]. 11 [1409]. 128-processor [1469]. 14 [422]. 15V [115]. 1987A [269]. 1988 [1417, 1447, 1479, 1505]. 1990 [482]. 1990s [457]. 1991 [474]. 1MB [113].

2 [33, 36, 268, 840, 1227, 1240]. 2.1 [1200]. 2.X [840]. 205 [259]. 2a [268]. 2M [62]. 2nd [765, 1173].

3 [118, 119, 177, 1489]. 3.0 [1263, 1287]. 3.0.1 [1358]. 320pp [357]. 3rd [37].

4.0 [1127, 1155]. 4.5 [1381]. 45 [36]. 488 [197, 473, 962, 990]. 488.1 [864]. 4Sight [106]. 4th [762]. 4W [525]. 4W-mixing [525].

5 [115, 848, 870]. 5.1 [1466]. 5th [721].

68030-Based [224].

7in [112].

86MM [63]. ’88 [20, 91]. 88000 [149]. ’89 [226]. 8x [240].

90 [1118, 1119, 1121, 1122, 1307, 1406]. ’91 [422, 466, 467, 497]. ’92 [588]. ’93 [636, 704, 779, 809]. ’94 [780, 859, 880, 883, 930, 954]. ’95 [958, 1016, 1050, 1102, 1334, 1404, 1449, 1461]. 95-based [1334]. ’96 [1163, 1188, 1192, 1262]. ’97 [1294, 1296, 1348, 1380].

A/D [35]. AAPT [635]. Abelian [459]. absorption [411, 1086, 1220]. Abstraction [1216]. Academic [17, 656]. Accelerator [541, 613]. Accents [635]. Access [78, 611, 907]. Accessing [551]. Accidental [1501]. Accompany [756]. accuracy [1290, 1409]. Accurate [28, 223, 503, 1177, 1312, 1387, 1493]. Achieve [226]. ACM97 [1272]. acoustic [1313]. Acoustics [974, 975]. Acquires [39, 196]. Acquisition [75, 128, 256, 292, 339, 563, 790, 799, 802, 1056, 1334, 1424, 1433]. action [1046]. Active [430, 759]. Active-Walker [759]. ActiveX [1298, 1462, 1466]. Activity [576]. Acton [1236]. Actual [1255]. ADA [409, 1404, 1448, 1449]. Adam [969, 1079]. Adams [1118]. adapted [1493]. Adapter [817]. Adaptive [236, 565, 890]. Add [113, 1460]. Add-On [113]. add-ons [1460]. Adding [252, 931]. Addison [474]. Addison-Wesley [474]. Additional [81, 782]. Additions [1191]. Addresses [1006]. Adds [36, 59, 169, 198, 1381]. Adept [24]. Adhikari [1457]. Advance [888, 931]. Advanced [202, 892, 926, 1034, 1503]. Advanced-Placement [1034]. Advances [535, 537, 554]. Advantage [952, 987]. Advantages [695]. Adventure [1318]. Adventures [270, 521]. Advertisement [360]. Advising [1249]. Affiliates [17]. After [841, 872]. Again [1027]. Age [4]. Aggregation [278, 327]. Aging [1172]. Ago [1417, 1447, 1479, 1505]. Aharonov [925]. Ahead [366, 1418]. Aid [35, 133]. Aided [25, 148, 276, 338]. Aimed [268]. Aims [170]. AIP [814, 863, 889, 1500]. Al [1079]. Alamos [246]. Alan [971]. Albuquerque [537, 704, 722]. Alejandro [973]. Aleksander [357]. Alexander [994]. Alexei [994]. Algebra [213, 263, 433, 819, 871, 992, 1440].

Algebraic [371, 963, 991]. Algorithm [188, 211, 231, 334, 350, 389, 392, 444, 526, 641, 680, 726, 776, 797, 877, 925, 949, 965, 982, 1266, 1292, 1474, 1527]. Algorithms [140, 258, 295, 346, 349, 408, 503, 597, 627, 688, 787, 858, 937, 963, 991, 1044, 1096, 1105, 1115, 1143, 1195, 1352, 1428, 1493]. Alhambra [1504]. Allan [975]. Alliant [89]. Allocation [1017]. alloys [460]. Along [1364]. Also [1287]. Alternates [142]. Alternating [914]. Alternative [311, 948, 1060, 1144, 1240, 1267]. Always [477, 884]. America [752]. Ametek [51]. Ami [740]. Amiga [155, 344]. Amigas [400]. Amorphous [426, 460, 1502]. Amorphous-Computing [1502]. Amusia [1403]. Analog [735]. Analog-To-Digital [735]. analyses [1460]. Analysis [103, 189, 191, 192, 235, 276, 301, 369, 370, 429, 436, 458, 464, 527, 594, 646, 656, 690, 723, 736, 744, 747, 849, 923, 934, 981, 1114, 1129, 1223, 1424, 1431, 1464, 1465, 1521]. Analytic [732]. analytically [901]. analyze [1384]. Analyzer [169, 926, 1244]. analyzing [391]. Andrew [1146]. Angular [688, 818]. Angular-Momentum [688]. Animated [582, 646, 1369]. Animating [344]. Animation [728, 904, 1357, 1384]. Animator [929]. anisotropic [1386]. Anne [767]. Annealing [253, 490, 582, 679, 855, 1024]. Anniversary [1225, 1238, 1239]. Announce [636]. Announces [532]. Annual [413, 421, 532, 652, 791, 940, 1077, 1213, 1350, 1371]. ANOVA [194]. Ansatz [1235, 1514]. Answer [450]. antialiased [332]. Anticipated [1255]. antiferromagnet [347]. antiprotonic [184]. Antonelli [1146]. Aperiodic [509, 651]. apertures [1222]. Apollo [67]. apparatus [313]. Appeal [1234]. Appear [660]. Apple [98]. Appliances [385]. Application [136, 260, 293, 296, 486, 656, 666, 801, 924, 925, 1426]. Applications [1, 35, 44, 231, 368, 424, 592, 603, 656, 690, 713, 714, 754, 823, 912, 916, 1038, 1050, 1100, 1130, 1173, 1256, 1296, 1317, 1375, 1385, 1491]. Applied [419, 771, 822, 889]. Approach [139, 290, 348, 373, 598, 599, 702, 965, 1039, 1197, 1362, 1388, 1485, 1522, 1525]. Approaches [207, 1253]. Approaching [780]. appropriate [778]. approximant [1292]. Approximants [548, 1021]. Approximate [743, 1472]. approximation [724, 1474]. April [763]. APS [1416]. Aptec [198]. aqueous [1386]. arbitrary [1222]. Architecture [69, 368]. Architectures [140, 206, 357, 686, 1142, 1232]. Ardent [104]. Argonne [150]. arising [260]. arm [99]. Arrangements [538]. Array [125, 407, 551]. arrays [187, 1265]. Arrive [55]. Arsenide [177]. Art [128, 604, 605, 765, 1196, 1347]. Article [360, 574, 1422, 1452]. Articles [863, 1343]. Artificial [13, 199, 235, 723, 829, 1124, 1282]. Arts [1324]. ASCI [1250]. ASECA [1386]. Asks [858]. aspects [331]. Aspray [1281]. Assault [1416]. assessment [831]. Assistance [697]. Assistant [356]. assisted [679, 831]. Assumptions [499]. Asteroids [1393]. Astronomer [1165]. Astronomers [1394]. Astronomical [457]. Astronomy [310, 338]. Astrophysical [454, 455]. Astrophysics [181, 1040, 1140, 1175, 1273]. asymptotic [1203]. AT&T [126]. Ate [437]. Athletes [1136]. Atlas [1427]. Atmospheric [1054]. ATOM [563, 1183, 1403]. Atomic [205, 472, 743, 1104, 1265, 1403, 1455, 1494].

Atomic-Level [205]. atoms [184, 296]. attraction [646]. Attractions [935, 1448]. attractor [1289]. Aug [1447]. Auger [313]. Author [888]. Authoring [886]. Authors [589]. Automata [446, 474, 524, 601, 920, 970, 1386, 1516]. Automated [755, 1039]. Automates [75]. Automatic [244, 1223]. averager [1437]. Avoid [485, 499]. AVS [1217]. Award [805]. Awarding [88]. Awards [286, 777, 1204]. Away [1126]. Awk [1405]. Axial [154].

B [895, 974]. Back [15, 228, 1478]. Background [568]. Bahcall [457]. Balancing [1135]. band [297, 1131]. Bang [631]. Bar [1456]. Bar-Yam [1456]. Barbara [1140]. bargain [709]. Barriers [1348]. Bartlett [656]. Based [224, 233, 256, 291, 391, 410, 545, 596, 628, 643, 692, 715, 773, 853, 855, 926, 971, 998, 1034, 1307, 1334, 1337, 1396, 1421]. Basic [344, 367, 603, 764, 773, 1375]. basins [646]. Basis [583]. Battle [120]. Baumann [1353]. BBEdit [1381]. Be [422, 906, 1004, 1051, 1249, 1270]. beam [902, 1494]. beams [445]. Beating [1216]. Beauty [3]. Become [587]. Becoming [955]. Beginner [765, 766, 990, 1489]. Beginners [263]. Beguelin [1079]. Begun [522]. Behavior [98, 642]. behaviors [648]. Being [1284]. Benefits [634, 1114]. Benjamin [995]. Berend [1306]. Berners [869]. Berners-Lee [869]. Bessel [28, 390, 463, 475, 1524]. Best [1327]. Beth [656]. Bethe [1235, 1514]. Better [203]. between [257, 313, 715]. Bevington [736]. Beylkin [656]. Beyond [1107, 1354]. bibliographies [1398]. Biconjugate [600]. Big [16, 17, 249, 631]. Bigelow [1145]. Bill [1376]. bin [614]. Binder [691, 1260]. binding [982]. Biological [1172–1174, 1214]. biologically [430]. Biology [916, 968]. Biophysics [577]. bistability [60]. Bit [114, 215, 312, 1193]. bit-map [312]. Black [1074]. blessing [1096]. Blind [247]. Block [231, 312]. block-tridiagonal [231]. Blocks [125]. Blue [927]. Board [32–34, 62, 66, 197, 1437, 1467]. Boards [67]. Boasts [114]. Body [151, 343, 456, 716, 1311]. Bohm [925]. boil [852]. Boltzmann [788, 946]. bombarded [272]. bombardment [295]. Book [43, 45, 118, 119, 302, 357, 605, 628, 654–656, 670–674, 690, 691, 713, 719, 736, 760– 767, 784–786, 821, 822, 844, 845, 868, 869, 894, 916, 942, 969– 975, 995, 1015, 1036, 1037, 1079, 1099, 1100, 1118–1124, 1145, 1146, 1148, 1150, 1152, 1173, 1196, 1197, 1236, 1260, 1283, 1284, 1303, 1305, 1307, 1326– 1328, 1353, 1354, 1375–1378, 1402, 1403, 1426, 1456, 1457]. Book [216, 474, 627, 629, 737, 895, 915, 994, 1147, 1149, 1151, 1174, 1215, 1237, 1259, 1281, 1282, 1306, 1352, 1425, 1427, 1488, 1489, 1515, 1516]. Bookshelf [1397]. Boost [225, 794, 1444]. Boosting [92]. Boosts [31, 32, 204, 1187]. Born [1027]. Born-Again [1027]. BOS [68]. Bose [1312]. Boston [656]. Both [581, 683]. Bottlenecks [639]. Boundary [508, 530, 546, 1157, 1202, 1256, 1292]. Boundary-element [1256]. boundary-value [1157]. Brace [656]. brachistochrone [373]. Brain [180, 357]. Brain-Like [357]. Brainerd [1118]. Brand [45]. Brandt [1147]. Bravo [1449]. Breadth [1069]. breakdown [1495]. Breaks [227]. breakup [314]. Brendan [765]. Brian [1146]. Bright [834]. Bring [49, 520, 1068, 1247]. Bringing [661]. Brings [34, 37, 247, 537, 554, 705, 751, 1034, 1273, 1380]. Broad [907]. Broaden [1163]. Brown [382, 971]. Browser [861]. Bruce [1036, 1150]. Bubble [852, 1463]. bubble-dynamics [1463]. Build [201]. Building [125, 492, 1138, 1186, 1420]. Bundle [638]. bundles [291]. Burkitt [627]. Bus [215, 473, 686, 817, 864, 962, 990, 1056, 1322, 1424]. Business [953]. Buy [29, 61, 103, 141, 164, 190, 217]. Buyer [998].

C [218, 219, 330, 612, 766, 845, 942, 943, 972, 1060, 1118, 1125, 1196, 1198, 1215, 1216, 1240, 1241, 1303, 1328, 1330, 1352, 1426, 1428]. C. [714, 1122, 1528]. CA [656]. CAD [42]. calculate [159]. Calculating [184, 488, 1001, 1132]. Calculation [390, 444, 447, 1076, 1159, 1354, 1523]. Calculations [187, 536, 579, 726, 743, 1046, 1236]. Calculator [73, 220]. calculators [1517]. Calculus [228, 433, 878, 1023, 1320, 1439]. Calendar [540, 557, 575, 610, 637, 664, 685, 708, 731, 783, 810, 837, 860, 885, 909, 960, 1230]. Calender [753]. Call [1238]. CAMAC [563, 1322]. Camera [23]. cameras [37]. Campbell [1281]. Campbell-Kelly [1281]. Can [52, 442, 469, 710, 858, 880, 997, 1093]. Can-Sized [469]. Capabilities [1200]. Capacitance [511]. capacity [31]. Captured [48]. Captures [517]. Car [726]. Care [729]. Careful [819, 1004]. Carlo [261, 301, 314, 352, 395, 434, 464, 564, 566, 644, 650, 691, 798, 980, 1002, 1064, 1172, 1221, 1260, 1351, 1363]. Carnegie [50]. Carnegie-Mellon [50]. Carol [763]. Case [235, 441, 1315]. cash [1067]. Catalog [761]. catalysts [618]. catalytic [618]. Catastrophes [491]. Catching [686, 817, 864, 962, 990, 1056, 1322, 1424]. Cautionary [1161]. cavity [273]. CCP [1445]. CD [48, 1173, 1436]. CD-ROM [1173, 1436]. CEBAF [128, 249]. Celebrating [1225]. Celerity [30]. cell [233, 727, 1042, 1158]. cells [725, 900]. Cellular [446, 524, 601, 920, 970, 1386, 1516]. cellular-automata [1386]. Center [16, 17, 146, 150, 174, 177, 202, 227, 245, 422, 558, 957, 1254, 1277, 1510]. Centers [127, 834, 1090]. cerebral [186]. CFD [747, 1190]. Chain [531, 1525]. Challenge [364, 879, 1026, 1269]. Challenges [401, 684, 705, 1283]. Champaign [865]. Champions [907]. Change [396, 522, 705]. Chaos [318, 354, 391, 402–404, 414, 418, 419, 421, 438, 462, 520, 549, 604, 605, 703, 774, 820, 882, 916, 942, 1098, 1113, 1266, 1521]. Chaotic [102, 315, 386, 391, 417, 506, 621, 737, 822, 1013, 1059, 1480]. chaotic-pendulum [1480]. characterization [233, 900, 1042]. Characterize [651]. characterized [619]. Charge [964]. charged [428, 925]. charged-particle [428]. charges [329]. Charles [628, 654, 656, 670, 895]. charm [595]. charting [1462]. Charts [105]. Cheap [551]. Chebyshev [333, 841]. Chebyshev/radiation [333]. Checks [280, 1058]. Chemical [623, 1084, 1114]. Chemistry [483, 500, 916, 1071, 1295]. Chen [714]. Chernysheva [1403]. Children [952]. Chip [20, 176]. Chivers [1119]. Choice [58, 968]. Chooses [929]. Chores [1241]. Chuck [1376]. Chui [654, 656]. Cinema [629]. CIP [413, 532, 652, 789, 791, 813, 911, 1009, 1032, 1052, 1053, 1077, 1139, 1167, 1191, 1208, 1213, 1225, 1228, 1238, 1239, 1251, 1274, 1297, 1304, 1342, 1350, 1368, 1371, 1395, 1397, 1419, 1422, 1450, 1452, 1481, 1483, 1507, 1509]. CIP’S [940]. circle [278]. circuit [410]. Circuits [276, 657, 773]. CiSE [1509]. Clarifies [964]. Clarifying [1137]. class [293]. Classes [638]. Classical [139, 531, 617, 1150, 1180, 1353]. classically [621]. Classics [629, 1238, 1304]. Classroom [155, 234, 438, 440, 739, 953]. Classrooms [49]. Clear [1366]. Clebsch [529, 877]. Climate [21, 364]. Clocks [1250]. Clones [1082]. Close [53]. closed [296, 615]. Closer [1367]. cloth [1173]. Cluster [683, 695]. Clusters [488, 663, 682]. CM [268, 848, 870]. CM-2a [268]. CM-5 [848, 870]. CO [1086]. coarse [949]. coated [797]. cobalt [545]. cobalt-based [545]. Code [100, 136, 184, 612, 644, 775, 789, 813, 1001, 1038, 1086, 1155, 1311, 1408, 1463]. Codes [150, 345, 547, 580, 1042, 1158]. Codevelop [17, 18]. coefficient [1022]. Coefficients [529, 688, 797, 877]. Coercivity [460, 826]. coexistence [1280]. coherently [296]. Coifman [656]. coils [647]. Cold [496]. Collaborate [1413]. Collaboratory [1186, 1420]. collecting [1051]. Collection [1017, 1398]. Collective [57, 1445]. College [1207, 1324]. collisionless [1311]. collisions [314, 563]. Colloidal [914]. Color [59, 65, 164, 178, 711, 1177]. Colorado [174]. Coloring [406]. Column [983]. Combine [807]. Come [908]. Comes [857]. Coming [889]. Command [591]. Comment [1414]. Comments [429, 1528]. commercial [214, 1265]. Commercializing [808]. Committee [457]. Communication [794, 887]. Communications [425]. Communities [809]. community [684]. compact [255]. Companies [750]. Companion [766]. comparative [723]. Comparison [346, 602, 624, 697, 772, 1458]. Compatible [74, 397]. Compatibles [76]. compensation [312]. Compete [123]. Competition [1187]. Compiler [38, 219, 1241, 1379, 1461]. Compilers [795]. Compiles [985]. Complete [1215, 1483]. Complex [237, 297, 474, 523, 732, 1037, 1173, 1174, 1199, 1256, 1280, 1456, 1516, 1519]. complex-plane [237]. Complexity [544, 576, 970, 1037]. components [1451]. Comprehensive [568, 583]. compression [312]. Compromise [772]. Comptonization [1363]. Comput [1409]. Computation [129, 140, 260, 289, 476, 577, 634, 746, 797, 807, 875, 877, 891, 899, 912, 915, 968, 1021, 1203, 1347, 1403, 1510]. Computational [49, 60, 137, 182, 199, 293, 320, 325, 353, 374, 396, 454, 498, 541, 543, 558, 627, 632, 639, 689, 705, 721, 750, 760, 777, 786, 824, 843, 847, 859, 896, 972, 974, 1005, 1010, 1027, 1099, 1141, 1161, 1175, 1243, 1268, 1294, 1305, 1390, 1416, 1422, 1425, 1442, 1452, 1500, 1504, 1522]. computational-physics [1422]. computations [235, 1022]. Compute [52, 1394]. Computer [4, 5, 11, 23, 25, 44, 53, 55, 61, 91, 92, 139, 148, 154, 159, 195, 206, 213, 225, 257, 262, 263, 269, 272, 273, 276, 285, 302, 313– 315, 338, 342, 360, 365, 378, 397, 424, 444, 455, 501, 522, 525, 562, 576, 587, 598, 628, 640, 642, 644, 660, 679, 726, 733, 819, 827, 831, 866, 871, 900, 946, 956, 964, 976, 992, 1006, 1034, 1048, 1086, 1133, 1158, 1173, 1174, 1180, 1197, 1207, 1244, 1246, 1248, 1250, 1281, 1324, 1325, 1360, 1384, 1398, 1400, 1440, 1496]. Computer-Aided [25, 148, 276, 338]. Computer-Algebra [992, 1440]. Computer-assisted [679, 831]. Computer-Based [1034]. Computer-Compatible [397]. Computer-Controlled [61]. Computer-enhanced [1400]. Computer-Generated [53, 55, 91, 598]. Computer-Rich [1324]. Computer-Supported [1207]. Computer-Video [733]. Computers [15, 22, 96, 120, 132, 133, 145, 151– 154, 230, 249, 286, 349, 380, 408, 425, 439, 442, 470, 487, 496, 514, 523, 544, 553, 571, 608, 635, 666, 738, 749, 778, 795, 804, 986, 1044, 1068, 1110, 1136, 1188, 1279, 1302, 1309, 1329, 1339, 1355, 1458, 1467, 1469, 1475, 1500, 1512]. Computin [1347]. Computing [21, 54, 123, 160, 179, 207, 209, 223, 238, 239, 271, 282, 290, 350, 357, 364, 371, 384, 422, 428, 457, 466, 467, 472, 497, 535, 558, 588, 620, 636, 690, 694, 695, 751, 758, 779, 781, 806, 857, 859, 880, 917, 943, 946, 954, 957, 967, 987, 1008, 1016, 1028, 1033, 1079, 1082, 1084, 1109, 1171, 1192, 1198, 1236, 1283, 1330, 1341, 1346, 1354, 1367, 1390, 1415, 1427, 1441, 1444, 1502]. conceptions [1267]. Concepts [367, 603, 710, 891]. Concrete [302]. Concurrent [17, 18, 216, 528, 799]. Condensed [502, 1260, 1416]. Conditions [163]. conductivity [350, 1180]. conductor [1498]. Conductors [964]. Conference [200, 362, 467, 482, 721, 806, 1478, 1503]. Configuration [329, 1015]. Confinement [132, 133]. conformational [1195]. conformations [1525]. Confronting [1267]. conjugate [293]. Connection [325, 334, 372, 471, 641, 645, 947]. Connoisseur [1327]. conservation [1410]. Consider [905]. Considers [754]. Consortium [550]. Conspiracy [758]. constants [159]. constrained [388]. constructing [1440]. Construction [171, 251]. Constructive [257]. Constructs [483]. Container [1428]. Container-free [1428]. containing [410, 1474]. Contest [413, 532, 652, 791, 940, 1077, 1204, 1210, 1213, 1253, 1350, 1371]. contestants [1067]. Continuation [732]. Continued [160, 390]. Continuous [332, 404, 490]. continuum [258, 350, 472, 562, 1522]. Contour [264, 510]. contribution [184]. Control [75, 249, 276, 563, 735, 802, 971, 1335, 1412, 1462, 1521]. Controlled [61, 111, 430, 623]. Controller [37]. Controlling [820, 1266, 1278]. Controls [78]. convection [1000]. convenes [704]. Convenient [758]. Convention [422]. convergent [431]. Conversion [735]. converter [254]. converters [35]. Converting [1318]. Convex [123]. Convolution [407, 1331]. Cooney [670]. Coordinates [581]. Coprocessor [32]. Core [1477]. Cornell [227, 1418]. Corner [1092]. Corners [667]. coronal [1220]. Corrected [82, 1260]. correlated [702]. Correlations [893, 1132, 1472]. Cosine [649]. Cosmic [11, 156, 882]. Cosmological [456]. Cost [313, 606, 608, 695, 711]. Cost-Effective [608]. Could [1133]. Counter [608]. Counts [928]. Coupled [214, 696, 775]. Coupling [688]. Course [737, 847, 892, 1025, 1422, 1425, 1452]. Courses [501, 514, 550, 1040, 1246]. Courseware [449, 676, 978, 1097]. Cover [573, 782]. Coverage [1449]. CPUs [595]. Crack [1416]. Cracking [1143, 1170]. Crafts [809]. Craig [764]. Crandall [915]. Crash [1091]. Crash-on [1091]. Cray [86, 123, 136, 177, 208, 244, 483, 884, 1311]. Create [559, 977, 1345]. Creates [518, 687]. Creating [996, 1275, 1369, 1482]. Credit [539]. credits [782]. Critical [279, 396, 801]. Criticality [491, 993]. CrossCode [218]. Crossing [767]. Crown [123]. CRUNCHER [100]. Crystal [923, 1366]. Crystallography [96, 804, 1164]. Crystals [159, 308, 1487]. Ct [687]. CTSS [136]. cubic [159, 615]. Cues [868]. Cultural [1026]. CUPS [756, 886, 1040]. current [1458]. curricula [1421]. Curriculum [353, 374, 1279, 1302, 1324]. Curve [185, 701, 948]. Curves [167, 444, 615, 1024]. custom [1462]. Cutting [667]. CVODE [1125]. CYBER [259]. CYBER-205 [259]. Cycles [75]. Cyclic [280]. cylindrical [876].

D [35, 119, 136, 155, 205, 216, 231, 332, 411, 428, 486, 559, 605, 691, 736, 836, 938, 976, 1151, 1375]. D. [627, 1196, 1303]. Daan [1306]. Daedalon [512]. damaging [275]. Damping [1245]. Daniel [1516]. DAQ [609, 1437]. DARPA [201]. Data [10, 39, 74, 75, 81–83, 128, 138, 166, 185, 188, 196, 204, 212, 215, 236, 252, 253, 256, 274, 280, 292, 298, 301, 307, 324, 337, 339, 389, 412, 432, 436, 477, 486, 507, 527, 563, 569, 581, 651, 653, 717, 733, 736, 790, 793, 794, 799, 802, 868, 890, 939, 1042, 1061, 1278, 1286, 1334, 1424, 1433, 1434, 1446, 1462, 1465, 1471, 1474]. Data-Acquisition [790, 1334, 1424, 1433]. Data-Adaptive [890]. Data-Networking [794]. data-processing [1434]. data-smoothing [236]. Data-Time [1286]. data-visualization [1471]. Database [173, 800, 1275]. Databases [22]. Date [814]. Daubechies [655, 656, 1002, 1085]. David [426, 895, 1237, 1283, 1377]. Dawson [241]. Day [307, 752]. Deadline [780, 958, 1253]. Dealing [1007]. Debut [173]. Debye [159, 488, 668]. Dec [1505]. decay [11, 1183]. Deciphers [631]. Decisions [485]. Decoder [556]. Decomposition [982, 1002, 1233]. deconvolution [1410]. dedicated [1360]. Defense [359]. Definitive [1376, 1377]. Delights [1287]. Delivers [1127, 1218]. demand [1390]. Demons [375]. Demonstrates [98]. Demos [1345]. Demoshield [1345]. Denis [673]. Denning [1354]. Dense [1331, 1525]. densities [1129]. density [801, 1159]. Department [441]. departmental [30]. Departments [1224]. dependent [210, 526, 925]. deposition [260]. Derivatives [447]. derive [314, 356]. derived [527]. derived-data [527]. DERIVE(R) [433]. describing [1132]. description [855]. Design [4, 125, 276, 341, 357, 587, 647, 710, 850, 1273, 1335]. designed [828]. Designers [609]. Designing [1298]. Designs [1142, 1232]. Desktop [53, 272, 644, 1309, 1441]. Detection [269]. detectors [1333, 1337]. Determination [529, 572]. Determining [211, 772, 1131]. Determinism [604, 605]. Deux [106]. Devaney [737]. Develop [1097]. Developer [805, 858]. developers [777]. Developing [205, 1324, 1421]. Development [71, 928, 1337]. Developments [248, 1054, 1380]. Develops [550, 558]. deviations [901]. Devices [310, 341, 481]. Devil [1007]. Diagnosing [1199]. diagnostic [430]. Diagonalization [297, 529, 734]. Diagram [1076]. Diego [656, 1271, 1277]. Dielectric [273, 445, 1495]. Diffeomorphism [405]. Difference [938, 949, 1003, 1313]. differences [1290, 1385, 1409]. Differential [214, 265, 299, 321, 336, 678, 681, 703, 775, 849, 963, 991, 1066, 1086, 1292, 1426, 1513]. differential-absorption [1086]. differential-algebraic [963, 991]. differentiation [1223]. Difficult [937]. Diffraction [488, 668, 1222]. diffractive [1182]. Diffusion [434, 623, 1065]. Diffusion-Controlled [623]. Digital [33, 77, 254, 735, 913, 1246, 1284]. digitizer [1437]. digitizer/averager [1437]. Digitizes [108]. dimension [37, 42, 186]. Dimensional [138, 187, 235, 309, 407, 432, 599, 742, 848, 870, 920, 947, 1042, 1181, 1222, 1313, 1408, 1459]. Dimensions [210, 298, 921, 1217, 1438]. Dimitri [869]. Dimitroyannis [869]. Ding [975]. dinosaurs [749]. Dirac [472]. Direct [231, 376, 1351, 1523]. directions [1512]. Director [907]. Directory [449, 676, 978, 1239]. Dirt [1206]. Disambiguated [1285]. Disasters [936]. discharges [727]. discipline [274]. discipline-independent [274]. Disciplines [1134]. Discontinuous [404]. Discrete [232, 332, 788, 1249, 1361]. Discuss [634]. Disease [1141]. disk [36]. Disneytown [1478]. Disparate [939]. display [310, 722]. displaying [1369]. dissipative [621]. Distance [102, 1454]. Distributed [715, 922, 1313, 1497]. distribution [183]. distributions [1105, 1472]. divergent [472, 1203]. Diverse [520]. Division [353, 374]. DNA [1444]. Do [308, 578, 746, 1013, 1160]. Do-it-Yourself [578]. Documents [1252]. DOE [146]. Does [33, 145, 191, 194, 578, 746, 1321]. Doesn’t [555]. Dog [437]. Doing [1205]. DOM4 [800]. Domain [369, 949, 1003, 1233, 1292]. Dome [1273]. Dominated [1199]. Don [243]. Donald [302]. Dongarra [1079]. Donnelly [673]. Don’t [477]. DORE [104]. dose [260]. Douglas [1147]. Down [279, 1273]. Download [1210]. Downloaded [1004]. Dräger [1488]. Draw [6]. Dripping [386]. Drive [148, 168]. driven [1258]. Driver [63]. DSP [35]. Duilds [176]. during [273]. Dynamic [604, 605, 613, 846, 1017, 1036, 1511]. Dynamical [57, 406, 419, 432, 505, 648, 737, 1519]. Dynamicists [1249]. Dynamics [51, 126, 139, 272, 295, 372, 417, 418, 454, 471, 474, 500, 502, 531, 562, 760, 785, 822, 916, 942, 981, 982, 1002, 1037, 1116, 1196, 1278, 1303, 1384, 1453, 1456, 1463, 1515, 1519]. dynamo [1046].

E-journal [1394]. Early [226, 453]. earth [366, 460, 939, 1091, 1273, 1300, 1477]. earth-transition [460]. Earthquake [934, 1374]. Eases [65]. Easy [27, 841, 1347]. echo [1399]. echo-reconstructive [1399]. Echoes [631]. Eclipse [112]. ed [761, 762, 765]. Eden [1226]. Edinburgh [590]. Editing [740]. Edition [1173, 1215, 1260, 1376, 1377]. editor [844, 1260, 1348, 1506]. Editorial [121, 1230]. editors [656]. Educate [1134]. Education [286, 461, 468, 522, 608, 986, 998, 1068, 1069, 1097, 1207, 1396, 1451]. Educational [554, 718, 748, 987, 1003, 1019, 1213, 1239, 1350, 1371, 1389, 1445].

Educators [634, 1392]. Educom [809]. Edward [1327, 1402]. Effect [365, 925]. Effective [555, 608]. Effects [642, 899, 1255]. Efficiency [136, 1444]. Efficient [183, 350, 1311, 1312, 1383]. Ehrlich [1148]. Eiffel [613]. eigenvalue [1143, 1170, 1372]. Eigenvalues [328, 757, 1131]. eigenvector [1266]. Eigenvectors [757]. Eighth [1350]. Einstein [1312]. Elastic [159, 562, 938]. Electric [385, 914, 1180, 1183, 1486]. Electric-Field [1486]. Electrical [576, 773]. Electricity [190, 1148]. Electrodynamics [397]. electromagnetic [184, 1405]. electromagnetics [1520]. Electron [99, 370, 644, 816]. Electron-Phonon [816]. Electronic [202, 276, 338, 399, 863, 887]. Electronics [657, 1246]. Electrons [99, 816]. electrostatic [428, 1157]. Element [328, 567, 1256]. Elementary [785]. Elements [9, 849, 941, 1182, 1202]. Elizabeth [764]. ellipse [329]. ellipsoidal [1498]. Elliptic [335, 1012]. ElPed [773]. Emerge [519]. Emphasizes [806]. Empirical [971]. Enables [1031]. End [20, 323, 585]. Energy [184, 311, 328, 542, 544, 553, 579, 726, 875, 1410]. Engeln [1352]. Engeln-Müllges [1352]. Engineer [821]. Engineering [168, 355, 671, 784, 786, 805, 916, 1100, 1130, 1135, 1236, 1257, 1390, 1500]. Engineers [118, 672, 674, 690, 822, 942, 1328]. enhanced [596, 1400, 1425]. Enhancement [64]. Enhances [913, 1422, 1452]. Enjoyable [161]. Enns [1328]. Enough [819, 943]. Enrich [1030]. Enters [4]. Entire [75]. Entrepreneur [1501]. Entropy [429, 744, 1258]. Entropy-driven [1258]. envelope [444]. Enviable [985]. Environment [38, 136, 384, 452, 568, 583, 718, 794, 922, 1054, 1075, 1357, 1406, 1420]. environment/application [136]. EOS [366]. epitaxial [980]. equals [833]. Equation [134, 388, 526, 740, 742, 903, 925, 1022, 1292, 1440, 1524]. Equations [214, 232, 265, 299, 321, 336, 337, 415, 606, 619, 678, 681, 703, 775, 849, 858, 876, 963, 991, 1066, 1290, 1409, 1426, 1513]. equilibria [1001]. equilibrium [329]. Equipment [409, 998]. Equipotential [921]. Era [43]. Eric [450]. Erratum [1409]. Error [464, 736, 963, 1412]. Errors [581, 678, 700, 735, 1236]. Espouses [1069]. Establishes [202]. Estimation [678, 890, 963]. ETA [51]. etching [1386]. ethane [618]. Euler [963]. European [1187]. EUV [563]. evaluate [1106]. Evaluates [733]. Evaluating [160]. Evaluation [408, 487, 1101, 1291, 1312, 1349]. Evans [147]. Everybody [969]. everyone [1459]. Evidence [1402]. Evolution [587, 699, 1183]. Evolving [426]. Exact [734, 949]. Exactly [193]. Examine [1317]. Example [157, 768, 1322]. Examples [1353]. Excel [672, 1460]. Excels [40, 1019]. excited [296, 1183]. Exercises [374]. Exhibit [554, 750]. Exhibition [883, 932]. Exhibitors [516, 722]. EXP [606]. expand [747]. Expanded [839]. expanding [261]. Expands [17, 42, 1500]. Expansion [66]. Expansions [867, 875, 1021]. Experience [292, 470, 558, 1330]. Experiment [180, 737, 807]. Experimental [137, 239, 666, 1270, 1420]. Experimentalists [1248]. Experimentation [710]. Experiments [7, 11, 249, 315, 544, 844]. Expert [24]. Experts [1108]. Explained [240, 1120]. Explanations [1402]. Explication [1184]. Explicit [526, 1524]. Exploits [695]. Explorations [1173, 1174]. Exploratory [412, 1081]. explorer [648]. Explores [859]. Exploring [732, 1035, 1237]. Exponential [160, 1105, 1526, 1527]. exponentially [1440]. Exponents [801]. Expression [1216, 1285]. Extended [551]. Extending [1176]. Extensions [1344]. exterior [878]. external [1256]. extracting [294]. extreme [563]. extreme-ultraviolet [563]. extremely [212]. extremes [723]. Eye [94, 326]. Eyesight [247].

F [1100, 1151]. F. [974]. Fabricate [20]. Face [936]. Facelift [1032]. Faces [1026]. Facets [54]. Facilitates [572]. Facilities [590]. Facility [86]. Facing [1269]. Fact [172]. Fall [1426]. Fare [1314]. farewell [1512]. Fast [314, 334, 389, 487, 501, 526, 680, 943, 1001, 1046, 1105, 1184, 1526]. FASTBUS [277]. Faster [1198, 1349]. Faucet [386]. fault [1374]. Favor [1205]. fcc [238]. FEA [1194]. Feagin [894]. Featured [1139]. Features [38, 69, 883, 1358]. Ferdi [1326]. Fermi [1515]. Fermilab [696]. ferromagnetic [211]. Fertile [1171]. Few [716]. Few-Body [716]. Fewer [556]. Feynman [1076, 1106, 1411]. Feynman-Diagram [1076]. FFT [223, 1193]. Fiber [201, 291]. Fiction [172]. fictitious [982]. Field [208, 296, 307, 392, 458, 543, 647, 689, 843, 923, 925, 1183, 1486]. field-line [392]. Fields [397, 428, 486, 711, 797, 914, 1018, 1331, 1405]. Fifth [940, 1022]. Fifth-order [1022]. Fifty [1354]. Files [715, 812, 996]. film [545, 642]. films [460, 1470]. filter [427]. Filtering [77]. Filters [435]. Finalists [779]. financial [1301]. Find [862, 1057, 1093, 1418]. Finding [776, 1028, 1195]. Fine [1314]. Finite [328, 567, 680, 849, 938, 941, 949, 1003, 1290, 1313, 1385, 1387, 1409]. Finite-Difference [938, 949, 1003, 1313]. Finite-Element [328, 567]. finite-interval [1387]. Fink [821]. First [413, 737, 748, 847, 887, 1234, 1319, 1406, 1414]. first-order [1414]. first-time [1406]. First-Year [1234]. Fischer [1146]. Fit [135, 772, 1040]. Fits [143, 167, 700, 1132]. Fitted [464, 1440, 1524]. Fitting [185, 311, 337, 569, 581, 615, 701, 712, 948, 1024]. Flagship [150]. Flanagan [1237, 1377]. Flexibly [849]. flight [802]. Floating [477, 1458]. floating-point [1458]. flow [562, 747, 1432, 1497, 1498]. Flows [179, 505, 758, 788, 1199]. Fluid [51, 179, 454, 500, 727, 760, 788, 1249, 1432, 1503, 1519]. fluid-dynamical [1519]. Fluid-flow [1432]. fluid-simulation [1503]. Fluids [1485]. fluxes [275]. Focus [2, 520, 781, 1389, 1392]. focused [1494]. focuses [748]. Focusing [682]. Folkers [1375]. following [563]. Force [1267]. Forecasting [604, 605, 1054]. Forecasts [1272]. forerunner [706]. Foresees [834]. Forest [836, 956]. Forever [1160]. Form [638, 759, 957, 1106, 1276]. formalism [371]. Forman [1236]. Formation [807, 882]. Formats [606]. forming [1384]. Forms [801, 1210, 1229]. formula [1523]. formulation [232]. FORTH [58, 289, 482]. Fortran [240, 345, 387, 620, 675, 693, 738, 775, 795, 805, 895, 977, 1118– 1123, 1155, 1205, 1307, 1309, 1352, 1406, 1408, 1461]. Forward [640]. Foster [605]. Foundation [225, 302]. Four [77, 1473]. Four-tap [1473]. Fourier [189, 212, 270, 298, 334, 407, 546, 866, 1100, 1144, 1184, 1193]. Fourth [421, 791]. Fox [216]. Fractal [186, 303, 327, 506, 822, 893]. Fractals [3, 354, 893, 1353]. fraction [390]. Fractional [463, 475]. Fractions [160]. Frame [65]. Framework [694, 1490]. Francis [822]. Frank [1352]. Franklin [1282]. Franz [1099]. Fredholm [415]. Free [9, 910, 1060, 1428, 1463]. Free-surface [1463]. Frenkel [1306]. frequency [727]. Fresnel [649, 1222]. Friendly [400, 1001]. Frugal [1082]. Fuel [500]. Full [952, 1031]. Full-Text [1031]. fully [1525]. Fun [1327]. Function [333, 406, 566, 787, 890, 1159, 1203]. Functionality [1127]. Functions [28, 41, 185, 258, 376, 390, 393, 463, 472, 475, 667, 732, 992, 1012, 1021, 1043, 1312, 1427]. Fund [383]. Fundamental [891]. Fundamentals [1124]. Funding [127, 225]. Fusion [132, 133, 341, 496, 1420]. Future [45, 660, 834, 881, 1247, 1261, 1283, 1293, 1321, 1338, 1346, 1478, 1499]. FY1989 [127].

G [216, 671, 767, 784, 1307, 1375]. G. [216, 1146, 1197]. GaAlAs [203]. GaAs [203]. GaAs/GaAlAs [203]. Gadfly [1027]. Gaffin [969]. Gallium [177]. Gambol [1110]. game [1496]. gamma [256, 1337, 1464]. gamma-ray [256, 1337, 1464]. Garbage [1017, 1051]. Garcia [973]. Garfinkel [713]. gas [261, 1159, 1516]. Gases [524]. Gash [786]. gasket [616]. Gauge [182, 345, 771, 843]. Gauss [319]. Gaussian [259, 393]. Gaylord [1173, 1174]. Gear [1262]. Geert [821]. Gehrke [1123]. Geist [1079]. General [7, 216, 392, 451, 473, 666, 1426]. General-purpose [7]. generality [451]. Generalized [1385]. Generate [205, 842]. Generated [53, 55, 91, 331, 598]. generating [183]. Generation [28, 89, 1182, 1227, 1315, 1329, 1355]. Generator [643, 1057]. Generators [259, 622, 830, 1473, 1526, 1528]. Generic [405]. genesis [1515]. Genetic [1114]. geodesic [619]. geometries [619]. Geometry [972, 1046]. Geophysicists [91]. Geophysics [148]. George [1328]. Gerard [474]. Gerd [1353]. Get [177, 477, 539, 789, 813, 933]. Gets [195, 1032, 1198, 1207, 1319]. Getting [763, 1172, 1295, 1336]. GeV [99]. Giant [1049]. GIFs [1369]. Gifted [1034]. Gigabit [383, 781]. Gigaflop [231]. Gigaflops [226]. Giles [1146, 1391]. Giordano [1305]. Gisela [1352]. Give [538]. Gives [752, 872]. Global [646, 678, 766, 832, 835, 850]. Glommable [1285]. glow [727]. Glueball [1109]. Glut [412, 1061, 1278]. Glynn [1489]. GNU [1316]. Gnuplot [977]. Go [960, 1126]. Goes [16, 17, 839]. Going [501]. Golay [435]. Gold [632]. Goldberg [1118]. Goldmine [1010, 1074, 1190, 1398]. Goldmines [935]. Good [745, 872, 936]. Goodbye [90]. Gordan [529]. Gordon [877]. Gould [1149, 1173]. Government [383, 933]. Grabber [65]. Gradient [293, 600, 647]. Graham [302, 1152]. Granada [1442]. Grand [364, 401, 1006]. grandly [684]. Granted [1126]. Graph [105, 373]. graphic [699]. Graphical [1130, 1344]. Graphics [6, 34, 76, 80, 104, 106, 119, 168, 222, 436, 501, 976]. Graphing [309, 1127, 1189, 1460, 1517]. Graphs [944]. gratings [1159]. gravitating [1453]. Gravitational [882, 1333]. Gravity [351]. Gray [1307, 1489]. Great [1091]. Greatest [1269]. Green [566, 732, 1159]. Greenhouse [365]. Gregory [656]. grid [949]. grids [1390]. Grosberg [994]. Ground [211, 227]. Group [1496]. Growing [1226]. Growth [303, 635, 759, 980, 1117]. GRPP [828]. GRTensor [924]. Guide [367, 712, 761, 762, 764– 766, 821, 969, 990, 998, 1015, 1079, 1100, 1118, 1123, 1327, 1376–1378, 1489]. Guidelines [573, 888]. Guides [609]. Gulmarg [256]. guns [370]. Guy [895].

H [605, 672, 714, 821, 895, 921, 974, 1118, 1124, 1197, 1328, 1488, 1516]. H. [844, 1375]. Haile [785]. halide [853]. Hall [1168]. Hamiltonian [414, 462, 549, 1098]. Hamiltonians [311]. hand [159]. hand-held [159]. Handbook [767, 845, 895, 1403]. Handles [79, 116, 863]. hands [728]. hands-on [728]. Hannon [1036]. Harcourt [656]. Hard [36]. hardcover [656]. Hardware [542, 799, 1015, 1244]. harmonic [1043]. Harrison [243]. Hartley [1033]. Harvey [1149, 1173]. Hassoun [1124]. Hawkins [1150]. haystack [1195]. HCP [538]. HDF [715]. HDTV [857]. Head [245]. Healthier [1133]. Heard [906]. Hearings [87]. Heats [1070]. Heaven [1294]. Heavy [1340]. Heermann [627, 691]. height [926, 1244]. Heisenberg [531]. Held [159, 422]. Help [1136, 1203]. Helped [1089]. Helps [657, 668, 814, 1071, 1165]. Her [1135]. Herbert [1215]. Here [833, 857, 908, 1093]. Hermitian [297]. Heterogeneous [694]. hierarchical [702]. High [34, 88, 114, 164, 364, 370, 542, 544, 553, 563, 579, 585, 614, 645, 751, 816, 874, 895, 946, 954, 1060, 1240, 1262, 1283, 1290, 1295, 1329, 1347, 1355, 1399, 1408, 1409, 1424]. High-Energy [542, 544, 553, 579]. High-Performance [585, 751, 954, 1060, 1240, 1329, 1355, 1399]. high-perveance [370]. High-resolution [563, 645, 946]. High-Speed [614, 1347]. High-Temperature [816]. Higher [991, 1493]. Higher-order [991, 1493]. Highlight [1296]. Highlights [482, 497, 1010, 1050]. Hilbert [732]. Hilborn [942]. Hill [134]. Hiller [1147, 1151]. HIP [579]. History [441, 755, 1281]. Holes [1074]. Holograms [53, 55, 205, 331, 598]. Holtje [1375]. Home [1032, 1052, 1053, 1167, 1191, 1300]. Homework [842]. Hone [1136]. honor [777]. Horbatsch [1259]. horizons [747]. hosts [1442]. House [87]. Housekeeping [936]. HPC [590]. HPCC [684]. HTML [1376, 1381]. Hubbard [641]. Hubble [250]. Human [229, 326, 576, 1272]. Hundreds [862]. Hunting [270]. Hv [1130]. Hybrid [1221]. Hydrocodes [561]. Hydrodynamics [848, 870, 1516]. hydrogen [472]. hydrogenic [875]. hydrogenolysis [618]. Hypatia [1398]. Hype [662]. HYPERCARD [317, 842]. Hypercube [226]. Hypergeometric [376]. Hyperimages [1343]. hypersphere [183].

I. [1119]. I/O [33, 196, 504, 1358]. Ian [785]. IBM [33, 36, 76, 126, 205, 504, 946]. ICN [833]. Icon [339]. Icon-Oriented [339]. ICs [556]. Ideas [481]. Identical [538]. Identity [280]. IDL [824, 897, 1466]. IEEE [197, 473, 864, 962, 990, 1500]. IEEE-488 [197, 473, 962, 990]. IEEE-488.1 [864]. If [319, 833, 1126]. Igor [357, 436]. II [34, 48, 98, 462, 530, 583, 624, 653, 719, 755, 769, 843, 870, 912, 991, 1011, 1170, 1232, 1302, 1355, 1514]. III [549, 594, 692, 864, 1075]. III-V [594]. IISS [874]. Illinois [1271]. Illuminate [707, 939]. Illustrate [524]. Illustrations [573]. Image [65, 170, 171, 310, 312, 845, 853, 874, 913, 1056, 1384, 1469]. Images [48, 155, 186, 506, 559, 584, 633, 1185, 1220, 1402]. imaginary [923]. Imagination [517]. Imagine [109]. Imaging [151, 152, 203, 431, 647, 1186, 1399]. IMD [1470]. Imitate [229]. Immunologically [601]. Impact [469, 1163]. Implement [1227]. Implementation [564, 641, 726, 1042, 1044, 1066]. Implementations [1170]. Implementing [529, 1311]. Implicit [912]. Impossible [1057, 1065]. Impresses [824, 944]. Improve [710, 1460]. Improved [688, 1200]. impulse [369]. includes [1466]. including [923]. Incomplete [995]. Increase [127]. incremental [38]. independent [274, 724]. Index [589, 659, 1023, 1483]. Indexer [63]. Indexer/Driver [63]. indices [28, 1439]. Indigo [585]. Individualized [842]. Industrial [985, 1256, 1294]. Industry [383, 933, 1048]. Ineffective [306]. Inertial [133]. Inevitable [936]. Inexpensive [137, 233, 817]. InfiniteReality [1107]. inflating [904]. Informal [469]. Information [706, 729, 764, 814, 989, 1029, 1281]. Informative [162]. Infrastructure [706, 729]. Ingrid [655, 656]. inhomogeneous [925, 1363]. Initial [163]. initio [726]. Innovations [859]. Input [1142, 1467]. Input/Output [1142]. Inputs [116]. Insight [633]. Install [86]. Installs [17, 19]. Institutes [148]. Instruction [913, 1034, 1042, 1158, 1166, 1369]. Instructional [547, 767, 1062]. Instructions [79]. Instrument [686, 755]. Instrumentation [61, 276, 1197, 1255, 1365, 1433]. Instruments [58, 78, 1337]. integrable [505]. Integral [241, 346, 415, 1106, 1331]. Integrals [160, 223, 335, 649, 1012, 1291, 1349, 1387]. Integrated [139, 583, 694]. Integrates [110, 192, 1491]. Integrating [265]. Integration [352, 376, 503, 565, 742, 876, 1388, 1493]. integrator [1412]. Integrity [280]. Intelligence [13, 235, 252]. Intelligent [385, 854, 1114, 1502]. interacting [1360]. Interaction [156, 257, 1081, 1106, 1408]. Interactions [816]. Interactive [274, 473, 479, 719, 773, 775, 874, 920, 1112, 1229, 1234, 1257, 1286, 1287, 1298, 1303, 1345, 1482]. Interactively [1009]. intercommutation [156]. Interdisciplinary [1317]. Interface [7, 33, 62, 197, 215, 313, 432, 473, 775, 817, 896, 977, 1130]. interfaces [445]. interfacing [11, 273]. interference [902]. interferometers [902]. internal [260]. International [721]. Internet [761–767, 789, 808, 811, 812, 838, 861, 881, 886, 910, 934, 935, 969, 984, 1007, 1008, 1010, 1027, 1072, 1074, 1092, 1138, 1190, 1227, 1315, 1326, 1398, 1413]. interpolation [701, 948, 1362, 1399]. Interpreted [1080]. interval [1387]. interview [1391, 1443]. intrinsic [60]. Introducing [129, 353, 811, 1112, 1119, 1224]. Introduction [44, 395, 397, 417, 419, 474, 628, 654, 674, 760, 822, 942, 1099, 1142, 1173, 1235, 1303, 1514]. Introductory [439, 514, 673, 1246]. intuitive [235]. invariant [1043]. invariants [924]. Invented [5]. Inventing [45]. inversion [1065]. Invert [937]. Investigation [311, 349, 642]. Investigations [25, 1141]. involving [1221]. Ion [272, 563, 1116]. ion-bombarded [272]. ions [923]. Iowa [16, 17]. IRAS [324]. IRIS [585]. Irix [1430]. Iron [1340]. Irvin [672]. ISBN [656, 1173]. Ising [965, 981, 1002]. Ising-like [965]. Ising-model [1002]. isosurfaces [618]. Issue [803, 888, 911, 960]. Issues [432]. Iterated [406]. Iteration [1372]. iterative [1410]. IV [962, 1095, 1098]. Ivory [1341].

J [118, 216, 690, 1099, 1118–1120, 1174, 1283, 1305, 1327, 1354, 1426, 1427]. J. [785, 1100, 1121, 1151]. Jack [1079]. James [894, 1100, 1146]. Jan [1149, 1173]. January [889]. Japanese [56, 591]. Jaroslaw [1148]. JAVA [143, 1237, 1356, 1377, 1379, 1421, 1451, 1520]. Java-Script [1377]. Jeanne [766]. Jefimenko [119]. Jensen [974]. Jerry [1489]. Jiang [1079]. John [760, 845, 1145, 1147]. Johnson [216]. Johnston [785, 1151]. Joins [1365]. Joint [96]. Jointly [383]. Jonathan [762]. Jones [656, 1150]. Jörg [1488]. Jose [422]. Joseph [972, 1145]. Josephson [410]. Joshua [1237]. Journal [399, 879, 1138, 1208, 1228, 1251, 1274, 1297, 1319, 1342, 1368, 1394, 1395, 1419, 1450, 1481, 1507]. Jovanovich [656]. Joy [769]. JPL [20, 226, 247]. Jr [821, 895, 1327]. Jr. [243]. Jul [1447]. Jul/Aug [1447]. Jun [1417]. junctions [410]. June [422].

K. [691, 1260]. Kalos [245]. Karin [43, 1443]. Keeler [1152]. Keeping [1048]. Kehoe [765]. Keith [736]. Keller [868]. Kelly [1281]. Kennedy [1376]. Kepler [1014]. Kernel [890]. Kerr [924]. Kerrigan [1121]. keV [295]. Khokhlov [994]. Kind [415, 1203]. Kinematics [944, 1435]. kinetic [618, 1084]. Kirkpatrick [763]. Klaus [1037]. Knowledge [995, 1069, 1226]. known [901]. Knuth [302]. Koch [1326]. Kochmer [762]. Koelbel [895]. KOH [1386]. Kolmogorov [138]. Kontron [170]. Korteweg [1022]. Kral [672]. Krol [761]. Kuck [1283]. Kuipers [995]. Kuperman [974]. Kutta [565, 891, 912].

L [302, 713, 737, 895, 973, 1173, 1403]. Lab [45, 53, 110, 112, 116]. Laboratories [394, 926, 1400]. Laboratory [39, 78, 79, 473, 719, 739, 872, 998, 1075, 1094, 1370]. Labs [146]. LabVIEW [69, 1334]. Lagrange [701, 948]. Lagrangian [562, 982]. Lanczos [293, 641]. Landau [821]. Landslides [930]. Lane [764]. Langevin [669]. Language [828, 1080, 1123, 1176]. Languages [217, 625]. Lantastic [872]. Laplace [903, 937, 1498]. LaQuey [766]. Large [26, 28, 71, 231, 324, 488, 572, 642, 1143, 1170, 1278, 1330, 1334, 1372, 1519]. Large-Scale [324, 642, 1143, 1170, 1278, 1334, 1372, 1519]. Laser [142, 547, 1086, 1318, 1408]. LASSPTools [1344]. Last [305]. Latest [242, 266, 284, 304, 322, 358, 379, 398, 420, 448, 465, 480, 495, 513, 534, 542, 552, 570, 586, 607, 626, 658, 677, 698, 770, 796, 825, 851, 873, 898, 919, 945, 979, 999, 1020, 1041, 1063, 1083, 1103, 1128, 1156, 1179, 1201, 1219, 1242, 1264, 1288, 1310, 1332, 1492]. Lattice [182, 206, 524, 536, 543, 650, 771, 843, 946, 1214, 1221, 1516, 1525]. Lattice-gas [1516]. Launch [1089]. Launches [150]. Laureate [1069]. Law [1255]. Lawrence [1279, 1302]. LCD [908]. Lead [148, 523, 1271, 1343, 1502]. Leads [590, 1358]. leaf [1429]. Learn [657, 880, 971, 986, 1205]. Learning [568, 583, 648, 773, 951, 1194, 1454]. Least [135, 255, 294, 311, 464, 597, 700, 712, 724, 1132, 1245, 1474]. Least-Squares [135, 294, 311, 464, 700, 712, 724, 1132, 1245, 1474]. Lecture [344, 1168]. Lectures [59, 263, 655]. Ledesma [1015]. Lee [869, 975]. length [1511]. Leonardo [1058]. Let [625, 906]. Lets [1009]. Letter [665, 1506]. Letters [889]. Level [205, 315, 1040]. levels [184, 311]. Levenberg [1245]. lex [828]. Liberal [1324]. Liberal-Arts [1324]. Liberate [1134]. Libraries [22, 141, 967]. Library [17, 19, 70, 1130, 1238, 1241, 1304, 1407]. Lie [1154]. Life [187]. lifetime [801]. light [1373]. Like [357, 728, 965]. likelihood [744]. Limitations [560, 922]. Limits [271, 543, 838]. Line [110, 135, 149, 346, 392, 581, 839, 992, 1331]. Line-Integral [1331]. Linear [433, 600, 681]. Lines [79, 296, 1086]. Lineshape [594]. Link [911]. linked [1244]. Links [68, 932, 1139]. Linsay [822]. Linux [1087, 1095, 1101, 1484]. lipid [1455]. Lipow [767]. Liquid [1487]. liquids [471]. List [1358]. Listen [1108]. Listing [721, 1179]. Listings [242, 266, 284, 304, 322, 358, 379, 398, 420, 448, 465, 480, 495, 513, 534, 552, 570, 586, 607, 626, 658, 677, 698, 770, 796, 825, 851, 873, 898, 919, 945, 979, 999, 1020, 1041, 1063, 1083, 1103, 1128, 1156, 1201, 1219, 1242, 1264, 1288, 1310, 1332, 1492]. Live [1160]. Lived [718]. local [1412]. Localization [1035]. logistic [643]. logit [643]. LogoLinks [1299]. Long [212, 345, 830, 1086, 1374]. Long-range [1374]. Look [640, 835, 956]. Looking [15, 366, 1338]. Looks [1478]. loops [1220]. Loose [1007]. Loosely [696]. Lotus [74, 119]. Lotus-Compatible [74]. Louise [656]. Loveman [895]. Low [313, 585, 606, 711, 923, 1485]. Low-Cost [313, 711]. Low-End [585]. Lyle [691, 1148]. Lynna [1149]. Lyzenga [216].

M [216, 785, 786, 844, 868, 894, 974, 1307, 1354]. M. [1120, 1403, 1426]. M.U.P.P.E.T. [130, 697]. M64 [31]. MA [656]. MAC [34, 61, 109, 113, 192, 569, 1082, 1178]. MacDose [593]. Machine [325, 334, 342, 372, 471, 574, 641, 645, 947, 1079, 1175, 1281]. Machine-2 [645]. Machines [268, 357, 683, 971, 1000, 1046]. Macintosh [29, 384, 436, 458, 461, 719, 829, 897]. Macintosh^{T M} [926]. Macintosh^{T M}-based [926]. Macro [401]. Macromolecules [994]. Macros [416]. MACSYMA [345, 1200]. Made [13, 27, 1236, 1294]. Magazine [1475, 1500]. Magellan [793]. Magnetic [132, 152, 153, 186, 341, 430, 545, 642, 647, 925]. magnetic-field [647]. Magnetism [190, 1148]. magnetization [372]. magnetization-reversal [372]. magneto [826]. magneto-optical [826]. magnetohydrodynamic [1001]. magnetohydrodynamics [136]. Mahoney [713]. Maintain [26]. Mainzer [1037]. Major [519, 1088]. Make [52, 997, 1133, 1178, 1393]. makes [1047, 1087]. Making [823, 1205, 1366, 1414]. Mallat [656]. Man [13]. Man-Made [13]. Manage [544]. Management [78, 252, 290, 717, 842]. Manchek [1079]. Mandelbrot [605]. Manifesto [130]. Manipulating [602, 624]. manipulation [1106]. Mankind [1037]. Mantle [1000]. Manuscript [437]. Many [54]. Map [312, 1058]. Maple [865, 1259, 1328, 1429]. Mapping [324, 1177, 1486]. Maps [91, 108, 404]. March [1416]. Marine [763]. Mark [755, 1307]. Market [268]. Marko [1259]. Markov [278]. Marquardt [772, 1245]. Marriage [1294]. Martin [1281]. Mary [656, 868, 895]. Maryland [1113]. mask [1182]. Masscomp [110]. Massive [536, 560]. Massively [176, 562, 564, 578, 726, 798, 1042, 1044]. Match [794]. MATCOM [1491]. Material [1318]. Materials [173, 202, 424, 730, 756, 815, 853, 1502]. Math [41, 72, 193, 533, 928, 953, 1241]. Math-Software [928]. Math/Science [72]. Mathcad [336, 596, 673, 674]. MATHCAD-based [596]. Mathematica [165, 440, 529, 599, 715, 801, 858, 865, 904, 950, 1085, 1173, 1174, 1353, 1358, 1378, 1489, 1521]. Mathematical [141, 221, 356, 1427, 1436]. mathematically [472]. Mathematica(R) [894]. Mathematica^{T M} [478, 493, 1182]. Mathematician [985]. Mathematics [291, 302, 1257, 1468]. Mathematics^{T M} [533]. MathHDF [715]. MathLink [715]. MathLink-Based [715]. MathStation [191]. MATLAB [192, 1241, 1491]. Matrices [264, 757]. matrix [9, 297, 348, 529, 599, 1131]. Matter [587, 855, 1037, 1045, 1260]. Matthias [1036]. Maxima [404]. maximum [429, 744]. Maxwell [232, 1290, 1409]. May [127, 660, 683, 882, 1126, 1417, 1502]. May/Jun [1417]. MB [36]. Mbits [92]. Mbits/s [92]. McGuire [1328]. MCNC [176]. Mean [458, 808]. Mean-field [458]. Means [323, 410]. Measurement [7, 326, 512, 1255]. measurements [273]. Mechanical [210, 328, 941, 1217]. Mechanics [378, 529, 562, 599, 617, 621, 892, 1019, 1150, 1151, 1180, 1194, 1259, 1267, 1353, 1522]. mechanisms [826]. Media [45, 53, 545, 826, 1157, 1313, 1363]. Medical [172]. Medicine [151]. Meeting [122, 421, 537, 635, 881, 1296, 1416]. Meetings [14, 47, 84, 117, 144]. Megabit [204]. Megabit/s [204]. Melissa [1326]. Mellon [50]. Melt [308]. membranes [1214, 1455]. Memoriam [243]. Memory [31, 64, 67, 113, 229, 551, 560, 1017]. Menus [168]. Message [917]. Message-Passing [917]. Metacomputing [905, 918]. metal [460]. Metcalf [1120]. Metcalfe [1354]. Method [60, 183, 294, 328, 333, 507, 564, 567, 596, 600, 733, 1066, 1078, 1086, 1212, 1221, 1245, 1256, 1260, 1336, 1351, 1361, 1387, 1524]. Methods [9, 44, 251, 293, 369, 408, 508, 530, 546, 723, 734, 772, 785, 786, 891, 894, 912, 963, 966, 967, 973, 991, 1173, 1296, 1301, 1396, 1440]. metric [924]. Meyer [656]. MIBEMOL [802]. Michael [713, 1147]. Micro [401, 1386]. microchannel [504]. Microcomputer [137, 185, 233, 331, 347, 625, 725, 743, 998]. Microcomputer-Based [233, 998]. Microcomputer-generated [331]. microcomputers [182]. microfractures [730]. Micromagnetics [545]. Microprocessor [149, 277]. Micros [72]. Microscope [111, 142, 394]. microwave [273]. Microworld [1267]. mid [86]. mid-1989 [86]. Mie [797]. Migrating [1121]. Millennium [1231]. Million [127, 227]. Millisecond [270]. Mind [1037]. Minds [13, 1282]. Minicomputer [112]. minima [835]. Minimal [572]. Minisupercomputers [89]. Minnesota [125]. MINSQ [167]. mirror [506]. Misner [670]. Mission [1057]. Mistrust [1248]. Mix [794]. mixed [1096]. mixing [525]. MM5 [1223]. mode [255, 744]. Model [63, 100, 238, 291, 502, 580, 641, 680, 727, 921, 964, 981, 1002, 1045, 1223, 1268, 1495]. modeled [1519]. Modeling [99, 124, 283, 364, 410, 411, 500, 525, 604, 605, 609, 613, 846, 902, 927, 995, 1000, 1036, 1133, 1248, 1313, 1375, 1470]. Modelling [966]. Models [21, 191, 221, 464, 759, 848, 870, 931, 975, 993, 1374, 1516, 1525]. Modern [213, 891, 912, 1147, 1197, 1372]. Modified [475, 1388]. Modules [277, 1257]. Mohamad [1124]. Molecular [17, 19, 124, 126, 195, 295, 471, 502, 562, 577, 785, 982, 1116, 1196, 1278, 1303, 1306, 1366, 1375, 1485, 1519]. Molecular-Dynamics [1278, 1303, 1519]. Molecule [563, 572]. molecules [1280]. Moloney [1147]. Momentum [688, 818]. money [1476]. Monitor [385, 423]. Monitoring [385]. Monitors [164]. Monolithic [683]. Monster [757]. Monte [261, 301, 314, 352, 395, 434, 464, 564, 566, 644, 650, 691, 798, 980, 1002, 1064, 1172, 1221, 1260, 1351, 1363]. Moon [822]. Moore [1255, 1476]. Mori [1002]. Morrison [605]. Mosaic [861]. Most [161, 1327]. Motion [237, 733, 914, 1267, 1388]. Motions [179]. Motivated [601]. Motorola [149]. Mouse [319]. Mouseclick [1209]. Move [640, 1418]. Moves [745, 1092, 1341]. Movies [929, 959]. Moving [1364]. MP [86]. MPI1 [917]. Mr. [1316]. MS [533]. MS-DOS [533]. MTL [249]. Much [890]. Müllges [1352]. Multi [432, 528]. Multi-Dimensional [432]. multi-Transputer [528]. multibody [314]. Multichannel [169, 926]. Multidimensional [309, 352, 1115, 1349]. Multidisciplinary [1039]. Multifractal [616]. Multifractals [354]. Multigrid [508, 530]. multilayer [275, 1470]. Multimedia [662, 739, 1257, 1482]. multiparameter [926]. multiparametric [799]. Multiple [292, 595, 1022, 1042]. multiple-instruction-multiple-data [1042]. multiple-processor [292]. Multiplexed [793]. multiprocessing [136]. multiprocessor [368]. Multiprocessors [126, 136]. Multiscale [1513]. Multivariate [1114, 1362]. muon [11, 1335]. Musciano [1376]. Mux [116]. Mysteries [577].

N [627]. NA48 [1335]. Name [988]. Named [245]. Nanocomputers [1477]. nanostructure [642]. Narrative [1402]. NASA [1, 961]. National [146, 225, 310, 706, 729]. Native [219]. Natural [199]. nature [703]. Navier [17, 18]. Navigating [432]. NCGA [362]. NCSA [17, 907]. NCUBE [268]. Near [660, 1498]. Nearly [738, 949, 1291]. Nears [958]. Neat [787]. Need [683]. needle [1195]. needs [782, 881]. Negroponte [1284]. Neou [763]. Neptune [307]. NetLearning [1326]. Network [20, 92, 201, 204, 250, 383, 682, 694, 723, 872]. Networked [1079, 1166]. Networking [441, 764, 766, 781, 794, 1430]. Networks [200, 377, 410, 474, 663, 714, 806, 950, 1124, 1325]. Neumann [28, 1524]. Neural [20, 200, 250, 357, 377, 714, 723, 1124, 1325]. Neuron [1117]. neurons [528]. Neutron [802, 1074]. neutrons [314]. new-products [839]. Newman [371, 924]. News [537, 554]. Newton [605]. Next [56, 224, 300, 342, 360, 1091, 1227, 1272, 1308, 1315, 1329, 1354, 1355]. Next-Generation [1227]. Nextstep [713, 718, 775, 1011]. Ni [618]. Nicholas [1284, 1305]. Niemeyer [1237]. Night [959]. NII [833]. Nine [550]. Ninth [1371]. NIST [752]. Niweek [1348]. NMR [166, 866]. No [662]. Nobel [1069]. noise [1474]. Nominations [1238]. Nonclassical [393]. noncompact [345]. Nonequilibrium [562, 669, 759, 1243]. nonintegrable [1361]. Nonlinear [185, 214, 418, 419, 525, 617, 646, 681, 703, 899, 916, 942, 950, 1132, 1157, 1203, 1245, 1292, 1328, 1521]. Nonlinearity [1456]. nonlocal [1022]. Nonperiodic [1035]. Nonproprietary [611]. nonstandard [1290, 1385, 1409]. Nonstiff [1125]. normal [1105, 1526]. normalizing [597]. Norris [43]. NorthWestNet [762]. not-so-global [832]. Notable [248]. notation [1023]. Note [102, 315, 1246]. Notes [121]. Nothing [501]. Nov [1505]. Nov/Dec [1505]. Novel [546]. Nowhere [501]. NREN [706]. NSF [2, 87, 174, 202, 834, 1090, 1271]. NSFNET [92]. Nuclear [100, 181, 238, 292, 368, 855, 1045, 1145, 1337]. nuclei [99]. nucleon [99]. Null [228]. Null-Strut [228]. Number [259, 505, 622, 643, 830, 1057, 1473, 1528]. Numbers [316, 1105, 1485, 1527]. Numeric [620]. Numerical [9, 134, 156, 283, 321, 343, 388, 447, 580, 678, 707, 742, 792, 899, 901, 937, 967, 973, 1014, 1046, 1055, 1078, 1084, 1153, 1273, 1291, 1301, 1321, 1344, 1352, 1388, 1407, 1428, 1431, 1457, 1493, 1494, 1524]. Numerical-analysis [1431]. Numerically [937]. Numerics [690, 792]. Numerous [787]. Numerov [876, 1336]. Nutshell [1237]. NYSERNet [204].

O [33, 196, 504, 1358]. Ober [767]. Oberon [1240]. Oberon-2 [1240]. Object [387, 518, 609, 611, 630, 713, 771, 792, 1064, 1176, 1226, 1307, 1330, 1404, 1438, 1522]. Object-Based [1307]. Object-Oriented [387, 518, 609, 611, 630, 713, 771, 792, 1064, 1176, 1226, 1330, 1404, 1522]. Objects [426, 613]. observable [642]. Observations [1232]. Observatory [310]. Observing [366]. Obsolete [1088]. Ocean [974]. Oceanography [975]. Oct [1479]. ODE [1125]. off [1221]. off-lattice [1221]. offer [1406]. Offered [1052, 1053]. Offers [588]. Ohio [86]. old [965, 1372, 1512]. older [1172]. Oleg [119]. Olympic [1136]. Omits [574]. On-demand [1390]. One [66, 187, 599, 713, 742, 1181, 1193, 1209]. one-dimensional [187, 599, 742, 1181]. One-Shot [1209]. Online [173, 762, 814, 862, 879, 889, 911, 1190, 1276, 1319, 1422, 1452, 1483]. ons [1460]. OOPP [387]. Open [69, 615]. OpenMol [854]. Operating [292, 409, 840, 1011, 1075, 1095]. Operation [568]. operator [1066]. opinions [1397]. opportunities [705, 1396]. Opportunity [1068]. Optical [4, 54, 60, 229, 282, 445, 594, 660, 826, 835, 850, 1182, 1470, 1494]. Optics [52, 525, 640, 1062, 1146]. Optimal [1412]. Optimization [253, 369, 388, 430, 490, 850, 1114, 1158, 1195]. Optimized [301]. Options [1030, 1056]. orbitals [1104]. orbits [901]. Order [463, 475, 949, 991, 1018, 1022, 1414, 1426, 1493, 1524]. Ordinary [265, 299, 321, 678, 703, 963, 991, 1292]. Oregon [1370]. Oren [302]. Organized [491, 801, 993]. Organizers [636]. Organizes [1263]. orientation [1438]. Oriented [339, 387, 518, 609, 611, 630, 713, 771, 792, 1064, 1176, 1226, 1330, 1404, 1522]. Origin [1127]. Orlando [554]. Orleans [1392]. O’Rourke [972]. Orthogonal [135, 255, 393, 724]. Orvis [118]. OS/2 [840]. oscillators [1493]. Oscilloscope [461]. Other [976]. Otto [216]. Our [380, 762, 952, 960, 1048, 1093, 1205, 1230]. Output [1142]. Overcomes [560]. Overcoming [279]. Oversees [1049]. Overview [686].

P [765, 1515]. pack [41]. Package [71, 73, 74, 80, 168, 194, 1521]. Packages [72, 103, 382, 835, 991]. Packet [1424]. packs [30]. Padé [548, 1021, 1292]. Page [960, 1032, 1052, 1053, 1093, 1167, 1191, 1300]. Pages [1298]. Paint [505]. Paleotechnic [1272]. PANACEA [580]. Panels [908]. Paper [382]. Paradigm [355, 387, 792, 1321]. Paralex [147]. Parallel [90, 147, 150, 174, 176, 207, 211, 244, 260, 452, 489, 504, 511, 523, 558, 562, 564, 571, 574, 578, 595, 627, 683, 696, 726, 738, 779, 798, 817, 917, 946, 980, 982, 992, 1000, 1042, 1044, 1046, 1079, 1117, 1269, 1360, 1404, 1484, 1495, 1497, 1510]. Parallel-Line [992]. Parallel-Processor [90]. Parallelism [536, 560]. Parallelization [1379, 1408]. Parallelizing [1038]. parameter [772]. Parametric [615, 948]. parametrized [311]. Park [1207]. Parker [43]. Parks [671, 784]. Parrinello [726]. Part [44, 568, 583, 602, 624, 817, 840, 843, 848, 870, 891, 912, 936, 962, 963, 991, 1011, 1056, 1075, 1095, 1142, 1143, 1170, 1232, 1279, 1302, 1322, 1329, 1355, 1424]. Partial [321, 681, 849, 1066, 1513]. Particle [97, 179, 295, 428, 727, 925, 1042, 1145, 1158, 1360]. particle-in-cell [727, 1042, 1158]. Particles [914]. Partnerships [1090, 1271, 1418]. Pascal [697]. pass [614]. Passing [917]. Passport [762]. Pasta [1515]. Patashnik [302]. Path [376, 1086, 1106]. paths [1411]. Patrick [670, 1237, 1378]. Pattern [488, 714, 807, 1384]. pattern-forming [1384]. patterns [902, 1182, 1243]. Paul [821, 822, 1173, 1174]. Pay [1206]. PBAR [184]. PC [59, 61, 76, 108, 111, 256, 410, 563, 704, 773, 780, 818, 883, 930, 958, 1015, 1101, 1155, 1163, 1178, 1194, 1255, 1294, 1296, 1337, 1365]. PC-based [256, 410, 773, 1337]. PC-Compatibles [76]. PC-Controlled [111]. PC/AT [563]. PC/Linux [1101]. PCs [40, 68, 196, 844]. Pd [618]. PDB [996]. PDEase2 [849]. PDEs [1323]. Peak [85, 487]. Peck [1237]. Peer [1138]. Peer-Reviewed [1138]. Penalty [1216]. Pendulum [892, 1480]. Penrose [371]. Penumbra [1504]. percolation [350]. Perfecting [53]. Performance [85, 364, 408, 585, 695, 751, 874, 895, 954, 1060, 1136, 1240, 1283, 1329, 1355, 1399, 1408, 1424, 1458]. Performer [1471]. Performing [212]. period [830]. Periodic [188, 348, 901, 1181]. Perl [768]. Personal [262, 313, 378, 795, 1028, 1244]. Perspectives [208]. perturbation [875, 947]. perturbed [1493]. perveance [370]. Petaflops [1162]. Peter [868, 1354].

Petroleum [148]. Pgperl [1189]. Phase [294, 505, 1258, 1280, 1399, 1414]. Phases [502, 1416]. PHENIX [1044]. Phenomena [303, 520, 561, 669, 1445, 1519]. Phenomenology [542]. Philip [736, 786]. Philosophy [320]. Philpott [1145]. Phonon [816]. Phonons [816]. photo [782]. photometry [679]. Phys [1409]. Physical [44, 253, 271, 369, 628, 710, 736, 892, 1173, 1174, 1199, 1301, 1385]. Physicist [205, 968, 1049, 1089, 1378]. Physicists [20, 122, 128, 131, 517, 535, 591, 632, 729, 750, 808, 814, 1082]. Physics [15, 24, 49, 50, 52, 59, 95, 129, 145, 150, 151, 155, 172, 178, 262, 288, 317, 325, 342, 353, 374, 380, 396, 439, 441, 449, 468, 470, 478, 479, 481, 493, 498, 501, 514, 516, 521, 522, 524, 543, 550, 553, 568, 578, 583, 590, 592, 635, 666, 676, 687, 716, 719, 721, 786, 842, 847, 859, 862, 889, 913, 916, 929, 931, 957, 978, 986, 994, 998, 1005, 1027, 1034, 1069, 1099, 1100, 1145, 1147, 1149, 1152, 1161, 1166, 1187, 1211, 1234, 1263, 1279, 1287, 1294, 1302, 1324, 1328, 1339, 1341, 1353, 1475, 1485]. Physics [7, 27, 101, 137, 213, 231, 257, 286, 292, 293, 344, 368, 422, 466, 467, 497, 579, 588, 593, 605, 608, 629, 636, 670, 673, 684, 691, 705, 707, 746, 778, 831, 844, 859, 880, 973, 1016, 1071, 1192, 1212, 1252, 1260, 1305, 1369, 1396, 1400, 1401, 1414, 1422, 1425, 1438, 1451, 1452, 1471, 1488, 1500, 1512]. physlets [1421]. PhysLINK [1423]. Picture [6, 143]. Pictures [1185]. Pipedream [918]. Pitches [174]. Pittsburgh [177, 1418]. placed [329]. Placement [1034]. Plan [49, 122]. Plane [237, 246, 562, 732]. plane-strain [562]. Planet [927, 1300]. Planets [1165]. Planning [87]. Plans [146, 636]. Plasma [341, 424, 707, 1070, 1094, 1383, 1408]. Plasmas [455]. plasmons [1202]. plastic [562]. Plates [511]. Platforms [533]. Playing [1185]. Plots [41, 105, 264, 510, 711]. Plotting [74, 296, 392, 569, 1431, 1434]. Plug [1111]. Plug-in [1111]. plus [1173, 1399]. Point [329, 477, 515, 1458]. points [183, 776]. poisoning [618]. Poisson [903]. poles [237]. polyatomics [1388]. polyexponential [185]. Polygonally [692]. polyhedral [1043]. Polymer [1254]. polymers [1525]. polynomial [724]. Polynomials [393, 841]. Polyp [368]. POOMA [1490]. populated [1525]. Portability [693]. Portable [622, 675, 687, 717, 738, 830, 917]. Porter [974]. Portland [751]. Ports [504, 1155]. Positive [1410]. Positivist [305]. Possibilities [922]. Possible [460, 1444]. Poster [1163]. Posting [1252]. potential [237, 1494]. potentials [348, 1256]. Potts [680]. Power [30, 524, 681, 801, 1082, 1218]. Powerful [483, 574, 591, 638, 675]. Powerstation [1155]. pp [474, 656, 1173]. Practical [239, 558, 868, 918, 937, 1269, 1362, 1474]. Practice [1233, 1391]. Praise [46]. Precautions [1008]. precision [620, 1022]. Predicting [136, 815]. Prediction [975, 1055]. predictions [1086]. Premiere [147]. Preprint [1209]. preprocessor [828]. Presentation [222]. Presents [633]. Press [357, 543, 656]. Preventing [1236]. Prevention [1008]. Previews [960]. price [709]. Primer [605, 764, 1357]. Princeton [17, 18]. Principal [1387]. Principal-value [1387]. Principle [572]. Principles [666, 971, 1375, 1457]. Print [1318]. Printer [817]. Prize [687, 748]. Prize-Winning [687]. prizes [1067]. Pro [671, 740]. Probabilistic [1301]. Probability [628, 1472].

Probability-Based [628]. probes [430]. Problem [343, 373, 523, 842, 967, 1014, 1065, 1212, 1480, 1515]. Problems [95, 139, 216, 293, 508, 530, 546, 1143, 1157, 1170, 1263, 1292, 1301, 1372, 1444, 1457]. procedure [594, 742]. Process [75, 846, 1386]. Processes [166, 628, 966, 1199, 1221, 1403]. Processing [12, 65, 150, 170, 174, 202, 244, 260, 424, 452, 696, 845, 913, 1269, 1359, 1384, 1434, 1484]. Processor [1, 90, 114, 169, 176, 196, 198, 292, 407, 489, 1469]. Processors [147, 216, 1158]. Product [12, 989]. production [99]. Products [720, 741, 839, 862, 888, 935, 1030, 1382, 1518]. Professionals [764]. Professor [953]. Profile [787]. Program [17, 70, 108, 131, 255, 281, 391, 444, 533, 588, 625, 773, 854, 888, 900, 953, 1004, 1265, 1386, 1440, 1445]. Programmable [63, 473, 823]. Programmer [1118]. Programmers [697]. Programming [71, 90, 157, 217, 314, 331, 387, 446, 518, 611, 668, 681, 697, 713, 771, 792, 828, 848, 870, 922, 962, 990, 1241, 1261, 1307, 1356, 1404, 1482, 1520]. Programs [8, 26, 158, 257, 337, 387, 559, 602, 615, 624, 675, 687, 738, 977, 987, 1062, 1403]. Progress [96, 541]. Project [756, 1058]. Projections [992, 1472]. Projectors [908]. Projects [71, 915, 1330, 1418]. Prolog [38]. Promise [13, 718]. Promises [203]. Propagation [949, 975, 1003, 1181, 1313]. properties [238, 1470]. Pro(R) [784]. Prototype [1]. Prove [24]. Proves [612]. Provide [697]. Provides [611]. Prudent [1008]. PS [33, 36]. PS/2 [33, 36]. Pseudorandom [1057, 1526]. pseudospectral [333, 1022]. Publication [80]. Publication-Quality [80]. Publications [961]. Publishing [879, 1025, 1026, 1031]. Pulsar [269, 270]. pulse [926, 1244]. Pulsed [547]. Pumping [1340]. Purpose [7, 310, 473, 1175, 1360]. Pursuing [1515]. Puts [1188]. PVM [694, 1079]. Python [1153, 1176].

QCD [206, 536]. Quadrature [393, 1115]. Qualitative [995]. Quality [80]. Quantifying [318, 426, 774]. Quantitative [1464]. Quantities [1402]. Quantum [210, 328, 348, 378, 395, 434, 500, 564, 599, 617, 621, 650, 689, 734, 798, 843, 894, 941, 1019, 1078, 1151, 1217, 1259, 1295, 1361, 1367, 1411, 1426]. Quantum-Field [689]. Quantum-Mechanical [210, 941, 1217]. Quark [855, 1045]. Quark-based [855]. quarks [595]. Quasi [316, 1181]. Quasi- [316]. quasi-periodic [1181]. quasicrystal [1129]. Quasicrystals [509]. quaternion [1131, 1388]. Quattro [671, 784]. quenched [680]. Question [450]. Quick [667]. Quieter [267]. Quotient [1372].

R [713, 736, 868, 975, 994, 1151, 1173, 1174, 1402]. R&D [985]. Rack [112]. radial [876]. radiating [924]. radiation [333, 1022]. radiative [9]. Radio [310, 727]. radiological [593]. radionuclides [260]. Rahman [748]. Randall [1150]. Random [259, 316, 350, 622, 628, 643, 830, 950, 1057, 1105, 1473, 1527, 1528]. random-number [1473]. Randomness [604, 605]. Randy [1307]. range [28, 1374]. Rapaport [1196, 1303]. Raphael [656]. Rapid [934]. rare [460, 1159]. rare-gas [1159]. rate [254]. rates [1221]. Ray [11, 256, 275, 598, 653, 1062, 1220, 1337, 1464]. Ray-Optics [1062]. Ray-Tracing [598, 653]. Rayleigh [1372]. Reaction [618, 921, 1084]. Reactions [314, 623]. Reader [911, 1112, 1397, 1480]. Readers [906, 1139]. readout [1335]. Ready [128]. Real [16, 17, 298, 484, 592, 733, 746, 1038, 1056, 1081, 1236, 1286, 1462, 1471]. Real-Time [484, 1056, 1081, 1286, 1462, 1471]. Reality [121, 592, 754, 918, 955, 1367]. Realm [228]. Reasoning [995]. Rebounds [1200]. Receives [805]. recipe [1372]. Recipes [1321]. Recognition [714, 1028, 1114]. Reconstruction [992]. reconstructive [1399]. Record [985]. recording [545, 826, 853]. Recursion [1137]. Recursive [28, 352, 1096]. Red [1250]. Redistribution [964]. reduced [1158]. Reduction [292, 736]. Redundancy [280]. Redwine [1122]. reference [1215, 1436]. Reflection [445, 1181]. Reflections [1347]. reflectors [275]. reform [748]. refraction [445]. register [1473]. regression [723]. Regular [216]. Reid [1120]. Related [912]. relativistic [1014]. Relativity [228, 451, 484, 827]. relax3d [903]. relaxation [273]. Releases [242, 266, 284, 304, 322, 358, 379, 398, 420, 448, 465, 480, 495, 513, 534, 552, 570, 586, 607, 626, 658, 677, 698, 770, 796, 825, 851, 873, 898, 919, 945, 979, 999, 1020, 1041, 1063, 1083, 1103, 1128, 1156, 1179, 1201, 1219, 1242, 1264, 1288, 1310, 1332, 1492]. Relies [541]. Rely [1248]. Remaining [936]. Remember [984]. Remote [385, 1420]. Remove [639]. Renderer [492]. Rendering [102, 692, 793, 1262]. Renders [104]. Renehan [1327]. Reno [1110]. Reorganized [1167]. repeller [102]. Replace [858, 1090, 1475]. Report [362, 457, 911, 1348, 1416]. Repositories [1209]. Representing [1232]. Requested [83]. Research [126, 133, 180, 208, 383, 425, 553, 608, 872, 887, 932, 1003, 1013, 1049, 1070, 1097, 1420, 1508, 1510]. Researchers [125, 520, 539, 639, 640, 836, 956]. resistor [950]. Resolution [34, 164, 563, 596, 645, 946, 1455]. Resonance [152, 153, 186, 617]. resonant [273, 1220]. Resource [519, 984]. Resources [591, 750, 871, 1491]. response [275, 369, 1045, 1159]. Rest [841]. Restoration [171]. Retinex [178]. Retrieval [74]. Return [1185]. Reusable [1451]. Reveal [576, 730]. reversal [372]. Reversed [1070]. Review [43, 45, 118, 119, 302, 357, 627– 629, 654, 670–674, 690, 691, 713, 719, 736, 737, 760, 762–767, 784– 786, 821, 822, 844, 845, 868, 869, 894, 916, 942, 969– 975, 995, 1015, 1036, 1037, 1079, 1099, 1118–1124, 1145, 1146, 1148– 1150, 1152, 1196, 1197, 1236, 1260, 1283, 1284, 1303, 1305, 1307, 1326– 1328, 1353, 1354, 1375–1378, 1402, 1403, 1426, 1456, 1457, 1488, 1516]. Review [216, 474, 895, 915, 930, 994, 1147, 1151, 1174, 1215, 1237, 1259, 1281, 1282, 1306, 1352, 1489, 1515]. Reviewed [1138]. Reviews [605, 655, 656, 761, 1100, 1173, 1397, 1427]. Revisited [1060, 1061, 1486]. Revolution [267, 518, 1089, 1255, 1365]. Revolutionizing [1005]. Reynolds [1485]. Rhythms [386]. Rich [1324]. Richard [713, 915, 1173, 1174, 1328]. Richardson [1266]. Rides [769]. Right [686, 817, 864, 962, 990, 1056, 1322, 1364, 1424]. rigid [1388]. RISC [149, 595]. Rise [1299]. Robert [671, 737, 784, 895, 942, 971, 1079, 1148, 1354, 1488].

Roberta [1145]. Roberts [1307]. Robin [1146]. Robinson [736, 975]. Rochester [482]. Rocket [500]. Roelofs [691, 1148]. Roger [1152]. Role [514]. Roles [1135]. Rollins [1152]. ROM [48, 1173, 1436]. Romania [381]. Ron [1015]. Ronald [302, 656, 1148]. Roots [667]. Roscoe [1391]. rose [884]. Rothberg [1145]. Rothman [1516]. Rough [872]. Round [12]. Round-up [12]. Roundtable [93]. Route [749]. routine [10, 236, 427, 487]. Roy [767]. Rubin [821]. Run [1, 1497]. run-time [1497]. Runge [565, 891, 912]. Running [51]. Runs [106]. Ruskai [656]. Russ [845]. Russians [957]. Ruth [1036]. Ryer [766].

S [92, 204, 216, 786, 822, 895, 1118, 1150, 1236, 1528]. Sadhan [1457]. safe [1404]. Safety [423]. Sake [952]. salesman [1212]. Salmon [216]. Sample [1343]. Sampling [241, 254, 352]. Samuel [786]. San [422, 656, 1271, 1277]. Sandia [175]. sandpiles [459]. Sane [936]. Santa [1140]. SATAN [1007]. satellites [829]. Sather [611, 1060]. saturable [411]. Savitzky [435]. Savitzky-Golay [435]. SC’93 [751]. SC97 [1317]. SC98 [1476, 1478]. Scalable [806, 946]. Scalar [486, 924]. Scale [324, 642, 1143, 1170, 1278, 1334, 1372, 1519]. scaler [926]. scales [1511]. scaling [135, 616]. Scanning [111, 394]. Scans [142]. scattered [797]. Scattering [99, 210, 237, 333, 506, 1003, 1059, 1457]. Scene [896]. Scheduling [250]. scheme [1132, 1291, 1399, 1426]. Schemes [406, 1084]. Scherrer [488, 668]. Schildt [1215]. Schmidt [974]. Scholarly [1026, 1029]. School [88]. Schreiber [895]. Schrödinger [388, 526, 742, 925, 1440, 1524]. Science [6, 22, 72, 105, 172, 173, 225, 257, 302, 320, 461, 627, 634, 984, 1005, 1010, 1068, 1088, 1231, 1255, 1257, 1327, 1390, 1398, 1442, 1500]. Sciences [1, 710, 736, 939]. Scientific [26, 29, 58, 70, 73, 92, 119, 140, 209, 217, 220, 252, 289, 290, 384, 469, 519, 638, 671, 697, 717, 784, 828, 915, 943, 967, 1025, 1031, 1048, 1101, 1130, 1138, 1171, 1198, 1236, 1261, 1330, 1346, 1356, 1404, 1420, 1441, 1465, 1517, 1520]. Scientist [94, 821, 997, 1006]. Scientists [88, 118, 239, 307, 672, 674, 690, 822, 942, 1030, 1328]. scintillation [1337]. Script [1377]. SCRs [115]. SCS [64]. SCSI [215]. SDSC [17, 19, 124]. Search [85, 188, 487, 1009, 1165]. Searchable [1275]. Searches [832, 1209]. Second [89, 415, 532, 949, 1018, 1203, 1215, 1260, 1376, 1377, 1426]. Second-Generation [89]. Second-Order [949, 1018, 1426]. Section [839, 960, 1208, 1228, 1251, 1274, 1297, 1319, 1342, 1368, 1395, 1419, 1450, 1481, 1507]. Secure [1276]. Security [1008]. See [477, 1230]. Seek [91, 639, 836]. Seeks [1090]. seen [706]. SEGS [168]. Segue [1499]. Seismic [91, 230, 938]. Seismic-Wave [938]. Selected [1353]. Selecting [840, 1011, 1075, 1095]. Self [491, 618, 801, 993, 1453]. self-gravitating [1453]. Self-Organized [491, 801, 993]. self-poisoning [618]. Semi [235]. Semi-intuitive [235]. Semiclassical [1045]. semiconductors [594, 982]. Send [804]. Senses [228]. Sensitivity [646, 1223]. Sep [1479]. Sep/Oct [1479]. separation [430]. sequence [1203, 1473]. sequential [594]. serial [680]. Series [429, 472, 656, 681, 867, 886, 1021, 1203]. Serim [1326]. Serve [542]. Server [68]. Service [814, 872, 1112, 1209]. Services [811]. sessions [721]. Set [856, 1158]. Seth [714]. Sets [315, 432, 842, 939]. Setting [193, 1162]. Seventh [1213]. Several [602, 624, 724]. SGML [1031]. Shared [560]. Shared-Memory [560]. sharing [136]. shattering [315]. Shear [758, 1070]. Sheets [935]. shell [100]. shift [1399, 1473]. shift-register-sequence [1473]. Shifts [294, 355, 1262]. Shines [897]. Ship [89]. Shock [515, 561]. Shock-Wave [561]. shop [1504]. Shot [1209]. Should [539, 573, 729, 1051]. Showcases [779]. Sidney [43, 1443]. Sierpinski [405, 616]. SIGGRAPH [287, 633]. Sights [1162]. SigmaPlot [569]. Signal [114, 169, 452, 1359]. Signal-processing [1359]. Signals [188, 776]. Silbar [713]. silicon [1386]. Silsbee [1488]. Silver [853, 1197]. silver-halide [853]. Silverman [1197]. Simple [215, 277, 313, 405, 502, 559, 924, 993, 1243, 1275, 1291, 1387, 1516]. simplex [594]. simplified [598, 1406, 1495]. Simplify [1241]. Simson [713]. simulate [314, 409]. Simulated [253, 275, 490, 582, 679, 730, 855, 1024]. Simulating [109, 327, 453, 725, 1091, 1373, 1525]. Simulation [16, 17, 19, 44, 46, 121, 158, 246, 261, 263, 272, 326, 347, 365, 372, 410, 455, 464, 471, 528, 562, 580, 593, 623, 642, 691, 702, 719, 785, 866, 950, 965, 995, 1173, 1180, 1196, 1220–1222, 1280, 1306, 1333, 1351, 1360, 1363, 1383, 1386, 1432, 1435, 1463, 1503, 1511, 1515]. Simulations [51, 122, 295, 341, 345, 456, 500, 515, 541, 543, 550, 576, 601, 628, 645, 669, 756, 798, 853, 914, 938, 980, 982, 1003, 1019, 1040, 1064, 1117, 1145–1152, 1172, 1174, 1214, 1234, 1246, 1278, 1303, 1325, 1455, 1482, 1487, 1488, 1494, 1497, 1519]. Simulator [195]. Sine [649]. Single [97, 99, 614, 1469]. Single-Particle [97]. single-pass [614]. Single-system-image [1469]. singular [1291]. SISAL [90]. Sisyphus [1289]. Site [1009, 1073, 1318]. Sites [923, 1139, 1271]. Sixth [1077, 1524]. sixth-order [1524]. Sized [469]. Skew [187]. slave [277]. Sleep [1111]. Sleep-in [1111]. Sleightholme [1119]. Slowing [279]. Small [249, 487, 901]. Smart [115]. Smirnov [138]. Smit [1306]. Smith [43]. Smog [363]. Smoothers [890]. Smoothing [10, 236, 389, 435, 890]. Smoothly [1155]. Social [469]. Society [363, 421, 1005, 1135]. Soda [469]. soft [1220]. Software [29, 39, 50, 59, 69, 75, 77, 97, 98, 103, 166, 175, 190, 205, 247, 251, 286, 330, 410, 413, 518, 521, 522, 532, 542, 550, 554, 631, 652, 717, 777, 791, 799, 854, 928, 940, 951, 963, 967, 991, 1030, 1049, 1067, 1077, 1142, 1169, 1187, 1204, 1210, 1213, 1232, 1239, 1244, 1253, 1269, 1335, 1350, 1371, 1407, 1414, 1431–1435, 1445, 1470]. Software-Contest [1204, 1210]. Sol [674]. Solar [233, 515, 725, 900, 1220]. Solaris [1015]. Solid [500, 1152, 1180, 1488]. Solids [966]. solitary [1022]. Solution [134, 165, 231, 321, 343, 388, 678, 681, 903, 1014, 1065, 1157, 1290, 1409, 1426, 1440, 1457, 1497, 1513, 1524]. Solutions [60, 619, 818, 956]. solve [925]. Solver [1125, 1263]. Solves [849, 1263]. Solving [214, 216, 336, 523, 775, 967, 1066, 1212, 1292, 1323, 1372]. Some [211, 429, 728, 745, 830, 1056, 1347, 1528]. Something [630]. Sonic [515]. sort [614]. sorting [349]. Sound [107]. Sounding [584]. Source [789, 813]. sources [460, 1373]. Soviet [209]. SP1 [946]. Space [1, 246, 250, 270, 332, 351, 619, 1211]. Spaced [188]. Spaces [309, 490]. Spacetime [494]. Spanning [1511]. Sparse [600]. spatial [1384]. speakers [721]. Spec [935]. Special [310, 827, 1175, 1360]. Special-Purpose [1175, 1360]. Specialists [663, 777, 781]. Specialized [1444]. spectra [1373]. Spectral [429, 801, 1066, 1408]. spectrometer [802]. Spectroscopy [153, 313, 563, 1265, 1337, 1464]. Speech [425, 1028]. Speed [114, 231, 614, 794, 838, 1347]. Speeds [175]. Spell [90]. Spelling [841]. sphere [507, 797]. spherical [28, 237, 390, 1046]. spherically [875]. Spicklemire [1152]. Spikes [115]. spin [211, 981]. Spinor [330]. Spline [258, 615, 1474]. spline-fitting [615]. splines [948]. splitting [1066]. Spoken [1108]. Spomick [1149]. Sponge [405]. Spontaneous [964]. Spotlight [1207]. Spotlights [954, 1503]. spray [1383]. SpreadSheet [40, 189, 281, 337, 416, 464, 567, 670, 672, 874, 1265]. Spreadsheets [95, 101, 214, 411, 428, 510, 671, 784]. Springer [1173]. Springer-Verlag [1173]. Spur [126, 132]. spurious [776]. Square [667]. Squares [135, 255, 294, 311, 464, 597, 700, 712, 724, 1132, 1245, 1474]. Squeezes [112]. Squelch [115]. SSC [609]. stability [60, 1453]. Stabilizing [1266]. Stable [797]. Stan [1282]. Standalone [1491]. Standard [693, 917]. Stanley [605]. star [699]. Stars [1074]. Start [872, 1412]. started [763, 1336]. Stat [194]. State [60, 128, 211, 1152, 1347, 1488]. State-Of-The-Art [128]. statistic [427]. Statistical [103, 691, 710, 994, 1149, 1212, 1460]. statistical-physics [1212]. Staudenmaier [844]. Stauffer [605]. Stay [814]. Steady [60]. Steady-state [60]. Steele [895]. steering [1497]. Step [713, 742, 1493]. Stephane [656, 1516]. Stephen [970]. Stepping [1367]. Steps [887, 1221]. Stepsize [565]. Stereo [976]. Steven [916, 1152]. Stewart [45]. sticky [278]. Stiff [265, 1125]. Stiff/Nonstiff [1125]. Still [1340]. stochastic [702]. Stock [535]. Stokes [17, 18]. Stoner [1148]. storage [255]. Stores [124]. STP [62]. STP-2M [62]. Straight [581]. strain [562]. strategies [311, 1245]. Stratified [352]. streams [212]. stretched [899]. Striking [1206]. String [899, 1171]. strings [156, 1401]. stripping [185]. Strogatz [916]. Structure [297, 406, 572, 743, 882, 1131, 1453]. Structures [538, 1181, 1366, 1446]. Struggling [921]. Strut [228]. Student [25, 831, 1058, 1100, 1267]. Students [50, 107, 129, 657, 733, 769, 1034, 1205, 1234, 1295]. Studies [583, 596, 807, 1254, 1456, 1480]. study [235, 595, 723, 1222, 1243]. Styer [1151]. style [157]. Sub [316]. Sub- [316]. Subcommittee [87]. Submission [780, 958]. Submissions [863]. subscribers [1509]. Subscription [1276]. subsets [1406]. Successful [953]. Suggest [882]. sum [1411]. sum-over-paths [1411]. summation [472]. Summerhill [764]. summing [1203]. Sun [391, 443, 1165]. SUN-based [391]. Sunderam [1079]. Super [469, 1386]. super-micro-etching [1386]. Super-Supercomputer [469]. Supercollider [122]. Supercomputer [17, 18, 30, 43, 88, 127, 175, 340, 469, 537, 578, 591, 1090, 1206, 1277]. Supercomputers [50, 56, 64, 140, 181, 234, 363, 798, 821, 907, 1042]. Supercomputing [2, 87, 177, 663, 779, 806, 834, 1271, 1418].

Superconducting [410, 660]. Superconductivity [146, 639]. Superconductors [816]. Superhighway [933]. Superlattice [203]. SUPERNET [175]. SuperQuest [88]. Supplies [197]. Supply [1218]. Support [561, 1062, 1379, 1466]. Supported [1207]. Supports [77, 521, 536, 694, 1509]. Surface [476, 1202, 1463]. Surfaces [186, 272, 618, 921, 966, 1394]. Survey [469, 911]. Sutherland [147]. SwissHPC [883]. switch [1414]. switching [1424]. SWRI [1211]. Symbol [602, 624]. Symbol-Manipulating [602, 624]. Symbolic [1, 521, 579, 634, 668, 681, 818, 875, 878, 950, 1023, 1043, 1076, 1084, 1106, 1358, 1439]. Symbolics [856]. symbols [1523]. Symmetric [403, 774, 875]. symmetries [1043, 1496]. symmetry [876, 923, 1043]. Symposium [754]. System [78, 79, 109, 128, 233, 256, 274, 290, 292, 366, 473, 474, 505, 512, 563, 568, 596, 621, 641, 646, 711, 790, 799, 800, 802, 824, 840, 896, 926, 992, 1011, 1075, 1095, 1334, 1335, 1384, 1469, 1471]. Systems [24, 44, 57, 199, 211, 229, 231, 328, 350, 369, 391, 406, 419, 528, 600, 604, 605, 612, 628, 686, 702, 703, 734, 737, 759, 788, 798, 801, 846, 875, 965, 1035, 1101, 1173, 1174, 1243, 1283, 1309, 1360, 1361, 1424, 1453, 1456, 1521]. Systems-Natural [199]. Systolic [125].

T. [869]. T3D [1038, 1311]. tables [253]. Tackle [95]. Take [523, 535, 952, 987, 1126]. Takes [592, 769, 1027]. Taking [451]. Talk [1178]. Talking [1504]. talks [721]. Tam [1378]. tap [1473]. Target [516]. tasks [409]. Taylor [1412]. Tcl [977]. Tcl/Tk [977]. Teach [442, 668]. Teachers [49, 986, 1071, 1110, 1326]. Teaches [866, 1435]. Teaching [131, 145, 182, 433, 778, 827, 926, 944, 1171, 1246, 1411]. Technical [12, 1380, 1434]. Technical-plotting [1434]. technique [902]. Techniques [182, 216, 310, 314, 987, 1330, 1502, 1503, 1522]. Technologies [396, 794, 976, 1430]. Technology [640, 764, 952, 957, 1227, 1508]. Teleconference [124]. Telemedicine [955]. Telescope [250, 256]. Telnet [989]. Temperature [273, 680, 723, 816]. temperatures [159]. template [8]. Templates [1216, 1285]. temporal [1384]. Ten [655, 1308, 1339, 1417, 1447, 1479, 1505]. Tennant [767]. Tension [476]. Tensor [878, 1018, 1023, 1439]. tensorial [1522]. Tenth [1225, 1239]. Teraflop [340]. Teraflops [930, 1250]. terms [923]. Terrain [492, 793]. Test [138, 359, 361]. testing [1396]. Text [756, 1031]. Textures [1177]. Their [60, 656, 714, 757, 966, 1136, 1523]. Them [1205]. Themes [1006]. Theodore [1489]. Theorem [241]. theoretic [373]. Theoretical [27, 262, 605, 1353]. theories [345]. Theorist [1270]. Theory [178, 180, 182, 227, 245, 486, 525, 656, 689, 737, 771, 843, 947, 1233, 1391, 1411, 1496]. There [700, 1319]. Thermal [203, 1149, 1159]. thin [460, 545, 642]. Thinking [268, 630, 1037, 1185]. Third [42, 652]. Thomas [1515]. Thompson [690, 1427]. Thorax [576]. Threats [608]. Three [66, 251, 298, 407, 661, 772, 848, 870, 920, 921, 1042, 1062, 1217, 1309, 1313, 1408, 1459]. Three-Dimensional [407, 848, 870, 920, 1042, 1313, 1408, 1459]. Three-in-One [66]. Threshold [767]. tight [982]. tight-binding [982]. tiling [431]. Tilings [509]. Time [16, 17, 136, 210, 369, 429, 484, 526, 619, 733, 752, 802, 925, 949, 1003, 1056, 1081, 1183, 1193, 1286, 1406, 1462, 1471, 1497]. Time-dependent [210, 526, 925]. time-domain [949, 1003]. time-of-flight [802]. time-sharing [136]. Titan [195]. TK [977, 1263]. TMS32020 [32]. Tobochnik [1149, 1173]. Today [571, 794]. Together [933]. Tokamak [612]. tokamaks [707]. Tomography [154, 230]. Too [162, 573]. Tool [468, 638, 944, 1003, 1019, 1468]. Toolkit [697, 977, 1359]. Tools [330, 353, 547, 625, 747, 836, 1381]. Top [805, 1445]. Topic [1171]. Topology [1018]. total [255, 726]. Tour [367]. Tower [1341]. towers [102]. Tracing [598, 653]. Tracking [829, 901, 1189]. Tracks [944]. Tracy [766]. Traditional [1088]. trajectories [428]. Tramp [790]. Transfer [348, 599, 812]. transfer-matrix [599]. Transfers [204, 215]. Transform [334, 643, 866, 1033]. transformations [332, 1203]. Transforming [1193]. Transforms [212, 298, 732, 937, 1100, 1144]. Transition [9, 460, 1265, 1414]. transitions [594, 1258]. transmission [348]. transmitted [797]. Transport [486, 644, 846, 1426]. Transportable [70]. Transputer [528, 853]. Transputer-based [853]. Traps [1116]. traveling [1212]. traveling-salesman [1212]. Tree [1226]. tridiagonal [231]. Trig [667]. trigger [1044]. TRIMM [489]. Trips [1074]. trouble [852]. True [976]. truncation [312]. Try [630, 663, 728]. TTC [878, 1439]. Tufte [1402]. Tune [1111]. Tune-in [1111]. Tunneling [111, 394, 941, 1078]. Turbulent [645, 758, 776]. Turing [359, 361]. Turn [22]. Turning [172]. Turns [953, 1429]. Tuszynski [1148]. Tutor [50]. Tutorial [656, 1079]. Twenty [421]. Twenty-Fourth [421]. twice [709]. twisted [351]. Two [72, 138, 187, 210, 298, 309, 703, 902, 1222, 1224, 1438]. two-beam [902]. Two-Dimensional [138, 187, 309, 1222]. Type [193, 527]. Types [77].

Uhlig [1352]. Ulam [1515]. Ultrasonic [512]. ultraviolet [563]. unbounded [1292]. uncertainties [772]. Uncovering [505]. undamping [1245]. Undergird [553]. Undergrad [49]. Undergraduate [11, 353, 374, 394, 461, 470, 892, 1088, 1166, 1425, 1508]. Undergraduates [1013]. understanding [593, 831, 1306]. Unevenly [188]. Unicos [1430]. Unified [568, 583]. uniform [183]. Uninterruptible [1218]. Union [209]. Uniting [1005]. Universe [324, 453, 769, 1415]. University [174, 268, 1370]. Unix [1075, 1344, 1357]. Unravels [577]. Unveils [149, 224]. Update [1073, 1393, 1394]. Updated [533, 1260]. Updates [1191]. Upgrade [92]. Upgrading [1122]. Upper [353, 374, 1040]. Upper-Division [353, 374]. Upper-Level [1040]. Uranus [48]. Use [159, 314, 317, 345, 461, 547, 992, 1101, 1326]. Used [341]. Useful [571, 578, 612, 997, 1327]. User [432, 761, 1001, 1079, 1130, 1142]. user-friendly [1001]. Users [68, 443, 516, 1287, 1406]. Uses [556, 613, 1062]. Using [23, 137, 182, 185–187, 189, 249, 336, 369, 409, 411, 464, 472, 473, 594, 595, 599, 642, 645, 646, 679, 743, 744, 817, 819, 828, 842, 844, 849, 890, 967, 977, 996, 1023, 1046, 1072, 1080, 1086, 1097, 1106, 1180, 1182, 1216, 1223, 1246, 1259, 1290, 1298, 1323, 1331, 1334, 1337, 1345, 1384, 1405, 1408, 1409, 1413, 1421, 1471, 1472, 1484, 1508, 1513, 1520, 1522, 1525]. Utilities [697]. Utilizes [934]. UVR [755].

V [233, 594, 990, 1403, 1429]. Vaidy [1079]. Value [508, 530, 546, 913, 1157, 1292, 1387]. variable [135, 643, 742, 1361, 1493]. variable-step [742, 1493]. variables [724]. Variance [306, 555]. variates [1526]. Variational [1157, 1457]. Varied [588, 1062]. Variety [349, 1296]. varying [1221]. Vascular [1141]. VAX [68]. Vector [198, 638, 1256, 1331]. Vector-Bundle [638]. Vendors [932]. Ventra [1426]. venture [1442]. verification [1043]. Verlag [1173]. Versatile [563]. Version [479, 1422, 1452, 1489]. Very [453, 830, 1424]. very-high-performance [1424]. very-long-period [830]. Vesely [1099]. VI [1056]. via [562, 681, 917]. Vibrating [1171, 1401]. Vibrational [744]. Video [37, 124, 733, 913]. Vie [632]. View [1161]. Viewing [285]. VII [1322]. VIII [1424]. Vinci [1058]. Virtual [592, 754, 1000, 1046, 1074, 1079, 1163, 1370, 1397, 1433]. Visiometrics [426, 1199]. Vision [178]. visual [868, 922, 1402]. Visualization [46, 93, 94, 274, 287, 290, 291, 367, 412, 443, 468, 477, 486, 517, 561, 603, 638, 653, 692, 715, 793, 868, 920, 997, 1039, 1081, 1102, 1126, 1129, 1154, 1220, 1262, 1286, 1289, 1308, 1380, 1383, 1405, 1438, 1459, 1468, 1471, 1497]. Visualizations [716, 939]. Visualize [983, 1331]. Visualizing [288, 426, 507, 617, 699, 1104, 1217, 1293]. Vivian [763]. VLSI [20]. VMEbus [1467]. Voice [1048]. Voigt [787]. Vol [216]. Voltage [115]. Volterra [415]. Volume [603, 653, 656, 659, 692, 1262]. volumetric [1159]. Vortex [1199]. Vortex-Dominated [1199]. Voyager [48, 307, 769]. Vries [1022]. vs [683].

W [627, 974]. W. [1118, 1123]. walk [350]. Walker [216, 759]. Wallace [1528]. War [412]. Ward [763]. Was [1347]. Watch [857]. Watcher [712]. Water [283]. Waterloo [865]. Wave [56, 98, 333, 472, 561, 876, 938, 949, 1003, 1194, 1313, 1333]. waveform [254]. Wavefunctions [1217]. Wavelet [656, 981]. Wavelets [651, 654–656, 1085, 1144, 1323, 1513]. Waves [107, 283, 515, 711, 1022, 1087, 1097, 1146, 1181, 1393]. Way [124, 590, 1028]. Weak [188, 1183]. weakly [1022]. Weather [931, 1047, 1055, 1223]. Web [869, 911, 960, 989, 1009, 1025, 1029– 1031, 1050, 1073, 1093, 1139, 1164, 1189, 1229, 1252, 1275, 1298, 1318, 1320, 1327, 1345, 1392, 1396, 1421, 1425, 1446, 1454, 1508]. Web-based [1396, 1421]. Web-enhanced [1425]. Website [1422, 1452, 1509]. Websites [1327]. Weicheng [1079]. Weight [393, 712]. Weight-Watcher [712]. weights [1132]. Weissert [1515]. Well [1040]. Wellin [1173, 1174]. Welsbuch [474]. Wendt [760]. Wenes [821]. Wesley [474]. Wheeler [134]. Where [29, 61, 103, 141, 164, 190, 217]. Whitney [628]. Whittaker [1021, 1203]. Who [5, 1108, 1316]. Whole [761]. Wide [869, 989, 1454]. widely [1221]. Wieder [674]. Will [90, 120, 226, 587, 722, 808, 1040, 1108, 1163, 1247, 1475]. William [118, 690, 1281, 1427]. win [1067]. Window [977]. Windows [824, 1287, 1334, 1465]. Wing [281]. winner [748]. Winners [413, 532, 652, 940, 1077, 1213, 1350]. Winning [687]. Winter [635]. WIS [1094]. within [183, 966]. Without [1348, 1482]. Woes [90]. Wolfram [970]. Wolpert [714]. Wong [786]. Word [12, 740]. Wordperfect [740]. Work [120, 145, 375, 555, 1058, 1188, 1436]. Workbench [657]. Workbook [690]. Working [997]. Works [1268]. Worksheet [76]. workshop [1071]. Workstation [170, 224, 452, 585, 699]. Workstations [67, 519, 821, 1101]. World [21, 285, 380, 402, 762, 869, 989, 1164, 1168, 1348, 1442, 1454]. worldline [564]. wormhole [904]. Would [728]. Write [345, 738]. Writing [8, 612].

X [275, 551, 977, 1220]. X-arRAY [551]. X-ray [275, 1220]. X-Window [977]. XML [1446].

Y-MP [86]. Ya [1403]. yacc [828]. Yam [1456]. Yaneer [1456]. Year [1225, 1234, 1505]. Years [1272, 1308, 1339, 1354, 1417, 1447, 1479]. Yields [175, 314, 1256]. Yorick [1080]. York [1173]. Yourself [578]. Yu [994]. Yves [656].

Z [281]. Zaleski [1516]. Zen [765]. zeros [237]. Zosel [895]. Zwanzig [1002].

References

Batacan:1987:NSP

[1] Peter Batacan. NASA symbolic processor prototype to run space sciences applications. Computers in Physics, 1(1):9–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

      //aip.scitation.org/doi/10.1063/1.4903441.

Davis:1987:FNS

[2] Stephen G. Davis. Focus on NSF supercomputing. Computers in Physics, 1(1):18–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

      org/doi/10.1063/1.4903429.

Anonymous:1987:BF

[3] Anonymous. The beauty of fractals. Computers in Physics, 1(1):26– ??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

      4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
      4903430.

Hecht:1987:ODE

[4] Jeff Hecht. Optical design enters the Computer Age. Computers in Physics, 1(1):34–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903431.

Slater:1987:WIC

[5] Robert Slater. Who invented the computer? Computers in Physics, 1 (1):44–49, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4903432.

King:1987:GDN

[6] John F. King. Graphics draw a new picture for science. Computers in Physics, 1(1):50–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903433.

Spencer:1987:GPM

[7] C. D. Spencer, P. Seligmann, and D. A. Briotta. General-purpose measurement interface for physics experiments. Computers in Physics, 1(1): 59–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4903434.

Carlson:1987:TWP

[8] Edward H. Carlson. A template for writing programs. Computers in Physics, 1(1):65–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903435.

Gao:1987:NMF

[9] Bo Gao and Anthony F. Starace. Numerical methods for free — free radiative transition matrix elements. Computers in Physics, 1(1):70– ??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4903436.

Hayden:1987:DSR

[10] Howard C. Hayden. Data smoothing routine. Computers in Physics, 1(1): 74–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4903437.

Sutton:1987:UCR

[11] C. Sean Sutton, Douglas A. MacIntire, Susan J. Egan, and Anne Caraley. Undergraduate cosmic ray muon decay experiments with computer interfacing. Computers in Physics, 1(1):76–??, November 1987. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4903438.

Mathes:1987:PRT

[12] Leslie Mathes. Product round-up: Technical word processing. Computers in Physics, 1(1):81–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903439.

Hardt:1987:MMM

[13] Shoshana L. Hardt. Man-made minds: The promise of artificial intelligence. Computers in Physics, 1(1):91–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4903440.

Anonymous:1987:M

[14] Anonymous. MEETINGS. Computers in Physics, 1(1):93–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4903442.

Borchers:1987:LBC

[15] Robert R. Borchers. Looking back at computers in physics. Computers in Physics, 1(1):94–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903443.

Anonymous:1987:ISC

[16] Anonymous. Iowa simulation center goes from real time to the big time. Computers in Physics, 1(1):96–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4903446.

Anonymous:1987:NEA

[17] Anonymous. NCSA expands academic affiliates program, SDSC installs molecular simulation library, Iowa simulation center goes from real time to the big time, Princeton and concurrent to codevelop Navier–Stokes supercomputer. Computers in Physics, 1(1):96–??, November 1987. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4903444.

Anonymous:1987:PCC

[18] Anonymous. Princeton and concurrent to codevelop Navier–Stokes supercomputer. Computers in Physics, 1(1):96–??, November 1987. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4903447.

Anonymous:1987:SIM

[19] Anonymous. SDSC installs molecular simulation library. Computers in Physics, 1(1):96–??, November 1987. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4903445.

Anonymous:1988:JPF

[20] Anonymous. JPL physicists to fabricate VLSI neural network chip by end of ’88. Computers in Physics, 2(1):7–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822652.

Hecht:1988:CMC

[21] Jeff Hecht. Climate models: Computing the world around us. Computers in Physics, 2(1):16–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822646.

Leaf:1988:DTC

[22] Jesse J. Leaf. Databases turn computers into science libraries. Computers in Physics, 2(1):24–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822647.

Centrella:1988:UCC

[23] Joan M. Centrella. Using the computer as a camera. Computers in Physics, 2(1):34–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822648.

Schwartz:1988:ESP

[24] Tom Schwartz. Expert systems prove adept at physics. Computers in Physics, 2(1):40–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822649.

Tinker:1988:CAS

[25] Robert F. Tinker. Computer-aided student investigations. Computers in Physics, 2(1):46–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822650.

Dazzo:1988:HML

[26] Genevieve M. Dazzo. How to maintain large scientific programs. Computers in Physics, 2(1):52–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822651.

Forinash:1988:TPM

[27] Kyle Forinash and Jeff Irick. Theoretical physics made easy. Computers in Physics, 2(1):59–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168295.

Gillman:1988:ARG

[28] E. Gillman and H. R. Fiebig. Accurate recursive generation of spherical Bessel and Neumann functions for a large range of indices. Computers in Physics, 2(1):62–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168296.

Mathes:1988:WBSa

[29] Leslie E. Mathes. Where to buy scientific software for the Macintosh. Computers in Physics, 2(1):75–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822653.

Anonymous:1988:CPP

[30] Anonymous. Celerity packs power in departmental supercomputer. Computers in Physics, 2(1):84–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822654.

Anonymous:1988:MBM

[31] Anonymous. M64 boosts memory capacity. Computers in Physics, 2(1): 84–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822655.

Anonymous:1988:CBB

[32] Anonymous. Coprocessor board boosts TMS32020. Computers in Physics, 2(1):84–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822656.

Anonymous:1988:IBD

[33] Anonymous. Interface board does digital I/O with IBM PS/2. Computers in Physics, 2(1):84–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822657.

Anonymous:1988:GBB

[34] Anonymous. Graphics board brings high resolution to Mac II. Computers in Physics, 2(1):85–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822658.

Anonymous:1988:DCA

[35] Anonymous. A/D converters aid DSP applications. Computers in Physics, 2(1):85–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822659.

Anonymous:1988:HDA

[36] Anonymous. Hard disk adds 45 MB to IBM PS/2. Computers in Physics, 2(1):85–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822660.

Anonymous:1988:CBV

[37] Anonymous. Controller brings video cameras into 3rd dimension. Computers in Physics, 2(1):85–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822661.

Anonymous:1988:PEF

[38] Anonymous. Prolog environment features incremental compiler. Computers in Physics, 2(1):86–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822662.

Anonymous:1988:SAL

[39] Anonymous. Software acquires laboratory data. Computers in Physics, 2(1):86–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822663.

Anonymous:1988:SEP

[40] Anonymous. Spreadsheet excels on PCs. Computers in Physics, 2(1): 86–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822664.

Anonymous:1988:MPP

[41] Anonymous. Math pack plots 10 functions. Computers in Physics, 2(1): 87–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822665.

Anonymous:1988:CET

[42] Anonymous. CAD expands to third dimension. Computers in Physics, 2 (1):87–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822666.

Hawley:1988:BRS

[43] John F. Hawley. Book review: Sidney Karin and Norris Parker Smith, The Supercomputer Era. Computers in Physics, 2(1):89–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822667.

Gould:1988:ICS

[44] Harvey Gould, Jan Tobochnik, and Don E. Harrison, Jr. An introduction to computer simulation methods: Applications to physical systems, part 1 and part 2. Computers in Physics, 2(1):90–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822668.

Crandall:1988:BRM

[45] R. E. Crandall. Book review: The Media Lab: Inventing the Future at MIT, by Stewart Brand. Computers in Physics, 2(1):91–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822669.

Borchers:1988:PVS

[46] Robert R. Borchers. In praise of visualization and simulation. Computers in Physics, 2(1):94–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822670.

Anonymous:1988:Ma

[47] Anonymous. Meetings. Computers in Physics, 2(1):95–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822671.

Anonymous:1988:VIU

[48] Anonymous. Voyager II Uranus images captured on CD ROM. Computers in Physics, 2(1):96–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822672.

Anonymous:1988:TPB

[49] Anonymous. Teachers plan to bring computational physics into undergrad classrooms. Computers in Physics, 2(1):96–??, January 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822673.

Anonymous:1988:SST

[50] Anonymous. Software and supercomputers tutor Carnegie-Mellon physics students. Computers in Physics, 2(1):96–??, January 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822674.

Anonymous:1988:FDS

[51] Anonymous. Fluid dynamics simulations up and running on ETA 10 and Ametek. Computers in Physics, 2(1):96–??, January 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822675.

Batacan:1988:CPM

[52] Peter Batacan. Can physics make optics compute? Computers in Physics, 2(2):9–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822703.

Batacan:1988:MLC

[53] Peter Batacan. MIT media lab close to perfecting computer-generated ‘desktop’ holograms. Computers in Physics, 2(2):17–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822676.

Wherrett:1988:MFO

[54] Brian S. Wherrett. The many facets of optical computing. Computers in Physics, 2(2):24–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822677.

Caulfield:1988:CGH

[55] H. John Caulfield. Computer-generated holograms arrive. Computers in Physics, 2(2):29–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822678.

Jenkins:1988:NWJ

[56] Richard A. Jenkins. The next wave of Japanese supercomputers. Computers in Physics, 2(2):34–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822679.

Gleick:1988:DSC

[57] James Gleick. The dynamical systems collective. Computers in Physics, 2 (2):40–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822680.

Palmer:1988:FCS

[58] Byron Palmer. FORTH: The choice for scientific instruments. Computers in Physics, 2(2):54–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822681.

Arndt:1988:PSA

[59] Richard A. Arndt and L. David Roper. PC software adds color to physics lectures. Computers in Physics, 2(2):62–??, March 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822682.

Haus:1988:CMI

[60] Joseph W. Haus and Tuyen K. Tran. Computational method for intrinsic optical bistability: Steady-state solutions and their stability. Computers in Physics, 2(2):69–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168297.

Mathes:1988:WBC

[61] Leslie E. Mathes. Where to buy computer-controlled instrumentation for PC and Mac. Computers in Physics, 2(2):75–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822683.

Anonymous:1988:SIB

[62] Anonymous. STP-2M interface board. Computers in Physics, 2(2):84–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822684.

Anonymous:1988:PID

[63] Anonymous. Programmable indexer/driver model 86MM. Computers in Physics, 2(2):84–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822685.

Anonymous:1988:MES

[64] Anonymous. Memory enhancement for SCS supercomputers. Computers in Physics, 2(2):84–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822686.

Anonymous:1988:FGE

[65] Anonymous. Frame grabber eases color image processing. Computers in Physics, 2(2):84–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822687.

Anonymous:1988:TOE

[66] Anonymous. Three-in-one expansion board. Computers in Physics, 2(2): 85–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822688.

Anonymous:1988:MBA

[67] Anonymous. Memory boards for Apollo Workstations. Computers in Physics, 2(2):85–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822689.

Anonymous:1988:BSL

[68] Anonymous. BOS server links VAX users to PCs. Computers in Physics, 2(2):85–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822690.

Anonymous:1988:LSF

[69] Anonymous. LabVIEW software features open architecture. Computers in Physics, 2(2):86–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822691.

Anonymous:1988:SPL

[70] Anonymous. Scientific program library is transportable. Computers in Physics, 2(2):86–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822692.

Anonymous:1988:DPL

[71] Anonymous. Development package for large programming projects. Computers in Physics, 2(2):86–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822693.

Anonymous:1988:TMS

[72] Anonymous. Two math/science packages for micros. Computers in Physics, 2(2):86–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822694.

Anonymous:1988:SCP

[73] Anonymous. Scientific calculator package. Computers in Physics, 2(2): 87–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822695.

Anonymous:1988:DRP

[74] Anonymous. Data retrieval and plotting package is Lotus-compatible. Computers in Physics, 2(2):87–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822696.

Anonymous:1988:DAP

[75] Anonymous. Data acquisition and process control software automates entire cycles. Computers in Physics, 2(2):88–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822697.

Anonymous:1988:GWI

[76] Anonymous. Graphics worksheet for IBM PC-compatibles. Computers in Physics, 2(2):88–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822698.

Anonymous:1988:SSF

[77] Anonymous. Software supports four types of digital filtering. Computers in Physics, 2(2):88–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822699.

Anonymous:1988:LMS

[78] Anonymous. Laboratory management system controls access to instruments. Computers in Physics, 2(2):89–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822700.

Anonymous:1988:LSH

[79] Anonymous. Laboratory system handles 100 lines of instructions. Computers in Physics, 2(2):89–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822701.

Anonymous:1988:PQG

[80] Anonymous. Publication-quality graphics package. Computers in Physics, 2(2):89–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822702.

Titus:1988:AD

[81] Jonathan A. Titus. Additional data. Computers in Physics, 2(2):92–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822704.

Duax:1988:CD

[82] William L. Duax. Corrected data. Computers in Physics, 2(2):92–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822705.

Morrison:1988:RD

[83] Michael A. Morrison. Requested data. Computers in Physics, 2(2): 92–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822706.

Anonymous:1988:Mb

[84] Anonymous. Meetings. Computers in Physics, 2(2):93–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822707.

Fernbach:1988:SPP

[85] Sidney Fernbach. In search of peak performance. Computers in Physics, 2 (2):94–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822708.

Anonymous:1988:OFI

[86] Anonymous. Ohio facility to install Cray Y-MP by mid-1989. Computers in Physics, 2(2):96–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822709.

Anonymous:1988:HSP

[87] Anonymous. House Subcommittee planning NSF supercomputing hearings. Computers in Physics, 2(2):96–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822710.

Anonymous:1988:SAS

[88] Anonymous. SuperQuest awarding supercomputer to high school scientists. Computers in Physics, 2(2):96–??, March 1988. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822711.

Anonymous:1988:ASS

[89] Anonymous. Alliant to ship second-generation minisupercomputers. Computers in Physics, 2(2):96–??, March 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822712.

Batacan:1988:WSS

[90] Peter Batacan. Will SISAL spell goodbye for parallel-processor programming woes? Computers in Physics, 2(3):9–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822728.

Anonymous:1988:GSC

[91] Anonymous. Geophysicists to seek computer-generated seismic maps in ’88. Computers in Physics, 2(3):12–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822719.

Anonymous:1988:NUB

[92] Anonymous. NSFNET upgrade boosting scientific computer network to 45 mbits/s. Computers in Physics, 2(3):14–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822720.

Anonymous:1988:VR

[93] Anonymous. The visualization roundtable. Computers in Physics, 2(3): 16–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822721.

Wolff:1988:VES

[94] Robert S. Wolff. Visualization in the eye of the scientist. Computers in Physics, 2(3):28–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822722.

Misner:1988:STP

[95] Charles W. Misner. Spreadsheets tackle physics problems. Computers in Physics, 2(3):37–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822723.

Frenz:1988:CCJ

[96] Bert Frenz. Computers and crystallography: Joint progress. Computers in Physics, 2(3):42–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822724.

Bernstein:1988:SPS

[97] Robert L. Bernstein and Gregory J. Chaitin. Single-particle software. Computers in Physics, 2(3):49–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822725.

Lane:1988:AIS

[98] Eric T. Lane. Apple II software demonstrates wave behavior. Computers in Physics, 2(3):53–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822726.

Lightbody:1988:MSA

[99] J. W. Lightbody, Jr. and J. S. O’Connell. Modeling single arm electron scattering and nucleon production from nuclei by GeV electrons. Computers in Physics, 2(3):57–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168298.

Resler:1988:NSM

[100] D. A. Resler and S. M. Grimes. The nuclear shell model code CRUNCHER. Computers in Physics, 2(3):65–??, May/June 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168299.

Dory:1988:SP

[101] R. A. Dory. Spreadsheets for physics. Computers in Physics, 2(3):70– ??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168300.

Pickover:1988:NRC

[102] Clifford A. Pickover. A note on rendering chaotic “repeller distance towers”. Computers in Physics, 2(3):75–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168301.

Mathes:1988:WBSb

[103] Leslie E. Mathes. Where to buy statistical analysis software packages. Computers in Physics, 2(3):79–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822727.

Anonymous:1988:DRA

[104] Anonymous. DORE renders Ardent graphics. Computers in Physics, 2 (3):93–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822729.

Anonymous:1988:GPS

[105] Anonymous. Graph plots science charts. Computers in Physics, 2(3):94– ??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822730.

Anonymous:1988:RGD

[106] Anonymous. 4Sight runs graphics à deux. Computers in Physics, 2(3):94– ??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822731.

Anonymous:1988:WSS

[107] Anonymous. Waves and sound for students. Computers in Physics, 2(3): 94–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822732.

Anonymous:1988:PPD

[108] Anonymous. PC program digitizes maps. Computers in Physics, 2(3):95– ??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822733.

Anonymous:1988:ISS

[109] Anonymous. Imagine simulating that system on the Mac. Computers in Physics, 2(3):95–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822734.

Anonymous:1988:MIL

[110] Anonymous. Masscomp integrates lab line. Computers in Physics, 2(3): 96–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822735.

Anonymous:1988:PCS

[111] Anonymous. PC-controlled scanning tunneling microscope. Computers in Physics, 2(3):96–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822737.

Anonymous:1988:ESM

[112] Anonymous. Eclipse squeezes minicomputer into 7in lab rack. Computers in Physics, 2(3):96–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822738.

Anonymous:1988:MMA

[113] Anonymous. 1MB Mac memory add-on. Computers in Physics, 2(3):97– ??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822739.

Anonymous:1988:BSP

[114] Anonymous. 32-bit signal processor boasts high speed. Computers in Physics, 2(3):97–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822740.

Anonymous:1988:SSSa

[115] Anonymous. Smart SCRs squelch 5–15V voltage spikes. Computers in Physics, 2(3):97–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822741.

Anonymous:1988:MHL

[116] Anonymous. Mux handles 64 lab inputs. Computers in Physics, 2(3):97– ??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822742.

Anonymous:1988:Mc

[117] Anonymous. Meetings. Computers in Physics, 2(3):98–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822743.

Cooney:1988:BRW

[118] Patrick J. Cooney. Book review: William J. Orvis, 1-2-3 For Scientists And Engineers. Computers in Physics, 2(3):99–??, May/June 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822744.

Cooney:1988:BRO

[119] Patrick J. Cooney. Book review: Oleg D. Jefimenko, Scientific Graphics With Lotus 1-2-3. Computers in Physics, 2(3):99–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822745.

Salpeter:1988:CBW

[120] Edwin E. Salpeter, David Bellin, and Gary Chapman. Computers in battle: Will they work? Computers in Physics, 2(3):100–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822713.

Crandall:1988:NSR

[121] Richard E. Crandall. Notes on simulation and reality [editorial]. Computers in Physics, 2(3):102–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822714.

Anonymous:1988:PPS

[122] Anonymous. Physicists plan supercollider simulations meeting. Computers in Physics, 2(3):104–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822715.

Anonymous:1988:CCC

[123] Anonymous. Cray and Convex compete for computing crown. Computers in Physics, 2(3):104–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822716.

Anonymous:1988:SMM

[124] Anonymous. SDSC molecular modeling teleconference on way to video stores. Computers in Physics, 2(3):104–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822717.

Anonymous:1988:MRD

[125] Anonymous. Minnesota researchers design systolic array building blocks. Computers in Physics, 2(3):104–??, May/June 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822718.

Batacan:1988:AIM

[126] Peter Batacan. AT&T and IBM multiprocessors spur molecular dynamics research. Computers in Physics, 2(4):9–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822757.

Anonymous:1988:SCF

[127] Anonymous. Supercomputer centers funding may increase by $20 million in FY1989. Computers in Physics, 2(4):10–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822746.

Anonymous:1988:CPR

[128] Anonymous. CEBAF physicists ready a state-of-the-art data acquisition system. Computers in Physics, 2(4):12–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822747.

Thompson:1988:ICP

[129] William J. Thompson. Introducing computation to physics students. Computers in Physics, 2(4):14–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822748.

MacDonald:1988:MUP

[130] William M. MacDonald, Edward F. Redish, and Jack M. Wilson. The M.U.P.P.E.T. manifesto. Computers in Physics, 2(4):23–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822749.

Layton:1988:TPH

[131] William Layton and Ralph Roskies. Teaching physicists how to program. Computers in Physics, 2(4):32–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822750.

Chen:1988:CSM

[132] Katherine T. Chen. Computers spur magnetic confinement fusion. Computers in Physics, 2(4):38–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822751.

Weber:1988:CAI

[133] Stephen V. Weber, Frederick W. Holloway, and Lamar W. Coleman. Computers aid inertial confinement fusion research. Computers in Physics, 2(4):46–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822752.

Wampler:1988:NSH

[134] K. Dean Wampler and Lawrence Wilets. On the numerical solution of the Hill–Wheeler equation. Computers in Physics, 2(4):53–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168302.

Miller:1988:OLS

[135] Benjamin P. Miller and Howard E. Dunn. Orthogonal least-squares line fit with variable scaling. Computers in Physics, 2(4):59–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168303.

Mirin:1988:PME

[136] A. A. Mirin. Predicting multiprocessing efficiency on the Cray multiprocessors in a (CTSS) time-sharing environment/application to a 3-D magnetohydrodynamics code. Computers in Physics, 2(4):62–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168304.

Peters:1988:ECP

[137] R. D. Peters. Experimental computational physics using an inexpensive microcomputer. Computers in Physics, 2(4):68–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168305.

Press:1988:KST

[138] William H. Press and Saul A. Teukolsky. Kolmogorov–Smirnov test for two-dimensional data. Computers in Physics, 2(4):74–??, July 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822753.

Leming:1988:CPC

[139] Charles Leming and Joseph K. Daugherty. Computer problems for classical dynamics: an integrated approach. Computers in Physics, 2(4):80–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822754.

Matsen:1988:SAA

[140] F. A. Matsen, T. Tajima, and Roger Haydock. Supercomputers: Algorithms, architectures, and scientific computation. Computers in Physics, 2(4):81–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822755.

Mathes:1988:WBM

[141] Leslie E. Mathes. Where to buy mathematical libraries. Computers in Physics, 2(4):83–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822756.

Anonymous:1988:LMA

[142] Anonymous. Laser microscope alternates scans. Computers in Physics, 2(4):90–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822758.

Anonymous:1988:JFP

[143] Anonymous. JAVA fits the picture. Computers in Physics, 2(4):91–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822759.

Anonymous:1988:Md

[144] Anonymous. Meetings. Computers in Physics, 2(4):92–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822760.

Risley:1988:TPC

[145] John S. Risley. Teaching physics with computers: Does it work? Computers in Physics, 2(4):94–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822761.

Anonymous:1988:DPS

[146] Anonymous. DOE plans superconductivity center at national labs. Computers in Physics, 2(4):96–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822762.

Anonymous:1988:ESP

[147] Anonymous. Evans & Sutherland and Paralex premiere parallel processors. Computers in Physics, 2(4):96–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822763.

Anonymous:1988:PIL

[148] Anonymous. Petroleum institutes lead drive for computer-aided geophysics. Computers in Physics, 2(4):96–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822764.

Anonymous:1988:MUR

[149] Anonymous. Motorola unveils 88000 RISC microprocessor line. Computers in Physics, 2(4):96–??, July 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822765.

Batacan:1988:ALF

[150] Peter Batacan. Argonne launches “flagship” parallel processing center for physics codes. Computers in Physics, 2(5):9–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822778.

Lancaster:1988:CPM

[151] Jack L. Lancaster and Gary D. Fullerton. Computers, physics, and medicine: Imaging the body. Computers in Physics, 2(5):16–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822772.

Jones:1988:CMR

[152] Robert D. Jones and James R. MacFall. Computers in magnetic resonance imaging. Computers in Physics, 2(5):25–??, September 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822773.

Levy:1988:CMR

[153] George C. Levy. Computers in magnetic resonance spectroscopy. Computers in Physics, 2(5):32–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822774.

Vannier:1988:CCA

[154] Michael W. Vannier. Computers in computer axial tomography. Computers in Physics, 2(5):39–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822775.

McKinstry:1988:DAI

[155] David M. McKinstry. 3-D Amiga images for the physics classroom. Computers in Physics, 2(5):44–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822776.

Matzner:1988:ICS

[156] Richard A. Matzner. Interaction of U(1) cosmic strings: Numerical intercommutation. Computers in Physics, 2(5):51–??, September 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168306.

Welch:1988:PSE

[157] L. C. Welch. Programming style: an example. Computers in Physics, 2 (5):65–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168307.

Good:1988:SP

[158] R. H. Good. Simulation programs. Computers in Physics, 2(5):76–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168308.

Delinger:1988:UHH

[159] W. G. Delinger. Use of a hand-held computer to calculate Debye temperatures of cubic crystals from elastic constants. Computers in Physics, 2(5):81–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168309.

Press:1988:ECF

[160] William H. Press and Saul A. Teukolsky. Evaluating continued fractions and computing exponential integrals. Computers in Physics, 2(5):88–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822777.

Greengrove:1988:ME

[161] Cheryl Greengrove. Most enjoyable. Computers in Physics, 2(5):90–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822779.

Lightbody:1988:IT

[162] John W. Lightbody, Jr. Informative, too. Computers in Physics, 2(5):90– ??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822780.

Washington:1988:IC

[163] Warren M. Washington. Initial conditions. Computers in Physics, 2(5): 90–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822781.

Mathes:1988:WBH

[164] Leslie E. Mathes. Where to buy high resolution color monitors. Computers in Physics, 2(5):91–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822782.

Anonymous:1988:MS

[165] Anonymous. Mathematica is the solution. Computers in Physics, 2(5): 97–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822783.

Anonymous:1988:SPN

[166] Anonymous. Software processes NMR data. Computers in Physics, 2(5): 98–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822784.

Anonymous:1988:FCM

[167] Anonymous. Fits curves with MINSQ. Computers in Physics, 2(5):98– ??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822785.

Anonymous:1988:MDS

[168] Anonymous. Menus drive SEGS engineering graphics package. Computers in Physics, 2(5):98–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822786.

Anonymous:1988:SPA

[169] Anonymous. Signal processor adds to multichannel analyzer. Computers in Physics, 2(5):99–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822787.

Anonymous:1988:KAW

[170] Anonymous. Kontron aims workstation at image processing. Computers in Physics, 2(5):99–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822788.

Bates:1988:IRC

[171] R. H. T. Bates, M. J. McDonnell, Wayne Christiansen, and David De-Graff. Image restoration and construction. Computers in Physics, 2(5): 100–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822766.

Ragan:1988:MPT

[172] Don Ragan. Medical physics: Turning science fiction into fact. Computers in Physics, 2(5):102–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822767.

Anonymous:1988:MSO

[173] Anonymous. Materials science online database to debut. Computers in Physics, 2(5):104–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822768.

Anonymous:1988:UCP

[174] Anonymous. University of Colorado pitches parallel processing center to NSF. Computers in Physics, 2(5):104–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822769.

Anonymous:1988:SSSb

[175] Anonymous. Sandia SUPERNET software yields supercomputer speeds. Computers in Physics, 2(5):104–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822770.

Anonymous:1988:MDMa

[176] Anonymous. MCNC duilds massively parallel processor on a chip. Computers in Physics, 2(5):104–??, September 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822771.

Bacon:1988:PSC

[177] Ben Bacon. Pittsburgh Supercomputing Center to get gallium arsenide Cray 3 in 1990. Computers in Physics, 2(6):9–??, November 1988. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822801.

Brecque:1988:RPT

[178] Mort La Brecque. Retinex: Physics and the theory of color vision. Computers in Physics, 2(6):16–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822794.

Aref:1988:CPM

[179] Hassan Aref, Scott W. Jones, and Oran M. Thomas. Computing particle motions in fluid flows. Computers in Physics, 2(6):22–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822795.

Cooper:1988:BRT

[180] Leon N. Cooper. Brain research: Theory and experiment. Computers in Physics, 2(6):29–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822796.

Malaney:1988:SNA

[181] Robert A. Malaney. Supercomputers and nuclear astrophysics. Computers in Physics, 2(6):40–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822797.

Dreher:1988:UMT

[182] Patrick Dreher. Using microcomputers for teaching lattice gauge theory computational techniques. Computers in Physics, 2(6):49–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168310.

Luban:1988:EMG

[183] Marshall Luban and Lawrence P. Staunton. An efficient method for generating a uniform distribution of points within a hypersphere. Computers in Physics, 2(6):55–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168311.

Borie:1988:PCC

[184] E. Borie and B. Jödicke. PBAR: a code for calculating the electromagnetic contribution to the energy levels of antiprotonic atoms. Computers in Physics, 2(6):61–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168312.

Kirkup:1988:CSN

[185] L. Kirkup and J. Sutherland. Curve stripping and nonlinear fitting of polyexponential functions to data using a microcomputer. Computers in Physics, 2(6):64–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168313.

Majumdar:1988:FDC

[186] Sharmila Majumdar and Rahul R. Prasad. The fractal dimension of cerebral surfaces using magnetic resonance images. Computers in Physics, 2 (6):69–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168314.

Siegman:1988:TDC

[187] Anthony E. Siegman. Two-dimensional calculations using onedimensional arrays, or “life on the skew”. Computers in Physics, 2(6): 74–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168315.

Press:1988:SAW

[188] William H. Press and Saul A. Teukolsky. Search algorithm for weak periodic signals in unevenly spaced data. Computers in Physics, 2(6): 77–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822798.

Dory:1988:FAU

[189] Robert A. Dory and Jeffrey H. Harris. Fourier analysis using a spreadsheet. Computers in Physics, 2(6):83–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822799.

Mathes:1988:WBE

[190] Leslie E. Mathes. Where to buy electricity and magnetism software. Computers in Physics, 2(6):89–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822800.

Anonymous:1988:MDMb

[191] Anonymous. MathStation does models and analysis. Computers in Physics, 2(6):96–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822802.

Anonymous:1988:MIA

[192] Anonymous. MATLAB integrates analysis on Mac. Computers in Physics, 2(6):96–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822803.

Anonymous:1988:SMT

[193] Anonymous. Setting math type exactly. Computers in Physics, 2(6):96– ??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822804.

Anonymous:1988:SPD

[194] Anonymous. Stat package does ANOVA. Computers in Physics, 2(6): 97–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822805.

Anonymous:1988:TCG

[195] Anonymous. Titan computer gets molecular simulator. Computers in Physics, 2(6):97–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822806.

Anonymous:1988:PAD

[196] Anonymous. I/O processor acquires data for PCs. Computers in Physics, 2(6):98–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822807.

Anonymous:1988:BSI

[197] Anonymous. Board supplies IEEE-488 interface. Computers in Physics, 2(6):98–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822808.

Anonymous:1988:AAV

[198] Anonymous. Aptec adds vector processor. Computers in Physics, 2(6): 98–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822809.

Haken:1988:CSN

[199] H. Haken and Arthur Lee. Computational systems-natural and artificial. Computers in Physics, 2(6):99–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822810.

Anonymous:1988:CNN

[200] Anonymous. A conference on neural networks. Computers in Physics, 2 (6):102–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822789.

Anonymous:1988:DBF

[201] Anonymous. DARPA to build fiber network. Computers in Physics, 2 (6):104–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822790.

Anonymous:1988:NEA

[202] Anonymous. NSF establishes advanced electronic materials processing center. Computers in Physics, 2(6):104–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822791.

Anonymous:1988:GGS

[203] Anonymous. GaAs/GaAlAs superlattice promises better thermal imaging. Computers in Physics, 2(6):104–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822792.

Anonymous:1988:NBN

[204] Anonymous. NYSERNet boosts network data transfers to 1.5 megabit/s. Computers in Physics, 2(6):104–??, November 1988. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822793.

Bacon:1989:IPD

[205] Ben Bacon. IBM physicist developing software to generate 3-D atomiclevel holograms. Computers in Physics, 3(1):9–??, January 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822817.

Batacan:1989:NCA

[206] Peter Batacan. New computer architectures for lattice QCD. Computers in Physics, 3(1):17–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822811.

Jenkins:1989:NAP

[207] Richard A. Jenkins. New approaches in parallel computing. Computers in Physics, 3(1):24–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822812.

Ewald:1989:PFC

[208] Robert H. Ewald. Perspectives from the field: Cray Research. Computers in Physics, 3(1):33–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822813.

Goodman:1989:SCS

[209] Seymour E. Goodman, William K. McHenry, and Peter Wolcott. Scientific computing in the Soviet Union. Computers in Physics, 3(1):39– ??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822814.

Sweeney:1989:TDQ

[210] Christopher J. Sweeney and Paul L. De Vries. Time-dependent quantummechanical scattering in two dimensions. Computers in Physics, 3(1): 49–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168336.

Inchiosa:1989:PAD

[211] Mario E. Inchiosa and Steven A. Janowsky. A parallel algorithm for determining the ground state of some ferromagnetic spin systems. Computers in Physics, 3(1):55–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168337.

Hocking:1989:PFT

[212] W. K. Hocking. Performing Fourier transforms on extremely long data streams. Computers in Physics, 3(1):59–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168338.

Ogilvie:1989:CAM

[213] J. F. Ogilvie. Computer algebra in modern physics. Computers in Physics, 3(1):66–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168339.

Enloe:1989:SCN

[214] C. L. Enloe. Solving coupled, nonlinear differential equations with commercial spreadsheets. Computers in Physics, 3(1):75–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168340.

Zammit:1989:SSB

[215] Ronald E. Zammit and Bahman Taheri. A simple SCSI bus interface for 8 bit data transfers. Computers in Physics, 3(1):77–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168359.

White:1989:BRG

[216] Richard L. White. Book review: G. Fox, M. Johnson, G. Lyzenga, S. Otto, J. Salmon, and D. Walker, Solving Problems On Concurrent Processors Vol. 1: General Techniques and Regular Problems. Computers in Physics, 3(1):83–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822815.

MacDonald:1989:WBS

[217] Joe MacDonald. Where to buy scientific programming languages. Computers in Physics, 3(1):85–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822816.

Anonymous:1989:CC

[218] Anonymous. CrossCode C. Computers in Physics, 3(1):90–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822818.

Anonymous:1989:CNC

[219] Anonymous. C native compiler. Computers in Physics, 3(1):90–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822819.

Anonymous:1989:SC

[220] Anonymous. Scientific calculator. Computers in Physics, 3(1):90– ??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822820.

Anonymous:1989:MM

[221] Anonymous. Mathematical models. Computers in Physics, 3(1):90– ??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822821.

Anonymous:1989:PG

[222] Anonymous. Presentation graphics. Computers in Physics, 3(1):90– ??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822822.

Press:1989:CAI

[223] William H. Press and Saul A. Teukolsky. Computing accurate integrals with the FFT. Computers in Physics, 3(1):91–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822823.

Anonymous:1989:NUB

[224] Anonymous. NeXT unveils 68030-based workstation. Computers in Physics, 3(1):96–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822824.

Anonymous:1989:NSF

[225] Anonymous. National Science Foundation to boost computer funding 18%. Computers in Physics, 3(1):96–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822825.

Anonymous:1989:JHW

[226] Anonymous. JPL hypercube will achieve 2 gigaflops in early ’89. Computers in Physics, 3(1):96–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822826.

Anonymous:1989:CBG

[227] Anonymous. Cornell breaks ground for $30 million Theory Center. Computers in Physics, 3(1):96–??, January 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822827.

Bacon:1989:RBR

[228] Ben Bacon. Relativity back in the realm of the senses: The null-strut calculus. Computers in Physics, 3(2):9–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822834.

Anderson:1989:OSI

[229] Dana Z. Anderson. Optical systems that imitate human memory. Computers in Physics, 3(2):18–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822828.

Myron:1989:CST

[230] James R. Myron, Larry R. Lines, and R. Philip Bording. Computers in seismic tomography. Computers in Physics, 3(2):26–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822829.

Anderson:1989:GSA

[231] David V. Anderson, Alan R. Fry, Ralf Gruber, and Alexandre Roy. Gigaflop speed algorithm for the direct solution of large block-tridiagonal systems in 3-D physics applications. Computers in Physics, 3(2):33–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168341.

Visscher:1989:DFM

[232] P. B. Visscher. Discrete formulation of Maxwell equations. Computers in Physics, 3(2):42–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168342.

Morales-Acevedo:1989:IMB

[233] Arturo Morales-Acevedo, Eduardo Herrera, and Heriberto Valencia. Inexpensive microcomputer-based system for solar cell I–V characterization. Computers in Physics, 3(2):46–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168343.

Decyk:1989:SC

[234] Viktor K. Decyk and Joan E. Slottow. Supercomputers in the classroom. Computers in Physics, 3(2):50–??, March 1989. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168316.

Hardt:1989:SIC

[235] S. L. Hardt and G. Chen. Semi-intuitive computations and artificial intelligence: Dimensional analysis as a case study. Computers in Physics, 3 (2):55–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168317.

Taylor:1989:ADS

[236] Clayborne D. Taylor and David P. Nicolas. An adaptive data-smoothing routine. Computers in Physics, 3(2):63–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168358.

Arndt:1989:SSP

[237] Richard A. Arndt and L. David Roper. Scattering in a spherical potential: Motion of complex-plane poles and zeros. Computers in Physics, 3(2): 65–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168318.

Cook:1989:CNP

[238] Norman D. Cook. Computing nuclear properties in the fcc model. Computers in Physics, 3(2):73–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168319.

Beasley:1989:PCE

[239] John D. Beasley and Randall D. Peters. Practical computing for experimental scientists. Computers in Physics, 3(2):81–??, March 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822830.

Metcalf:1989:FE

[240] Michael Metcalf, John Reid, and Jeanne T. Martin. Fortran 8x explained. Computers in Physics, 3(2):82–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822831.

Rybicki:1989:DIS

[241] George B. Rybicki. Dawson’s integral and the sampling theorem. Computers in Physics, 3(2):85–87, March 1989. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822832.

Anonymous:1989:LLRa

[242] Anonymous. Listings of the latest releases. Computers in Physics, 3(2): 89–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822833.

Smith:1989:MDH

[243] Roger Smith. In memoriam: Don E. Harrison, Jr. 1927–1988. Computers in Physics, 3(2):94–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822835.

Anonymous:1989:APP

[244] Anonymous. Automatic parallel processing from Cray. Computers in Physics, 3(2):96–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822836.

Anonymous:1989:KNH

[245] Anonymous. Kalos named to head theory center. Computers in Physics, 3(2):96–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822837.

Anonymous:1989:SPS

[246] Anonymous. Space plane simulation at Los Alamos. Computers in Physics, 3(2):96–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822838.

Anonymous:1989:JSB

[247] Anonymous. JPL software brings eyesight to the blind. Computers in Physics, 3(2):96–??, March 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822839.

Ogilvie:1989:ND

[248] J. F. Ogilvie. Notable developments. Computers in Physics, 3(3):9–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822848.

Bacon:1989:USC

[249] Ben Bacon. Using small computers to control big experiments: MTL and CEBAF. Computers in Physics, 3(3):11–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822840.

Anonymous:1989:NNS

[250] Anonymous. Neural network scheduling for Hubble space telescope. Computers in Physics, 3(3):14–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822841.

Carlson:1989:TSC

[251] Edward H. Carlson. Three software construction methods. Computers in Physics, 3(3):16–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822842.

Campbell:1989:AIS

[252] William J. Campbell, Nicholas M. Short, Jr., and Lloyd A. Treinish. Adding intelligence to scientific data management. Computers in Physics, 3(3):26–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822843.

Corey:1989:OPD

[253] Ellen M. Corey and David A. Young. Optimization of physical data tables by simulated annealing. Computers in Physics, 3(3):33–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168320.

Feibig:1989:DWS

[254] Peg L. Feibig, Leslie I. Brown, and Delores M. Etter. Digital waveform sampling rate converter. Computers in Physics, 3(3):38–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168321.

Branham:1989:PTO

[255] Richard L. Branham, Jr. A program for total (orthogonal) least squares in compact storage mode. Computers in Physics, 3(3):42–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168322.

Sawhney:1989:PBD

[256] K. J. S. Sawhney, I. K. Kaul, R. Koul, C. L. Bhat, H. S. Rawat, and S. K. Kaul. PC-based data acquisition system for the Gulmarg gammaray telescope. Computers in Physics, 3(3):47–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168323.

Forinash:1989:CIB

[257] Kyle Forinash and Ray Wisman. Constructive interaction between computer science and physics programs. Computers in Physics, 3(3):50–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168324.

FroeseFischer:1989:SAC

[258] Charlotte Froese Fischer and Muhammad Idrees. Spline algorithms for continuum functions. Computers in Physics, 3(3):53–??, May 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168325.

Black:1989:GRN

[259] S. C. Black and A. D. Kennedy. Gaussian random number generators on a CYBER-205. Computers in Physics, 3(3):59–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168326.

Johnson:1989:APP

[260] T. K. Johnson and R. L. Vessella. On the application of parallel processing to the computation of dose arising from the internal deposition of radionuclides. Computers in Physics, 3(3):69–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168327.

Boyd:1989:MCS

[261] Iain D. Boyd. Monte Carlo simulation of an expanding gas. Computers in Physics, 3(3):73–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168328.

Schmid:1989:TPP

[262] Erich W. Schmid, Gerhard Spitz, Wolfgang Losch, A. H. Armstrong, and Roger P. Kohin. Theoretical physics on the personal computer. Computers in Physics, 3(3):81–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822844.

Stauffer:1989:CSC

[263] D. Stauffer, F. W. Hehl, V. Winkelmann, J. G. Zabolitzky, and W. J. Thompson. Computer simulation and computer algebra: Lectures for beginners. Computers in Physics, 3(3):83–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822845.

Dory:1989:MCP

[264] Robert A. Dory. Matrices and contour plots. Computers in Physics, 3 (3):84–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822846.

Press:1989:ISO

[265] William H. Press and Saul A. Teukolsky. Integrating stiff ordinary differential equations. Computers in Physics, 3(3):88–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822847.

Anonymous:1989:LLRb

[266] Anonymous. Listings of the latest releases. Computers in Physics, 3 (3):92–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822849.

Borchers:1989:QR

[267] Robert R. Borchers. A quieter revolution. Computers in Physics, 3(3): 96–??, May 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822850.

Bacon:1989:NTM

[268] Ben Bacon. NCUBE 2 and Thinking Machines CM-2a aimed at university market. Computers in Physics, 3(4):9–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822860.

Middleditch:1989:CDP

[269] John Middleditch. Computer detection of pulsar 1987A. Computers in Physics, 3(4):14–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822855.

Middleditch:1989:AFS

[270] John Middleditch. Adventures in Fourier space: Hunting the millisecond pulsar. Computers in Physics, 3(4):19–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822856.

Hecht:1989:PLC

[271] Jeff Hecht. Physical limits of computing. Computers in Physics, 3(4): 34–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822857.

Smith:1989:SDI

[272] R. Smith and D. E. Harrison, Jr. Simulation of the dynamics of ionbombarded surfaces on a desktop computer. Computers in Physics, 3 (4):41–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168329.

Dahiya:1989:CIM

[273] J. N. Dahiya and R. W. Freeman. Computer interfacing of a microwave resonant cavity for temperature measurements during dielectric relaxation. Computers in Physics, 3(4):49–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168330.

Treinish:1989:IDI

[274] Lloyd A. Treinish. An interactive, discipline-independent data visualization system. Computers in Physics, 3(4):55–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168331.

Gray:1989:SRM

[275] K. J. Gray, B. G. Peterson, L. V. Knight, and J. M. Thorne. Simulated response of multilayer X-ray reflectors to damaging fluxes. Computers in Physics, 3(4):65–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168332.

Castano:1989:CAD

[276] Victor M. Castaño and Adolfo Bribiesca. Computer-aided design and analysis of electronic circuits for instrumentation and control. Computers in Physics, 3(4):69–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168333.

Meyer:1989:SMF

[277] W. T. Meyer and M. S. Gorbics. A simple microprocessor for FASTBUS slave modules. Computers in Physics, 3(4):73–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168334.

Pickover:1989:MAS

[278] Clifford A. Pickover. Markov aggregation on a sticky circle. Computers in Physics, 3(4):79–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168335.

Gould:1989:OCS

[279] Harvey Gould and Jan Tobochnik. Overcoming critical slowing down. Computers in Physics, 3(4):82–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822858.

Press:1989:CRC

[280] William H. Press and Saul A. Teukolsky. Cyclic redundancy checks for data integrity or identity. Computers in Physics, 3(4):88–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822859.

Dory:1989:WZS

[281] Robert A. Dory. Wing Z spreadsheet program. Computers in Physics, 3 (4):93–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822861.

Feitelson:1989:OC

[282] D. G. Feitelson and David Casasent. Optical computing. Computers in Physics, 3(4):102–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822851.

Mader:1989:NMW

[283] Charles L. Mader and Robert L. Street. Numerical modeling of water waves. Computers in Physics, 3(4):103–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822852.

Anonymous:1989:LLRc

[284] Anonymous. Listings of the latest releases. Computers in Physics, 3(4): 104–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822853.

Clark:1989:VWT

[285] Barry Clark. Viewing the world through a computer. Computers in Physics, 3(4):112–??, July 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822854.

Anonymous:1989:CPE

[286] Anonymous. The Computers in Physics Education Software Awards. Computers in Physics, 3(5):8–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822870.

Bacon:1989:VS

[287] Ben Bacon. Visualization at SIGGRAPH. Computers in Physics, 3(5): 11–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822866.

Rapaport:1989:VP

[288] Dennis Rapaport. Visualizing physics. Computers in Physics, 3(5):18– ??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822868.

Noble:1989:SCF

[289] Julian Noble. Scientific computation in FORTH. Computers in Physics, 3(5):31–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822869.

Butler:1989:VMS

[290] D. M. Butler and M. H. Pendley. The visualization management system approach to visualization in scientific computing. Computers in Physics, 3(5):40–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168344.

Butler:1989:VMB

[291] D. M. Butler and M. H. Pendley. A visualization model based on the mathematics of fiber bundles. Computers in Physics, 3(5):45–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168345.

Kutt:1989:OEM

[292] P. H. Kutt and D. P. Balamuth. Operating experience with a multipleprocessor system for data acquisition and reduction in nuclear physics. Computers in Physics, 3(5):52–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168346.

Vasavada:1989:ALC

[293] Kashyap V. Vasavada and Jack H. Freed. Application of Lanczos and conjugate gradient methods to a class of computational problems in physics. Computers in Physics, 3(5):61–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168347.

Sebhatu:1989:LSM

[294] Mesgun Sebhatu and W. Edward Gettys. A least-squares method for extracting phase shifts. Computers in Physics, 3(5):65–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168348.

Smith:1989:AMD

[295] R. Smith and D. E. Harrison, Jr. Algorithms for molecular dynamics simulations of keV particle bombardment. Computers in Physics, 3(5):68– ??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168349.

Westerveld:1989:PCF

[296] W. B. Westerveld. Plotting of closed field lines with an application to coherently excited atoms. Computers in Physics, 3(5):74–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168350.

Camarda:1989:DCH

[297] H. S. Camarda. Diagonalization of a complex Hermitian matrix with a band structure. Computers in Physics, 3(5):78–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168351.

Press:1989:FTR

[298] William H. Press and Saul A. Teukolsky. Fourier transforms of real data in two and three dimensions. Computers in Physics, 3(5):84–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822871.

Dory:1989:ODE

[299] Robert A. Dory. Ordinary differential equations. Computers in Physics, 3(5):88–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822872.

Cralle:1989:WN

[300] R. K. Cralle. What’s NeXT? Computers in Physics, 3(5):93–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822873.

Ferrenberg:1989:OMC

[301] Alan M. Ferrenberg and Robert H. Swendsen. Optimized Monte Carlo data analysis. Computers in Physics, 3(5):101–??, September 1989. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822862.

Liu:1989:BRR

[302] Stanley Liu. Book review: Ronald L. Graham, Donald E. Knuth, and Oren Patashnik, Concrete Mathematics: a Foundation for Computer Science. Computers in Physics, 3(5):106–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822863.

Vicsek:1989:FGP

[303] Tamás Vicsek and Harvey Gould. Fractal growth phenomena. Computers in Physics, 3(5):108–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822864.

Anonymous:1989:LLRd

[304] Anonymous. Listings of the latest releases. Computers in Physics, 3(5): 109–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822865.

Bacon:1989:LP

[305] Ben Bacon. The last positivist. Computers in Physics, 3(5):112–??, September 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822867.

Kalantar:1989:IV

[306] A. H. Kalantar. Ineffective variance. Computers in Physics, 3(6):10– ??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822874.

Karpf:1989:VSF

[307] Andreas Karpf. Voyager scientists have a field day with Neptune data. Computers in Physics, 3(6):13–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822876.

Phillpot:1989:HDC

[308] Simon R. Phillpot, Sidney Yip, and Dieter Wolf. How do crystals melt? Computers in Physics, 3(6):20–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822877.

Mihalisin:1989:MGT

[309] T. Mihalisin, E. Gawlinski, J. Timlin, and J. Schwegler. Multidimensional graphing in two-dimensional spaces. Computers in Physics, 3(6): 32–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822878.

Rots:1989:SPI

[310] Arnold H. Rots. Special purpose image display devices and techniques at the national radio astronomy observatory. Computers in Physics, 3(6):

41–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168352.

Chan:1989:LSF

[311] F. L. Chan, K. M. Lau, C. L. Li, H. M. Mok, D. J. Newman, and Betty Ng. Least-squares fitting of parametrized Hamiltonians to energy levels: an investigation into alternative strategies. Computers in Physics, 3(6): 47–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168353.

Siedband:1989:BTI

[312] Melvin P. Siedband, Huilian Zhang, and James Gray. Block truncation image compression with bit-map compensation. Computers in Physics, 3(6):52–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168354.

Santucci:1989:SLC

[313] S. Santucci, P. Picozzi, L. Lozzi, and R. Diamanti. A simple and low-cost interface between an apparatus for Auger spectroscopy and a personal computer. Computers in Physics, 3(6):57–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168355.

Dickens:1989:UMC

[314] J. K. Dickens. Use of Monte Carlo techniques to derive yields for n+^{12} C multibody breakup reactions: Programming the computer to simulate collisions by fast neutrons. Computers in Physics, 3(6):62–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168356.

Pickover:1989:NCE

[315] Clifford A. Pickover. A note on computer experiments with chaotic shattering of level sets. Computers in Physics, 3(6):69–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168357.

Press:1989:QSR

[316] William H. Press and Saul A. Teukolsky. Quasi- (that is, sub-) random numbers. Computers in Physics, 3(6):76–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822879.

Syed:1989:UHP

[317] M. Q. Syed and M. A. Hastaoglu. The use of HYPERCARD in physics. Computers in Physics, 3(6):81–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822880.

Tobochnik:1989:QC

[318] Jan Tobochnik and Harvey Gould. Quantifying chaos. Computers in Physics, 3(6):86–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822881.

Cralle:1989:WIG

[319] R. K. Cralle. What if Gauss had a mouse? Computers in Physics, 3(6): 92–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822882.

Thagard:1989:CPS

[320] Paul Thagard and Peter Skiff. Computational philosophy of science. Computers in Physics, 3(6):96–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822883.

Sewell:1989:NSO

[321] Granville Sewell and Harvey Gould. The numerical solution of ordinary and partial differential equations. Computers in Physics, 3(6):98– ??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822884.

Anonymous:1989:LLRe

[322] Anonymous. Listings of the latest releases. Computers in Physics, 3(6): 99–??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822885.

Donnelly:1989:ME

[323] Denis Donnelly. A means to an end. Computers in Physics, 3(6):112– ??, November 1989. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822875.

Karpf:1990:LSM

[324] Andreas Karpf. Large-scale mapping of the Universe with IRAS data. Computers in Physics, 4(1):9–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822892.

Boghosian:1990:CPC

[325] Bruce M. Boghosian. Computational physics on the Connection Machine. Computers in Physics, 4(1):14–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822889.

Petrig:1990:MSH

[326] Benno L. Petrig and Charles E. Riva. Measurement and simulation of the human eye. Computers in Physics, 4(1):36–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822890.

Family:1990:SFA

[327] Fereydoon Family and Tamás Vicsek. Simulating fractal aggregation. Computers in Physics, 4(1):44–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822891.

Ram-Mohan:1990:FEM

[328] L. R. Ram-Mohan, Sunil Saigal, Don Dossa, and J. Shertzer. The finiteelement method for energy eigenvalues of quantum mechanical systems. Computers in Physics, 4(1):50–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168374.

Aguirregabiria:1990:ECP

[329] J. M. Aguirregabiria, A. Hernández, M. Rivas, and M. A. Valle. On the equilibrium configuration of point charges placed on an ellipse. Computers in Physics, 4(1):60–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168375.

Dreitlein:1990:SST

[330] Joseph F. Dreitlein and Jon R. Sauer. Spinor software tools in C++. Computers in Physics, 4(1):64–??, January 1990. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168376.

Chia:1990:MGH

[331] T. T. Chia and Y. K. Lai. Microcomputer-generated holograms — the programming aspects. Computers in Physics, 4(1):73–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168377.

Snider:1990:CAD

[332] A. Snider and T. R. Mackie. Continuous antialiased 2-D and 3-D transformations in discrete space. Computers in Physics, 4(1):78–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168378.

Boyd:1990:CRF

[333] John P. Boyd. A Chebyshev/radiation function pseudospectral method for wave scattering. Computers in Physics, 4(1):83–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168395.

Hertz:1990:AFF

[334] Paul Hertz. An algorithm for the fast Fourier transform on a Connection Machine. Computers in Physics, 4(1):86–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168379.

Press:1990:EI

[335] William H. Press and Saul A. Teukolsky. Elliptic integrals. Computers in Physics, 4(1):92–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822893.

Donnelly:1990:SDE

[336] Denis Donnelly. Solving differential equations using MathCAD. Computers in Physics, 4(1):97–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822894.

Orvis:1990:FDE

[337] William J. Orvis. Fitting data to equations with spreadsheet programs. Computers in Physics, 4(1):99–??, January 1990. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822895.

McLean:1990:ECA

[338] Ian S. McLean and Michal Simon. Electronic and computer-aided astronomy. Computers in Physics, 4(1):104–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822886.

Tietz:1990:IOD

[339] Jim Tietz. Icon-oriented data acquisition. Computers in Physics, 4(1): 106–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822887.

Fox:1990:TS

[340] Geoffrey Fox. The teraflop supercomputer. Computers in Physics, 4(1): 112–??, January 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822888.

Karpf:1990:PSU

[341] Andreas Karpf. Plasma simulations used to design new magnetic fusion devices. Computers in Physics, 4(2):121–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822896.

Crandall:1990:NCP

[342] Richard E. Crandall. The NeXT computer as physics machine. Computers in Physics, 4(2):132–141, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822897.

Greengard:1990:NSB

[343] Leslie Greengard. The numerical solution of the N-body problem. Computers in Physics, 4(2):142–152, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822898.

Smith:1990:APL

[344] Frank A. Smith, Jr. Animating the physics lecture with Amiga BA-SIC. Computers in Physics, 4(2):155–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168396.

Cahill:1990:HUM

[345] Kevin Cahill. How to use MACSYMA to write long FORTRAN codes for noncompact simulations of gauge theories. Computers in Physics, 4(2):159–165, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168380.

Mitchell:1990:CLI

[346] J. Ross Mitchell, Peter Dickof, and Alan G. Law. A comparison of line integral algorithms. Computers in Physics, 4(2):166–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168381.

Yamazaki:1990:SAM

[347] M. Yamazaki and Y. Murai. Simulation of an antiferromagnet by microcomputer. Computers in Physics, 4(2):173–??, March–April 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168360.

Yu:1990:QTP

[348] K. W. Yu. Quantum transmission in periodic potentials: a transfer matrix approach. Computers in Physics, 4(2):176–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168361.

Bishop:1990:ISA

[349] Marvin Bishop, Jordan D. Perel, and Victoria Swan. Investigation of sorting algorithms on a variety of computers. Computers in Physics, 4(2):179–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168362.

Tobochnik:1990:ERW

[350] Jan Tobochnik. Efficient random walk algorithm for computing conductivity in continuum percolation systems. Computers in Physics, 4(2):181– ??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168363.

Farrar:1990:GTS

[351] Kelly A. Farrar and Adrian L. Melott. Gravity in twisted space. Computers in Physics, 4(2):185–??, March–April 1990. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168364.

Press:1990:RSS

[352] William H. Press and Glennys R. Farrar. Recursive stratified sampling for multidimensional Monte Carlo integration. Computers in Physics, 4(2):190–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822899.

Cook:1990:ICT

[353] David M. Cook. Introducing computational tools in the upper-division undergraduate physics curriculum. Computers in Physics, 4(2):197–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822900.

Gould:1990:MFC

[354] Harvey Gould and Jan Tobochnik. More on fractals and chaos: Multifractals. Computers in Physics, 4(2):202–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822901.

Wolff:1990:EPS

[355] Robert S. Wolff. Engineering paradigm shifts. Computers in Physics, 4(2):208–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822902.

Maier:1990:DMA

[356] W. L. Maier. Derive — a mathematical assistant. Computers in Physics, 4(2):210–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822903.

Nosenchuck:1990:BRN

[357] Daniel M. Nosenchuck. Book review: Neural Computing Architectures: The Design of Brain-Like Machines, Igor Aleksander, 320pp., MIT Press, 1989 $45. Computers in Physics, 4(2):212–??, March–April 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822904.

Anonymous:1990:LLRa

[358] Anonymous. Listings of the latest releases. Computers in Physics, 4 (2):213–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822905.

Bacon:1990:DTT

[359] Ben Bacon. In defense of the Turing test. Computers in Physics, 4 (2):216–??, March–April 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822906.

Sowell:1990:NCA

[360] Glenn Sowell. The NeXT Computer: Article or advertisement? Computers in Physics, 4(3):224–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822907.

Hamilton:1990:TT

[361] Russell J. Hamilton. Turing test. Computers in Physics, 4(3):224–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822908.

Wolff:1990:RNC

[362] Robert S. Wolff. Report from NCGA conference. Computers in Physics, 4 (3):225–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822909.

McRae:1990:SSS

[363] Gregory J. McRae and Armistead G. Russell. Smog, supercomputers and society. Computers in Physics, 4(3):227–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822910.

Chervin:1990:HPC

[364] Robert M. Chervin. High performance computing and the grand challenge of climate modeling. Computers in Physics, 4(3):234–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822911.

Washington:1990:CSG

[365] Warren M. Washington and Thomas W. Bettge. Computer simulation of the greenhouse effect. Computers in Physics, 4(3):240–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822912.

Dozier:1990:LAE

[366] Jeff Dozier. Looking ahead to EOS: The Earth observing system. Computers in Physics, 4(3):248–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822913.

Wolff:1990:VTG

[367] Robert S. Wolff. Visualization 101: a tour guide of basic concepts. Computers in Physics, 4(3):260–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822914.

Manner:1990:PMA

[368] R. Männer and O. Stucky. The Polyp multiprocessor: Architecture and applications in nuclear physics. Computers in Physics, 4(3):267–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168365.

Saleh:1990:TDA

[369] A. H. M. Saleh. Time domain analysis of the impulse response of physical systems using optimization methods. Computers in Physics, 4(3):275–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168366.

Calitz:1990:AHP

[370] M. F. Calitz, A. G. du Toit, and W. Drijfhout. Analysis of high-perveance electron guns. Computers in Physics, 4(3):280–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168367.

Esteban:1990:ACN

[371] E. P. Esteban and E. Ramos. Algebraic computing and the Newman– Penrose formalism. Computers in Physics, 4(3):285–??, May 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168368.

Mansuripur:1990:SMR

[372] Masud Mansuripur and Roscoe Giles. Simulation of the magnetizationreversal dynamics on the Connection Machine. Computers in Physics, 4 (3):291–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168369.

Roomany:1990:GTA

[373] Hamzeh H. Roomany. A graph theoretic approach to the brachistochrone problem. Computers in Physics, 4(3):303–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168370.

Cook:1990:CEU

[374] David M. Cook. Computational exercises for the upper-division undergraduate physics curriculum. Computers in Physics, 4(3):308–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822915.

Harris:1990:DW

[375] R. Harris. Demons at work. Computers in Physics, 4(3):314–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822916.

Press:1990:HFD

[376] William H. Press and Saul A. Teukolsky. Hypergeometric functions by direct path integration. Computers in Physics, 4(3):320–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822917.

Dory:1990:NN

[377] Robert A. Dory. Neural networks. Computers in Physics, 4(3):324–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822918.

Brandt:1990:QMP

[378] Siegmund Brandt, Dieter Dahmen, and Roger P. Kohin. Quantum mechanics on the personal computer. Computers in Physics, 4(3):329–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822919.

Anonymous:1990:LLRb

[379] Anonymous. Listings of the latest releases. Computers in Physics, 4(3): 332–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822920.

Borchers:1990:CPO

[380] Robert R. Borchers. Computers, physics and our world. Computers in Physics, 4(3):336–??, May 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822922.

Strugaru:1990:R

[381] Crisan Strugaru. Romania. Computers in Physics, 4(4):344–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822923.

Cothary:1990:BPP

[382] Louise Cothary. Brown paper packages. Computers in Physics, 4(4): 344–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822924.

Bacon:1990:GIJ

[383] Ben Bacon. Government and industry jointly fund gigabit network research. Computers in Physics, 4(4):345–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822925.

Wolff:1990:MSC

[384] Robert S. Wolff. The Macintosh scientific computing environment. Computers in Physics, 4(4):348–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822926.

Morrison:1990:RME

[385] Richard A. Morrison. Remote monitoring of electric appliances with an intelligent monitor. Computers in Physics, 4(4):362–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822927.

Cahalan:1990:CRD

[386] Robert F. Cahalan, Henning Leidecker, and Gabriel D. Cahalan. Chaotic rhythms of a dripping faucet. Computers in Physics, 4(4):368–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822928.

Wampler:1990:OOP

[387] K. Dean Wampler. The object-oriented programming paradigm (OOPP) and FORTRAN programs. Computers in Physics, 4(4):385–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168371.

Garner:1990:NSS

[388] J. Garner. Numerical solution to the Schrödinger equation by constrained optimization. Computers in Physics, 4(4):395–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168372.

Jacobson:1990:FDS

[389] L. A. Jacobson. A fast data smoothing algorithm. Computers in Physics, 4(4):400–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168373.

Lentz:1990:CFC

[390] William J. Lentz. Continued fraction calculation of spherical Bessel functions. Computers in Physics, 4(4):403–407, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168382.

Aronson:1990:CSB

[391] J. W. Aronson. CHAOS: a SUN-based program for analyzing chaotic systems. Computers in Physics, 4(4):408–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168383.

Horowitz:1990:GFL

[392] E. J. Horowitz. A general field-line plotting algorithm. Computers in Physics, 4(4):418–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168384.

Press:1990:OPG

[393] William H. Press and Saul A. Teukolsky. Orthogonal polynomials and Gaussian quadrature with nonclassical weight functions. Computers in Physics, 4(4):423–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822929.

Sears:1990:STM

[394] R. K. Sears, B. G. Orr, and T. M. Sanders, Jr. A scanning tunneling microscope for undergraduate laboratories. Computers in Physics, 4(4): 427–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822930.

Tobochnik:1990:IQM

[395] Jan Tobochnik, Harvey Gould, and Kenneth Mulder. An introduction to quantum Monte Carlo. Computers in Physics, 4(4):431–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822931.

Watson:1990:CCT

[396] Val Watson. The change in critical technologies for computational physics. Computers in Physics, 4(4):436–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822932.

Visscher:1990:FEC

[397] Pieter B. Visscher and Hilliard K. Macomber. Fields and electrodynamics a computer-compatible introduction. Computers in Physics, 4(4):441–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822933.

Anonymous:1990:LLRc

[398] Anonymous. Listings of the latest releases. Computers in Physics, 4(4): 443–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822934.

Bacon:1990:EJ

[399] Ben Bacon. The electronic journal? Computers in Physics, 4(4):448–??, July 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822935.

Coats:1990:FA

[400] C. F. Coats. Friendly Amigas. Computers in Physics, 4(5):456–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822936.

Bacon:1990:MMG

[401] Ben Bacon. Micro to macro grand challenges. Computers in Physics, 4(5):457–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822937.

Pickover:1990:WC

[402] Clifford A. Pickover. The world of chaos. Computers in Physics, 4(5): 460–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822938.

Field:1990:SC

[403] Mike Field and Martin Golubitsky. Symmetric chaos. Computers in Physics, 4(5):470–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822939.

Markus:1990:CMC

[404] Mario Markus. Chaos in maps with continuous and discontinuous maxima. Computers in Physics, 4(5):481–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822940.

Rossler:1990:SSS

[405] O. E. Rössler, M. Klein, G. Baier, J. Parisi, J. Peinke, M. Krug, and J. L. Hudson. Sierpinski sponge in a simple generic diffeomorphism. Computers in Physics, 4(5):494–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822941.

Frame:1990:CSD

[406] Michael Frame and Lynne Erdman. Coloring schemes and the dynamical structure of iterated function systems. Computers in Physics, 4(5):500– ??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822942.

Wells:1990:TDF

[407] N. H. Wells, C. S. Burrus, G. E. Desobry, and A. L. Boyer. Threedimensional Fourier convolution with an array processor. Computers in Physics, 4(5):507–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168385.

Arnold:1990:MPE

[408] Clifford N. Arnold. Methods for performance evaluation of algorithms and computers. Computers in Physics, 4(5):514–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168386.

DeAcetis:1990:UAT

[409] Louis A. DeAcetis, Oron Schmidt, and Kumar Krishen. Using ADA tasks to simulate operating equipment. Computers in Physics, 4(5):521– ??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168387.

Blackburn:1990:MSN

[410] James A. Blackburn and H. J. T. Smith. Modeling superconducting networks containing Josephson junctions by means of PC-based circuit simulation software. Computers in Physics, 4(5):526–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168388.

Ning:1990:MSA

[411] Xiaohui Ning. Modeling saturable absorption using 3-D spreadsheets. Computers in Physics, 4(5):532–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168389.

Wolff:1990:EVW

[412] Robert S. Wolff. Exploratory visualization — the “war” on data glut. Computers in Physics, 4(5):536–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822943.

Donnelly:1990:CFA

[413] Denis Donnelly. CIP’s first annual software contest: The winners. Computers in Physics, 4(5):540–??, September 1990. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822944.

Srivastava:1990:HC

[414] Niraj Srivastava, Charles Kaufman, and Gerhard Müller. Hamiltonian chaos. Computers in Physics, 4(5):549–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822945.

Press:1990:FVI

[415] William H. Press and Saul A. Teukolsky. Fredholm and Volterra integral equations of the second kind. Computers in Physics, 4(5):554–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822946.

Dory:1990:MS

[416] Robert A. Dory. Macros in a spreadsheet. Computers in Physics, 4(5): 558–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822947.

Baker:1990:CDI

[417] G. L. Baker, J. P. Gollub, and Rick Turner. Chaotic dynamics — an introduction. Computers in Physics, 4(5):562–??, September 1990. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822948.

Thompson:1990:NDC

[418] J. M. T. Thompson, H. B. Stewart, and Rick Turner. Nonlinear dynamics and chaos. Computers in Physics, 4(5):562–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822949.

Wiggins:1990:IAN

[419] Stephen Wiggins and David S. Mazel. Introduction to applied nonlinear dynamical systems and chaos. Computers in Physics, 4(5):563–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822950.

Anonymous:1990:LLRd

[420] Anonymous. Listings of the latest releases. Computers in Physics, 4(5): 564–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822951.

Pickover:1990:TFA

[421] Clifford A. Pickover. The twenty-fourth annual meeting of Chaos Society. Computers in Physics, 4(5):568–??, September 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822952.

Anonymous:1990:PCH

[422] Anonymous. Physics Computing ’91 to be held June 10–14 at San Jose Convention Center. Computers in Physics, 4(6):577–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822953.

Banko:1990:MS

[423] Brad Banko. Monitor safety. Computers in Physics, 4(6):581–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822954.

Graves:1990:CAP

[424] David B. Graves and Richard A. Gottscho. Computer applications in plasma materials processing. Computers in Physics, 4(6):584–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822955.

Flanagan:1990:CSC

[425] J. L. Flanagan. Computers in speech communications research. Computers in Physics, 4(6):592–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822956.

Bitz:1990:DVV

[426] François J. Bitz and Norman J. Zabusky. David and “visiometrics”: Visualizing and quantifying evolving amorphous objects. Computers in Physics, 4(6):603–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822957.

Younan:1990:SFR

[427] Nicolas H. Younan and William G. Frazier. A statistic filter routine. Computers in Physics, 4(6):615–??, November 1990. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168390.

Leclerc:1990:SCC

[428] G. Leclerc and L. Sanche. Spreadsheets for computing charged-particle trajectories in 3-D electrostatic fields. Computers in Physics, 4(6):617– ??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168391.

Laeri:1990:SCM

[429] Franco Laeri. Some comments on maximum entropy spectral analysis of time series. Computers in Physics, 4(6):627–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168392.

Sachar:1990:OCS

[430] Kenneth S. Sachar and Barry Goldstein. Optimization of the controlled separation of biologically active diagnostic magnetic probes. Computers in Physics, 4(6):637–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168393.

Welch:1990:ICT

[431] Daniel W. Welch. Imaging by convergent tiling. Computers in Physics, 4 (6):645–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168394.

Wolff:1990:NTM

[432] Robert S. Wolff. Navigating through multi-dimensional data sets: Dynamical user interface issues. Computers in Physics, 4(6):650–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822958.

Horbatsch:1990:TCL

[433] Marko Horbatsch. Teaching calculus and linear algebra with DE-RIVE(R). Computers in Physics, 4(6):656–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822959.

Reynolds:1990:DQM

[434] Peter J. Reynolds, Jan Tobochnik, and Harvey Gould. Diffusion quantum Monte Carlo. Computers in Physics, 4(6):662–??, November 1990.

CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822960.

Press:1990:SGS

[435] William H. Press and Saul A. Teukolsky. Savitzky-Golay smoothing filters. Computers in Physics, 4(6):669–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822961.

Duffy:1990:IGD

[436] Philip B. Duffy. Igor: Graphics and data analysis for the Macintosh. Computers in Physics, 4(6):673–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822962.

Borchers:1990:DAM

[437] Robert R. Borchers. The dog ate the manuscript. Computers in Physics, 4(6):680–??, November 1990. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822963.

Zeyher:1991:CC

[438] Allen Zeyher. Chaos in the classroom. Computers in Physics, 5(1): 9–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822974.

Jong:1991:CIP

[439] Marvin L. De Jong. Computers in introductory physics. Computers in Physics, 5(1):12–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822964.

Taylor:1991:MC

[440] Cyrus Taylor. Mathematica in the classroom. Computers in Physics, 5 (1):16–21, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822966.

Cline:1991:NPD

[441] Richard W. Cline. Networking a physics department: a case history. Computers in Physics, 5(1):22–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822967.

Mazur:1991:CWT

[442] Eric Mazur. Can we teach computers to teach? Computers in Physics, 5 (1):31–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822968.

Whitton:1991:VSU

[443] Mary C. Whitton. Visualization for Sun users. Computers in Physics, 5 (1):40–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822969.

McClain:1991:ACP

[444] Marjorie McClain, Albert Feldman, David Kahaner, and Xuantong Ying. An algorithm and computer program for the calculation of envelope curves. Computers in Physics, 5(1):45–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168405.

Regan:1991:RRO

[445] Jeffery J. Regan and David R. Andersen. Reflection and refraction of optical beams at dielectric interfaces. Computers in Physics, 5(1):49–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168406.

Stauffer:1991:PCA

[446] Dietrich Stauffer. Programming cellular automata. Computers in Physics, 5(1):62–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822970.

Press:1991:NCD

[447] William H. Press and Saul A. Teukolsky. Numerical calculation of derivatives. Computers in Physics, 5(1):68–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822971.

Anonymous:1991:LLRa

[448] Anonymous. Listings of the latest releases. Computers in Physics, 5(1): 70–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822972.

Anonymous:1991:DPC

[449] Anonymous. 1991 directory of physics courseware. Computers in Physics, 5(1):71–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822973.

Bacon:1991:AEQ

[450] Ben Bacon. An answer to Eric’s question. Computers in Physics, 5(1): 128–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822965.

Bernstein:1991:TGG

[451] R. Bernstein. Taking the generality out of general relativity. Computers in Physics, 5(1):463–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168431.

Mahonen:1991:PSP

[452] Petri Mähönen. Parallel signal processing workstation environment. Computers in Physics, 5(1):611–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168418.

Zeyher:1991:SVE

[453] Allen Zeyher. Simulating the very early Universe. Computers in Physics, 5(2):135–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822975.

Norman:1991:CAF

[454] Michael L. Norman, David A. Clarke, and James M. Stone. Computational astrophysical fluid dynamics. Computers in Physics, 5(2):138–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822976.

Max:1991:CSA

[455] Claire E. Max. Computer simulation of astrophysical plasmas. Computers in Physics, 5(2):152–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822977.

Bertschinger:1991:CBS

[456] Edmund Bertschinger and James M. Gelb. Cosmological N-body simulations. Computers in Physics, 5(2):164–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822978.

Smarr:1991:ACB

[457] Larry L. Smarr and William H. Press. Astronomical computing in the 1990s: The Bahcall Committee Report. Computers in Physics, 5(2): 180–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822979.

Alex:1991:MFA

[458] Michael Alex and Jason Crain. Mean-field analysis on the Macintosh. Computers in Physics, 5(2):193–197, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168407.

Creutz:1991:AS

[459] Michael Creutz. Abelian sandpiles. Computers in Physics, 5(2):198–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168408.

Giles:1991:PSC

[460] Roscoe Giles and Masud Mansuripur. Possible sources of coercivity in thin films of amorphous rare earth-transition metal alloys. Computers in Physics, 5(2):204–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168409.

Huggins:1991:UMO

[461] Elisha Huggins, Samuel Vélez, Keith Dunleavy, and Anuj Gupta. Use of the Macintosh oscilloscope in undergraduate science education. Computers in Physics, 5(2):220–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822980.

Srivastava:1991:HCI

[462] Niraj Srivastava, Charles Kaufman, and Gerhard Müller. Hamiltonian chaos II. Computers in Physics, 5(2):239–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822981.

Press:1991:BFF

[463] William H. Press and Saul A. Teukolsky. Bessel functions of fractional order. Computers in Physics, 5(2):244–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822982.

Poshusta:1991:SEA

[464] R. D. Poshusta. Spreadsheet error analysis of least-squares fitted models using Monte Carlo simulation. Computers in Physics, 5(2):248–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822983.

Anonymous:1991:LLRb

[465] Anonymous. Listings of the latest releases. Computers in Physics, 5(2): 253–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822984.

Bacon:1991:PC

[466] Ben Bacon. Physics Computing ’91. Computers in Physics, 5(2):256–??, March 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822985.

Anonymous:1991:PCC

[467] Anonymous. Physics Computing ’91 Conference. Computers in Physics, 5(3):263–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822986.

Wolff:1991:VTP

[468] Robert S. Wolff. Visualization as a tool for physics education. Computers in Physics, 5(3):278–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822987.

Pickover:1991:ISS

[469] Clifford A. Pickover. An informal survey of the scientific and social impact of a soda can-sized super-supercomputer. Computers in Physics, 5 (3):290–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822988.

Martin:1991:CPU

[470] Richard F. Martin, Jr., George Skadron, and Robert D. Young. Computers, physics and the undergraduate experience. Computers in Physics, 5 (3):302–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822989.

Melcuk:1991:MDS

[471] Andrew I. Mel’čuk, Roscoe C. Giles, and Harvey Gould. Molecular dynamics simulation of liquids on the Connection Machine. Computers in Physics, 5(3):311–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168420.

Belinfante:1991:CDA

[472] F. J. Belinfante. Computing Dirac’s atomic hydrogen wave functions of the continuum, using summation of mathematically divergent series. Computers in Physics, 5(3):319–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168410.

Hall:1991:GPI

[473] B. D. Hall. A general purpose interactive programmable laboratory interface system using the IEEE-488 bus. Computers in Physics, 5(3):323–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168411.

Stauffer:1991:BRC

[474] Dietrich Stauffer. Book review: Complex System Dynamics — An Introduction to Automata Networks, Gerard Welsbuch 208 pp., Addison-Wesley, 1991. Computers in Physics, 5(3):329–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822990.

Press:1991:MBF

[475] William H. Press and Saul A. Teukolsky. Modified Bessel functions of fractional order. Computers in Physics, 5(3):330–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822991.

Potvin:1991:CST

[476] Jean Potvin. Computation of the surface tension. Computers in Physics, 5(3):333–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822992.

Wolff:1991:VFP

[477] Robert S. Wolff. Visualization of floating point data: You don’t always get what you see. Computers in Physics, 5(3):339–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822993.

Tam:1991:PM

[478] Patrick Tam. Physics and Mathematica^{T M}. Computers in Physics, 5(3): 342–348, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822994.

Mallinckrodt:1991:IPV

[479] A. John Mallinckrodt. Interactive physics version 1.2. Computers in Physics, 5(3):349–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822995.

Anonymous:1991:LLRc

[480] Anonymous. Listings of the latest releases. Computers in Physics, 5(3): 355–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822996.

Rigden:1991:PID

[481] John S. Rigden. Physics: Ideas and devices? Or, devices and ideas? Computers in Physics, 5(3):360–??, May 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822997.

Noble:1991:HRF

[482] Julian V. Noble. Highlights of the 1990 Rochester FORTH Conference. Computers in Physics, 5(4):367–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822998.

Zeyher:1991:CCP

[483] Allen Zeyher. Cray constructs powerful chemistry. Computers in Physics, 5(4):369–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822999.

Gekelman:1991:RTR

[484] Walter Gekelman, James Maggs, and Lingyu Xu. Real-time relativity. Computers in Physics, 5(4):372–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823000.

Noble:1991:AD

[485] Julian V. Noble. Avoid decisions. Computers in Physics, 5(4):386–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823001.

Krueger:1991:ATT

[486] Wolfgang Krueger. The application of transport theory to visualization of 3-D scalar data fields. Computers in Physics, 5(4):397–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168412.

Sanchez:1991:NPS

[487] Héctor Jorge Sánchez. A new peak search routine for fast evaluation on small computers. Computers in Physics, 5(4):407–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168413.

Hall:1991:CDS

[488] B. D. Hall and R. Monot. Calculating the Debye–Scherrer diffraction pattern for large clusters. Computers in Physics, 5(4):414–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168397.

Shoemaker:1991:TPP

[489] R. L. Shoemaker, H. H. Barrett, A. Landesman, R. H. Seacat III, and B. B. Taylor. The TRIMM parallel processor. Computers in Physics, 5 (4):418–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168398.

Press:1991:SAO

[490] William H. Press and Saul A. Teukolsky. Simulated annealing optimization over continuous spaces. Computers in Physics, 5(4):426–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823002.

Bak:1991:CSO

[491] Per Bak. Catastrophes and self-organized criticality. Computers in Physics, 5(4):430–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823003.

Hughes:1991:BTR

[492] Peter Hughes. Building a terrain renderer. Computers in Physics, 5(4): 434–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823004.

Tam:1991:PMU

[493] Patrick Tam. Physics and Mathematica^{T M}. Computers in Physics, 5(4): 438–442, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823005.

Gastineau:1991:S

[494] John E. Gastineau. Spacetime. Computers in Physics, 5(4):443–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823006.

Anonymous:1991:LLRd

[495] Anonymous. Listings of the latest releases. Computers in Physics, 5(4): 446–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823007.

Rigden:1991:CFC

[496] John S. Rigden. Cold fusion with computers? Computers in Physics, 5 (4):448–??, July 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823008.

Anderson:1991:HPC

[497] David Anderson and Elaine Oran. Highlights of Physics Computing ’91. Computers in Physics, 5(5):455–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823009.

Gustafson:1991:CP

[498] Karl Gustafson. Computational physics. Computers in Physics, 5(5): 457–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823010.

Carter:1991:AA

[499] Paul A. Carter. Avoid assumptions. Computers in Physics, 5(5):458–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823011.

Seel:1991:MSR

[500] Max Seel. Modeling of solid rocket fuel: From quantum chemistry to fluid dynamics simulations. Computers in Physics, 5(5):460–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823012.

Harter:1991:NGN

[501] William G. Harter. Nothing going nowhere fast: Computer graphics in physics courses. Computers in Physics, 5(5):466–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823013.

Pawley:1991:SMM

[502] G. S. Pawley. A simple model for the molecular dynamics of condensed phases. Computers in Physics, 5(5):479–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168399.

Folland:1991:AIA

[503] N. O. Folland. On accurate integration algorithms. Computers in Physics, 5(5):483–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168400.

Spencer:1991:PPI

[504] C. D. Spencer and E. J. Herrmann. Parallel I/O ports for the IBM microchannel. Computers in Physics, 5(5):488–??, September 1991. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168401.

Healy:1991:PNU

[505] Liam Healy and Étienne Deprit. Paint by number: Uncovering phase flows of an integrable dynamical system. Computers in Physics, 5(5):491– ??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168402.

Korsch:1991:FMI

[506] H. J. Korsch and A. Wagner. Fractal mirror images and chaotic scattering. Computers in Physics, 5(5):497–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168403.

Hon:1991:NMV

[507] David Hon. A new method for visualizing data on a sphere. Computers in Physics, 5(5):505–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168404.

Press:1991:MMBa

[508] William H. Press and Saul A. Teukolsky. Multigrid methods for boundary value problems. I. Computers in Physics, 5(5):514–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823014.

Strandburg:1991:QAT

[509] Katherine J. Strandburg. Quasicrystals and aperiodic tilings. Computers in Physics, 5(5):520–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823015.

Dory:1991:CPS

[510] Robert A. Dory. Contour plots for spreadsheets. Computers in Physics, 5(5):529–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823016.

Parker:1991:WCP

[511] G. W. Parker. What is the capacitance of parallel plates? Computers in Physics, 5(5):534–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823017.

Gatland:1991:DUM

[512] Ian R. Gatland. The Daedalon ultrasonic measurement system. Computers in Physics, 5(5):541–??, September 1991. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823018.

Anonymous:1991:LLRe

[513] Anonymous. Listings of the latest releases. Computers in Physics, 5(5): 550–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823019.

Laws:1991:RCI

[514] Priscilla W. Laws. The role of computers in introductory physics courses. Computers in Physics, 5(5):552–??, September 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823020.

Zeyher:1991:SSP

[515] Allen Zeyher. Solar simulations point to sonic shock waves. Computers in Physics, 5(6):559–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823021.

Holmes:1991:ETP

[516] Lewis M. Holmes. Exhibitors target physics users. Computers in Physics, 5(6):561–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823022.

Smarr:1991:VCI

[517] Larry Smarr. Visualization captures the imagination of physicists. Computers in Physics, 5(6):564–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823023.

Dubois:1991:OOP

[518] Paul F. Dubois. Object-oriented programming creates a “software revolution”. Computers in Physics, 5(6):568–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823024.

Grubb:1991:WEM

[519] David P. Grubb and Robert R. Borchers. Workstations Emerge as major scientific resource. Computers in Physics, 5(6):571–??, November/ December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823025.

Davidson:1991:CRB

[520] Arthur Davidson. Chaos researchers bring diverse phenomena into focus. Computers in Physics, 5(6):574–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823026.

Crandall:1991:SSS

[521] Richard E. Crandall. Symbolic software supports new adventures in physics. Computers in Physics, 5(6):576–579, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823027.

Wilson:1991:CSB

[522] Jack M. Wilson. Computer software has begun to change physics education. Computers in Physics, 5(6):580–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823028.

Leuze:1991:PCT

[523] Michael R. Leuze and Robert C. Ward. Parallel computers take lead in complex problem solving. Computers in Physics, 5(6):582–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823029.

Boghosian:1991:LGI

[524] Bruce M. Boghosian. Lattice gases illustrate the power of cellular automata in physics. Computers in Physics, 5(6):585–??, November/ December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823030.

Weiszmann:1991:CMN

[525] A. N. Weiszmann. Computer modeling in nonlinear optics: 4w-mixing theory. Computers in Physics, 5(6):591–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168414.

Visscher:1991:FEA

[526] P. B. Visscher. A fast explicit algorithm for the time-dependent Schrödinger equation. Computers in Physics, 5(6):596–??, November/ December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168415.

Metcalf:1991:DDT

[527] Michael Metcalf. A derived-data type for data analysis. Computers in Physics, 5(6):599–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168416.

Migliore:1991:CSN

[528] M. Migliore, G. F. Ayala, and S. L. Fornili. Concurrent simulation of neurons on multi-transputer systems. Computers in Physics, 5(6): 605–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168417.

Dill:1991:IMM

[529] Dan Dill. Implementing matrix mechanics in Mathematica: Determination of Clebsch–Gordan coefficients by matrix diagonalization. Computers in Physics, 5(6):616–625, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168419.

Press:1991:MMBb

[530] William H. Press and Saul A. Teukolsky. Multigrid methods for boundary value problems. II. Computers in Physics, 5(6):626–??, November/ December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823031.

Slanic:1991:DCH

[531] Zoran Slanic, Harvey Gould, and Jan Tobochnik. Dynamics of the classical Heisenberg chain. Computers in Physics, 5(6):630–??, November/ December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823032.

Donnelly:1991:CAW

[532] Denis Donnelly. CIP announces winners of second annual software contest. Computers in Physics, 5(6):636–??, November/December 1991. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823033.

Foster:1991:MUM

[533] Kenneth R. Foster. Mathematics^{T M} 2.0: An updated math program for MS-DOS platforms. Computers in Physics, 5(6):643–644, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823034.

Anonymous:1991:LLRf

[534] Anonymous. Listings of the latest releases. Computers in Physics, 5 (6):645–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823035.

Borchers:1991:PTS

[535] Robert R. Borchers. Physicists take stock of computing advances. Computers in Physics, 5(6):648–??, November/December 1991. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823036.

Zeyher:1992:MPS

[536] Allen Zeyher. Massive parallelism supports lattice QCD calculations. Computers in Physics, 6(1):7–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823045.

Borchers:1992:AMB

[537] Robert R. Borchers. Albuquerque meeting brings news of supercomputer advances. Computers in Physics, 6(1):8–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823046.

DeBoer:1992:HAG

[538] Barry G. DeBoer. HCP arrangements give identical structures. Computers in Physics, 6(1):10–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823037.

Hsiung:1992:RSG

[539] Ping-Kang Hsiung. Researchers should get credit. Computers in Physics, 6(1):10–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823038.

Anonymous:1992:Ca

[540] Anonymous. Calendar. Computers in Physics, 6(1):11–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823040.

Bourianoff:1992:APR

[541] George Bourianoff and Richard Talman. Accelerator progress relies on computational simulations. Computers in Physics, 6(1):14–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823041.

Baer:1992:LSH

[542] Howard Baer and William F. Long. Latest software, hardware serve highenergy phenomenology. Computers in Physics, 6(1):24–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823042.

Bitar:1992:LFS

[543] Khalil M. Bitar and Urs M. Heller. Lattice field simulations Press the limits of computational physics. Computers in Physics, 6(1):33–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823043.

Corden:1992:CMC

[544] M. J. Corden, C. H. Georgiopoulos, S. L. Linn, and S. Youssef. Computers manage complexity of high-energy experiments. Computers in Physics, 6(1):41–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823044.

Giles:1992:MTF

[545] R. Giles, P. S. Alexopoulos, and M. Mansuripur. Micromagnetics of thin film cobalt-based media for magnetic recording. Computers in Physics, 6(1):53–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168445.

Cameron:1992:NFM

[546] Seth A. Cameron. Novel Fourier methods for boundary value problems. Computers in Physics, 6(1):71–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168446.

Stone:1992:PLC

[547] David H. Stone. Pulsed laser codes for use as instructional tools. Computers in Physics, 6(1):77–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168447.

Press:1992:PA

[548] William H. Press and Saul A. Teukolsky. Padé approximants. Computers in Physics, 6(1):82–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823047.

Srivastava:1992:HCI

[549] Niraj Srivastava, Charles Kaufman, and Gerhard Müller. Hamiltonian chaos III. Computers in Physics, 6(1):84–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823048.

Ehrlich:1992:SCD

[550] Robert Ehrlich, Maria Dworzecka, and William M. MacDonald. Software consortium develops simulations for nine physics courses. Computers in Physics, 6(1):90–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823049.

Lane:1992:XAA

[551] Barton G. Lane. X-arRAY: Accessing extended memory on the cheap. Computers in Physics, 6(1):97–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823050.

Anonymous:1992:LLRa

[552] Anonymous. Listings of the latest releases. Computers in Physics, 6(1): 99–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823051.

Lannutti:1992:CUH

[553] Joseph E. Lannutti. Computers undergird high-energy physics research. Computers in Physics, 6(1):104–??, January 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823039.

Anonymous:1992:OEB

[554] Anonymous. Orlando exhibit brings news of educational software advances. Computers in Physics, 6(2):111–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823052.

Kalantar:1992:EVD

[555] A. H. Kalantar. Effective variance doesn’t work. Computers in Physics, 6(2):116–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823053.

Karlow:1992:DUF

[556] Edwin A. Karlow. Decoder uses fewer ICs. Computers in Physics, 6(2): 116–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823054.

Anonymous:1992:Cb

[557] Anonymous. Calendar. Computers in Physics, 6(2):120–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823055.

Giles:1992:CCD

[558] Roscoe Giles and Claudio Rebbi. Computational center develops practical experience in parallel computing. Computers in Physics, 6(2):122–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823056.

Sprott:1992:SPC

[559] J. C. Sprott. Simple programs create 3-D images. Computers in Physics, 6(2):132–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823057.

Brooks:1992:MPO

[560] Eugene D. Brooks III. Massive parallelism overcomes shared-memory limitations. Computers in Physics, 6(2):139–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823058.

Zukas:1992:HSV

[561] J. A. Zukas, J. R. Furlong, and S. B. Segletes. Hydrocodes support visualization of shock-wave phenomena. Computers in Physics, 6(2): 146–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823059.

Hoover:1992:MPC

[562] William G. Hoover, Anthony J. De Groot, and Carol G. Hoover. Massively parallel computer simulation of plane-strain elastic–plastic flow via nonequilibrium molecular dynamics and Lagrangian continuum mechanics. Computers in Physics, 6(2):155–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168448.

Liu:1992:VCP

[563] G. Liu, S. Fuelling, and R. Bruch. Versatile CAMAC–PC/AT system for data acquisition and control in high-resolution extreme-ultraviolet (EUV) spectroscopy following ion–atom and ion–molecule collisions. Computers in Physics, 6(2):168–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168449.

Somsky:1992:MPI

[564] W. R. Somsky and J. E. Gubernatis. A massively parallel implementation of the worldline quantum Monte Carlo method. Computers in Physics, 6(2):178–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168450.

Press:1992:ASR

[565] William H. Press and Saul A. Teukolsky. Adaptive stepsize Runge–Kutta integration. Computers in Physics, 6(2):188–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823060.

Lee:1992:GFM

[566] Michael A. Lee and Kevin E. Schmidt. Green’s function Monte Carlo. Computers in Physics, 6(2):192–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823061.

Dory:1992:FEM

[567] Robert A. Dory. Finite-element method in a spreadsheet. Computers in Physics, 6(2):198–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823062.

Wilson:1992:CUPa

[568] Jack M. Wilson and Edward F. Redish. The comprehensive unified physics learning environment: Part I. Background and system operation. Computers in Physics, 6(2):202–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823063.

Renwick:1992:PFY

[569] Stephen P. Renwick. Plotting and fitting your data: SigmaPlot for the Mac. Computers in Physics, 6(2):210–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823064.

Anonymous:1992:LLRb

[570] Anonymous. Listings of the latest releases. Computers in Physics, 6(2): 213–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823065.

Bailey:1992:HUT

[571] David H. Bailey. How useful are Today’s parallel computers? Computers in Physics, 6(2):216–??, March 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823066.

Zeyher:1992:MPF

[572] Allen Zeyher. Minimal principle facilitates structure determination of large molecule. Computers in Physics, 6(3):223–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823067.

Hecht:1992:GSC

[573] Frederick Hecht. Guidelines should cover illustrations too. Computers in Physics, 6(3):226–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823068.

Bowline:1992:AOP

[574] Stephen Bowline, James Utt, and Lowell F. Williams. Article omits powerful parallel machine. Computers in Physics, 6(3):226–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823069.

Anonymous:1992:Cc

[575] Anonymous. Calendar. Computers in Physics, 6(3):227–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823070.

Johnson:1992:CSR

[576] Christopher R. Johnson, Robert S. MacLeod, and Mike A. Matheson. Computer simulations reveal complexity of electrical activity in the human thorax. Computers in Physics, 6(3):230–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823071.

Briggs:1992:CUM

[577] James M. Briggs and J. Andrew McCammon. Computation unravels mysteries of molecular biophysics. Computers in Physics, 6(3):238–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823072.

Jacob:1992:DIY

[578] Robert Jacob and John Anderson. Do-it-yourself massively parallel supercomputer does useful physics. Computers in Physics, 6(3):244–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823073.

Hsieh:1992:HSH

[579] Alexander Hsieh and Eran Yehudai. HIP: Symbolic high-energy physics calculations. Computers in Physics, 6(3):253–261, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168451.

Brown:1992:PMN

[580] Stewart A. Brown. PANACEA: a model for numerical simulation codes. Computers in Physics, 6(3):262–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168424.

Press:1992:FSL

[581] William H. Press and Saul A. Teukolsky. Fitting straight line data with errors in both coordinates. Computers in Physics, 6(3):274–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823074.

Silverman:1992:ASA

[582] Amihai Silverman and Joan Adler. Animated simulated annealing. Computers in Physics, 6(3):277–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823076.

Wilson:1992:CUPb

[583] Jack M. Wilson, Edward F. Redish, and Denis Donnelly. The comprehensive unified physics learning environment: Part II. The basis for integrated studies. Computers in Physics, 6(3):282–??, May 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823077.

Wolff:1992:SI

[584] Robert S. Wolff. Sounding out images. Computers in Physics, 6(3): 287–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823078.

Mandell:1992:IIH

[585] Myron Mandell and Margaret Gjertsen. IRIS indigo: High-performance low-end workstation. Computers in Physics, 6(3):290–??, May 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823079.

Anonymous:1992:LLRc

[586] Anonymous. Listings of the latest releases. Computers in Physics, 6(3): 293–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823080.

Wolynes:1992:WCD

[587] Peter G. Wolynes. Will computer design become a matter of evolution? Computers in Physics, 6(3):296–??, May 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823081.

Anonymous:1992:PCO

[588] Anonymous. Physics Computing ’92 offers varied program. Computers in Physics, 6(4):303–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823082.

Anonymous:1992:IA

[589] Anonymous. Index of authors. Computers in Physics, 6(4):328–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823083.

Bowler:1992:PLW

[590] K. C. Bowler, L. K. Chantler, D. C. Heggie, R. D. Kenway, D. J. Tildesley, A. S. Trew, and D. J. Wallace. Physics leads the way at Edinburgh HPC facilities. Computers in Physics, 6(4):334–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823084.

Mortensen:1992:JPC

[591] Paul Mortensen. Japanese physicists command powerful supercomputer resources. Computers in Physics, 6(4):339–345, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823085.

Bryson:1992:VRT

[592] Steve Bryson. Virtual reality takes on real physics applications. Computers in Physics, 6(4):346–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823086.

Hobbie:1992:MSU

[593] Russell K. Hobbie. MacDose: a simulation for understanding radiological physics. Computers in Physics, 6(4):355–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168425.

Estrera:1992:LAO

[594] J. P. Estrera, W. M. Duncan, and S. R. Slaughter. Lineshape analysis for optical transitions in III-v semiconductors using a sequential simplex procedure. Computers in Physics, 6(4):360–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168426.

Stoughton:1992:UMR

[595] C. Stoughton and D. J. Summers. Using multiple RISC CPUs in parallel to study charm quarks. Computers in Physics, 6(4):371–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168427.

Kennedy:1992:MBM

[596] Jack Kennedy. A MATHCAD-based method for enhanced system resolution studies. Computers in Physics, 6(4):377–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168452.

Thompson:1992:ANL

[597] William J. Thompson. Algorithms for normalizing by least squares. Computers in Physics, 6(4):386–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168428.

Stein:1992:CGH

[598] Alan D. Stein, Zhiyue Wang, and John S. Leigh, Jr. Computer-generated holograms: a simplified ray-tracing approach. Computers in Physics, 6 (4):389–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168429.

Walker:1992:TMA

[599] James S. Walker and J. Gathright. A transfer-matrix approach to one-dimensional quantum mechanics using Mathematica. Computers in Physics, 6(4):393–399, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168430.

Press:1992:BGM

[600] William H. Press and Saul A. Teukolsky. Biconjugate gradient method for sparse linear systems. Computers in Physics, 6(4):400–??, July 1992.

CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823087.

Stauffer:1992:IMS

[601] Dietrich Stauffer and Ras B. Pandey. Immunologically motivated simulations of cellular automata. Computers in Physics, 6(4):404–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823088.

Cook:1992:CSSa

[602] David M. Cook, Russell Dubisch, Glenn Sowell, Patrick Tam, and Denis Donnelly. A comparison of several symbol-manipulating programs: Part I. Computers in Physics, 6(4):411–420, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823089.

Wolff:1992:VVB

[603] Robert S. Wolff. Volume Visualization I: Basic concepts and applications. Computers in Physics, 6(4):421–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823090.

Morrison:1992:AMD

[604] Foster Morrison and Susan McKay. The art of modeling dynamic systems: Forecasting for chaos, randomness, and determinism. Computers in Physics, 6(4):424–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823091.

Gastineau:1992:BRF

[605] John Gastineau, A. John Mallinckrodt, and Susan McKay. Book reviews: Foster Morrison, The Art 01 Modeling Dynamic Systems: Forecasting for Chaos, Randomness, and Determinism. D. Stauffer and H. E. Stanley, From Newton to Mandelbrot: a Primer in Theoretical Physics. Computers in Physics, 6(4):424–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823092.

Roper:1992:EFE

[606] L. David Roper and Margaret Gjertsen. EXP formats equations at low cost. Computers in Physics, 6(4):426–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823093.

Anonymous:1992:LLRd

[607] Anonymous. Listings of the latest releases. Computers in Physics, 6(4): 428–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823094.

deGroot:1992:CEC

[608] Robert de Groot. Cost-effective computers counter threats to physics research and education. Computers in Physics, 6(4):432–??, July 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823095.

Zeyher:1992:OOM

[609] Allen Zeyher. Object-oriented modeling guides SSC DAQ designers. Computers in Physics, 6(5):439–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823096.

Anonymous:1992:Cd

[610] Anonymous. Calendar. Computers in Physics, 6(5):441–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823097.

Omohundro:1992:SPN

[611] Stephen M. Omohundro. Sather provides nonproprietary access to object-oriented programming. Computers in Physics, 6(5):444–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823098.

Haney:1992:CPU

[612] Scott W. Haney and James A. Crotinger. C++ proves useful in writing a tokamak systems code. Computers in Physics, 6(5):450–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823099.

Nishimura:1992:DAM

[613] Hiroshi Nishimura. Dynamic accelerator modeling uses objects in Eiffel. Computers in Physics, 6(5):456–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823100.

Cant:1992:HSS

[614] G. P. Cant. The high-speed single-pass bin sort. Computers in Physics, 6(5):469–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168432.

Smith:1992:PCS

[615] D. A. Smith. Parametric cubic spline-fitting programs for open and closed curves. Computers in Physics, 6(5):472–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168433.

Tong:1992:MSS

[616] P. Y. Tong and K. W. Yu. Multifractal scaling in a Sierpiński gasket. Computers in Physics, 6(5):478–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168434.

Burns:1992:VNR

[617] Marshall Burns. Visualizing nonlinear resonance in classical and quantum mechanics. Computers in Physics, 6(5):483–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168435.

Kristyan:1992:RKS

[618] Sandor Kristyan and Janos Szamosi. Reaction kinetic surfaces and isosurfaces of the catalytic hydrogenolysis of ethane and its self-poisoning over Ni and Pd catalysts. Computers in Physics, 6(5):494–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168436.

Andrew:1992:STG

[619] Keith Andrew and Charles G. Fleming. Space–time geometries characterized by solutions to the geodesic equations. Computers in Physics, 6 (5):498–505, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168437.

Meredith:1992:NPF

[620] Roger W. Meredith. Numeric precision in FORTRAN computing. Computers in Physics, 6(5):506–512, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168438.

Savage:1992:QMC

[621] C. M. Savage. The quantum mechanics of a classically chaotic dissipative system. Computers in Physics, 6(5):513–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168439.

Press:1992:PRN

[622] William H. Press and Saul A. Teukolsky. Portable random number generators. Computers in Physics, 6(5):522–524, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823101.

Argyrakis:1992:SDC

[623] Panos Argyrakis, Harvey Gould, and Jan Tobochnik. Simulation of diffusion-controlled chemical reactions. Computers in Physics, 6(5): 525–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823102.

Cook:1992:CSSb

[624] David M. Cook, Russell Dubisch, Glenn Sowell, Patrick Tarn, and Denis Donnelly. A comparison of several symbol-manipulating programs: Part II. Computers in Physics, 6(5):530–??, September/October 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823103.

Gjertsen:1992:LTL

[625] Margaret Gjertsen. Languages and tools let you program your microcomputer. Computers in Physics, 6(5):541–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823105.

Anonymous:1992:LLRe

[626] Anonymous. Listings of the latest releases. Computers in Physics, 6 (5):551–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823106.

Rapaport:1992:BRD

[627] Dennis C. Rapaport, A. John Mallinckrodt, and Susan McKay. Book review: D. W. Heermann and A. N. Burkitt, Parallel Algorithms in Computational Science. Computers in Physics, 6(5):554–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823107.

Mallinckrodt:1992:BRC

[628] A. John Mallinckrodt. Book review: Charles A. Whitney, Random Processes in Physical Systems: An Introduction to Probability-Based Computer Simulations. Computers in Physics, 6(5):555–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823108.

Robson:1992:BRP

[629] John W. Robson, A. John Mallinckrodt, and Susan McKay. Book review: Physics: Cinema Classics. Computers in Physics, 6(5):556–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823109.

Dubois:1992:TSN

[630] Paul F. Dubois. Try something new: Object-oriented thinking. Computers in Physics, 6(5):560–??, September/October 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823110.

Zeyher:1992:SDE

[631] Allen Zeyher. Software deciphers echoes of the big bang. Computers in Physics, 6(6):565–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823111.

Stauffer:1992:CPV

[632] Dietrich Stauffer. Computational physicists vie for the gold. Computers in Physics, 6(6):567–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823112.

Giles:1992:SPI

[633] Roscoe Giles. SIGGRAPH presents insight through images. Computers in Physics, 6(6):568–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823113.

Donnelly:1992:SED

[634] Denis Donnelly. Science educators discuss benefits of symbolic computation. Computers in Physics, 6(6):571–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823114.

Bajwa:1992:AWM

[635] Param D. Bajwa. AAPT winter meeting accents growth of computers in physics. Computers in Physics, 6(6):572–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823115.

Bajwa:1992:OAP

[636] Param D. Bajwa. Organizers announce plans for Physics Computing ’93. Computers in Physics, 6(6):573–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823116.

Anonymous:1992:Ce

[637] Anonymous. Calendar. Computers in Physics, 6(6):574–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823117.

Butler:1992:VBC

[638] David M. Butler and Steve Bryson. Vector-bundle classes form powerful tool for scientific visualization. Computers in Physics, 6(6):576– ??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823118.

Ghoshal:1992:SRS

[639] Uttam Ghoshal and T. Van Duzer. Superconductivity researchers seek to remove computational bottlenecks. Computers in Physics, 6(6):585– ??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823119.

McAulay:1992:RLO

[640] Alastair D. McAulay. Researchers look to optics to move computer technology forward. Computers in Physics, 6(6):594–??, November 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823120.

Leung:1992:ILA

[641] P. W. Leung and Paul E. Oppenheimer. Implementation of the Lanczos algorithm for the Hubbard model on the Connection Machine system. Computers in Physics, 6(6):603–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168440.

Fu:1992:IEN

[642] Hong Fu, R. Giles, M. Mansuripur, and G. Patterson. Investigation of the effects of nanostructure on the observable behavior of magnetic thin film using large-scale computer simulation. Computers in Physics, 6(6):610– ??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168441.

Collins:1992:RNG

[643] J. J. Collins, M. Fanciulli, R. G. Hohlfeld, D. C. Finch, G. v. H. Sandri, and E. S. Shtatland. A random number generator based on the logit transform of the logistic variable. Computers in Physics, 6(6):630– ??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168442.

Al-Beteri:1992:MCE

[644] A. A. Al-Beteri and D. E. Raeside. A Monte Carlo electron transport code for the desktop computer. Computers in Physics, 6(6):633– ??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168443.

Chen:1992:HRT

[645] Shiyi Chen and Xiaowen Shan. High-resolution turbulent simulations using the Connection Machine-2. Computers in Physics, 6(6):643–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168444.

Cusumano:1992:GSA

[646] Joseph P. Cusumano, Derchyan Lin, Kevin Morooney, and Louis J. Pepe. Global sensitivity analysis of a nonlinear system using animated basins of attraction. Computers in Physics, 6(6):647–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168421.

Quine:1992:DMF

[647] Richard W. Quine, George Rinard, Donald E. Rugg, Gareth R. Eaton, and Sandra S. Eaton. Design of magnetic-field gradient coils for imaging. Computers in Physics, 6(6):656–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168422.

Fromont:1992:LDB

[648] Bruno Fromont and Pascale Beaufumé. Learning dynamical behaviors with explorer. Computers in Physics, 6(6):660–??, November 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168423.

Press:1992:FIC

[649] William H. Press and Saul A. Teukolsky. Fresnel integrals, cosine and sine integrals. Computers in Physics, 6(6):670–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823121.

Tobochnik:1992:QMC

[650] Jan Tobochnik, George Batrouni, and Harvey Gould. Quantum Monte Carlo on a lattice. Computers in Physics, 6(6):673–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823122.

Dory:1992:WCA

[651] Robert A. Dory and Jeffrey H. Harris. Wavelets characterize aperiodic data. Computers in Physics, 6(6):681–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823123.

Donnelly:1992:CTA

[652] Denis Donnelly. CIP’s third annual software contest: The winners. Computers in Physics, 6(6):686–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823124.

Wolff:1992:VVI

[653] Robert S. Wolff. Volume visualization II: Ray-tracing of volume data. Computers in Physics, 6(6):692–??, November 1992. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823125.

Heil:1992:BRC

[654] Christopher Heil. Book review: Charles K. Chui, An Introduction to Wavelets. Computers in Physics, 6(6):697–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823126.

Heil:1992:BRI

[655] Christopher Heil. Book reviews: Ingrid Daubechies, Ten Lectures on Wavelets. Computers in Physics, 6(6):697, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823127.

Lundberg:1992:BRW

[656] Matthew Lundberg. Book reviews: Wavelets: a Tutorial in Theory and Applications (Volume 2 in series ‘Wavelet Analysis and its Applications’) Charles K. Chui Academic Press, Harcourt Brace Jovanovich, San Diego, CA, 1992; ISBN 0-12-174590-2, 736 pp., hardcover, $69.95. Wavelets and Their Applications Mary Beth Ruskai, Gregory Beylkin, Ronald Coifman, Ingrid Daubechies, Stephane Mallat, Yves Meyer and Louise Raphael, editors Jones and Bartlett, Boston, MA, 1992; ISBN 0-86720-225-4, 480 pp., hardcover, $59.95. Computers in Physics, 6(6):698– ??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823128.

Locklair:1992:EWH

[657] Gary Locklair. Electronics workbench helps students to learn about circuits. Computers in Physics, 6(6):700–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823129.

Anonymous:1992:LLRf

[658] Anonymous. Listings of the latest releases. Computers in Physics, 6(6): 706–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823130.

Anonymous:1992:IV

[659] Anonymous. Index to volume 6. Computers in Physics, 6(6):709–??, November 1992. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823131.

Whalen:1992:SOC

[660] Barry Whalen. A superconducting optical computer may appear in the near future. Computers in Physics, 6(6):712–??, November 1992. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823132.

Holmes:1993:BYT

[661] Lewis M. Holmes. Bringing you three new departments. . .. Computers in Physics, 7(1):3–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823139.

Fuller:1993:MHN

[662] Robert G. Fuller. Multimedia — hype or no? Computers in Physics, 7(1): 5–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823147.

Holmes:1993:SST

[663] Lewis M. Holmes. Supercomputing specialists try out clusters and networks. Computers in Physics, 7(1):6, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823148.

Anonymous:1993:Ca

[664] Anonymous. Calendar. Computers in Physics, 7(1):10–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823133.

Anonymous:1993:L

[665] Anonymous. Letter. Computers in Physics, 7(1):11–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823134.

Matey:1993:GPA

[666] James R. Matey. General principles for application of computers in experimental physics. Computers in Physics, 7(1):12–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823135.

Thompson:1993:CCQ

[667] William J. Thompson. Cutting corners: Quick square roots and trig functions. Computers in Physics, 7(1):18–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823136.

Zypman:1993:SPH

[668] Fredy R. Zypman and Denis Donnelly. Symbolic programming helps to teach Debye–Scherrer diffraction. Computers in Physics, 7(1):22–26, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823137.

Elder:1993:LSN

[669] Ken Elder, Harvey Gould, and Jan Tobochnik. Langevin simulations of nonequilibrium phenomena. Computers in Physics, 7(1):27–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823138.

Aubrecht:1993:BRC

[670] Gordon J. Aubrecht II, A. John Mallinckrodt, and Susan McKay. Book review: Charles Misner and Patrick Cooney, Spreadsheet Physics. Computers in Physics, 7(1):35–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823140.

Aubrecht:1993:BRR

[671] Gordon J. Aubrecht II, A. John Mallinckrodt, and Susan McKay. Book review: Robert G. Parks, Quattro Pro for Scientific and Engineering Spreadsheets. Computers in Physics, 7(1):35–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823141.

Aubrecht:1993:BRI

[672] Gordon J. Aubrecht II, A. John Mallinckrodt, and Susan McKay. Book review: Irvin H. Kral, The Excel Spreadsheet for Engineers and Scientists. Computers in Physics, 7(1):35–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823142.

Burr:1993:BRD

[673] A. F. Burr, A. John Mallinckrodt, and Susan McKay. Book review: Denis Donnelly, MathCAD for Introductory Physics. Computers in Physics, 7(1):36–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823143.

Burr:1993:BRS

[674] A. F. Burr, A. John Mallinckrodt, and Susan McKay. Book review: Sol Wieder, Introduction to Mathcad for Scientists and Engineers. Computers in Physics, 7(1):36–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823144.

Dubois:1993:PPF

[675] Paul F. Dubois and Lee Busby. Portable, powerful Fortran programs. Computers in Physics, 7(1):38–44, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823145.

Gjertsen:1993:DPC

[676] Margaret H. Gjertsen. 1993 directory of physics courseware. Computers in Physics, 7(1):45–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823146.

Anonymous:1993:LLRa

[677] Anonymous. Listings of the latest releases. Computers in Physics, 7(1): 77–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823149.

Brunini:1993:EGE

[678] A. Brunini and P. E. Zadunaisky. Estimation of global errors in the numerical solution of ordinary differential equations. Computers in Physics, 7(1):81–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168457.

Groisman:1993:CAP

[679] G. Groisman and J. R. Parker. Computer-assisted photometry using simulated annealing. Computers in Physics, 7(1):87–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168458.

Hassold:1993:FSA

[680] G. N. Hassold and Elizabeth A. Holm. A fast serial algorithm for the finite temperature quenched Potts model. Computers in Physics, 7(1): 97–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168481.

Pickett:1993:PSS

[681] Robert C. Pickett, Gordon E. Lindgren, and Ronald K. Anderson. Power series solution of linear and nonlinear partial differential equations via symbolic programming. Computers in Physics, 7(1):108–??, January 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168456.

Holmes:1993:FCN

[682] Lewis M. Holmes. Focusing on clusters and network. Computers in Physics, 7(2):115–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823150.

Mirchandaney:1993:CVM

[683] Ravi Mirchandaney. Cluster vs. monolithic parallel machines: You may need both. Computers in Physics, 7(2):117–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823151.

Lewin:1993:HGC

[684] David I. Lewin. HPCC grandly challenges physics community. Computers in Physics, 7(2):118–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823152.

Anonymous:1993:Cb

[685] Anonymous. Calendar. Computers in Physics, 7(2):128–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823153.

Matey:1993:CRB

[686] James R. Matey and Phillip Stein. Catching the right bus I: An overview of architectures for instrument systems. Computers in Physics, 7(2): 130–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823154.

Sherwood:1993:CCP

[687] Bruce Arne Sherwood, David M. Andersen, and Denis Donnelly. Ct creates prize-winning portable physics programs. Computers in Physics, 7 (2):136–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823155.

Shriner:1993:AMC

[688] John F. Shriner and William J. Thompson. Angular-momentum coupling coefficients: New and improved algorithms. Computers in Physics, 7(2): 144–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823156.

Potvin:1993:CQF

[689] Jean Potvin, Harvey Gould, and Jan Tobochnik. Computational quantum-field theory. Computers in Physics, 7(2):149–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823157.

Mallinckrodt:1993:BRW

[690] A. John Mallinckrodt and Susan McKay. Book review: William J. Thompson, Computing for Scientists and Engineers: a Workbook of Analysis, Numerics, and Applications. Computers in Physics, 7(2):155– ??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823158.

Mallinckrodt:1993:BRK

[691] A. John Mallinckrodt and Susan McKay. Book review: K. Binder and D. Heermann and Lyle Roelofs, Monte Carlo Simulation in Statistical Physics. Computers in Physics, 7(2):156–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823159.

Wolff:1993:VVI

[692] Robert S. Wolff. Volume visualization III: Polygonally based volume rendering. Computers in Physics, 7(2):158–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823160.

Dubois:1993:PFS

[693] Paul F. Dubois and Lee Busby. Portability and the Fortran Standard. Computers in Physics, 7(2):162–165, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823161.

Dongarra:1993:IPF

[694] Jack Dongarra, G. A. Geist, Robert Manchek, and V. S. Sunderam. Integrated PVM framework supports heterogeneous network computing. Computers in Physics, 7(2):166–174, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL http://

www.netlib.org/utk/papers/comp-phy7/comp-phy7.html;   http:
//www.netlib.org/utk/papers/comp-phy7/comp-phy7.ps;   http:
//www.netlib.org/utk/people/JackDongarra/pdf/comp-phy7.pdf;
 https://aip.scitation.org/doi/10.1063/1.4823162.

Duke:1993:CCE

[695] Dennis W. Duke. Cluster computing exploits performance and cost advantages. Computers in Physics, 7(2):176–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823163.

Rinaldo:1993:LCP

[696] Frank Rinaldo and Stephen Wolbers. Loosely coupled parallel processing at Fermilab. Computers in Physics, 7(2):184–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823164.

Christian:1993:PUP

[697] Wolfgang Christian and Margaret H. Gjertsen. Pascal utilities provide programming assistance: Comparison of M.U.P.P.E.T. and Scientific Programmers Toolkit. Computers in Physics, 7(2):191–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823165.

Anonymous:1993:LLRb

[698] Anonymous. Listings of the latest releases. Computers in Physics, 7(2): 197–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823166.

Noels:1993:VES

[699] Arlette Noels, Roland Papy, and François Remy. Visualizing the evolution of a star on a graphic workstation. Computers in Physics, 7(2): 202–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168459.

Jolivette:1993:LSF

[700] Peter L. Jolivette. Least-squares fits when there are errors in X. Computers in Physics, 7(2):208–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168460.

Backstrom:1993:CFL

[701] G. Bäckström. Curve fitting by Lagrange interpolation. Computers in Physics, 7(2):213–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168461.

Visscher:1993:HAS

[702] P. B. Visscher, Judy Dye, and Jen-Ho Fang. A hierarchical approach to stochastic simulation of correlated systems. Computers in Physics, 7(2): 217–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168462.

Pachner:1993:NCT

[703] J. Pachner. The nature of chaos in two systems of ordinary nonlinear differential equations. Computers in Physics, 7(2):226–??, March 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168463.

Holmes:1993:PCA

[704] Lewis M. Holmes. PC ’93 convenes in Albuquerque. Computers in Physics, 7(3):251–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823167.

Negele:1993:CPC

[705] John W. Negele. In computational physics, change brings opportunities and challenges. Computers in Physics, 7(3):253–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823168.

Lewin:1993:NSF

[706] David I. Lewin. NREN seen as forerunner of national information infrastructure. Computers in Physics, 7(3):254–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823169.

Zeyher:1993:NTI

[707] Allen Zeyher. Numerical tokamaks illuminate plasma physics. Computers in Physics, 7(3):256–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823170.

Anonymous:1993:Cc

[708] Anonymous. Calendar. Computers in Physics, 7(3):258–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823171.

Thompson:1993:BTP

[709] William J. Thompson. A bargain at twice the price. Computers in Physics, 7(3):261–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823172.

Gunter:1993:HSD

[710] Berton H. Gunter and James R. Matey. How statistical design concepts can improve experimentation in the physical sciences. Computers in Physics, 7(3):262–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823173.

Ceperley:1993:LCS

[711] Peter H. Ceperley and Denis Donnelly. Low-cost system plots fields and waves in color. Computers in Physics, 7(3):274–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823174.

Gatland:1993:WWG

[712] Ian R. Gatland and William J. Thompson. A weight-watcher’s guide to least-squares fitting. Computers in Physics, 7(3):280–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823175.

Mallinckrodt:1993:BRSa

[713] A. John Mallinckrodt and Susan McKay. Book review: Simson L. Garfinkel and Michael K. Mahoney and Richard R. Silbar, NeXTSTEP Programming, Step One: Object-Oriented Applications. Computers in Physics, 7(3):287–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823176.

Mallinckrodt:1993:CHC

[714] A. John Mallinckrodt and Susan McKay. C. H. Chen and Seth Wolpert, Neural Networks in Pattern Recognition and their Applications. Computers in Physics, 7(3):288–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823177.

Janhunen:1993:MMB

[715] Pekka Janhunen, Douglas Stein, and Robert S. Wolff. MathHDF: MathLink-based distributed visualization between Mathematica and HDF files. Computers in Physics, 7(3):290–294, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823178.

Landau:1993:VFB

[716] R. H. Landau, T. Mefford, G. He, and Robert S. Wolff. Visualizations in few-body physics. Computers in Physics, 7(3):296–??, May 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823179.

Brown:1993:SPS

[717] Stewart A. Brown, Mike Folk, Gregory Goucher, Russ Rew, and Paul F. Dubois. Software for portable scientific data management. Computers in Physics, 7(3):304–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823180.

Christian:1993:NLP

[718] Wolfgang Christian and Margaret H. Gjertsen. Nextstep has lived up to its promise in an educational environment. Computers in Physics, 7(3): 309–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823181.

Mallinckrodt:1993:BRI

[719] A. John Mallinckrodt. Book review: Interactive Physics II: a physics simulation laboratory for the Macintosh. Computers in Physics, 7(3): 312–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823182.

Anonymous:1993:NPa

[720] Anonymous. New products. Computers in Physics, 7(3):317–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823183.

Anonymous:1993:LSS

[721] Anonymous. Listing of sessions, speakers, and talks at the 5th international conference on computational physics. Computers in Physics, 7(3): 319–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823184.

Anonymous:1993:WEW

[722] Anonymous. What exhibitors will have on display in Albuquerque. Computers in Physics, 7(3):328–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823185.

Roy:1993:CST

[723] Swati Roy, W. L. (Nancy) Lee, James W. Mentele, and Richard W. Nau. A comparative study of temperature extremes analysis by regression methods and by an artificial neural network. Computers in Physics, 7 (3):330–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168453.

Caprari:1993:LSO

[724] Robert S. Caprari. Least-squares orthogonal polynomial approximation in several independent variables. Computers in Physics, 7(3):336–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168454.

Morales-Acevedo:1993:SSC

[725] Arturo Morales-Acevedo and Humberto Arizpe. Simulating solar cells with a microcomputer. Computers in Physics, 7(3):343–??, May 1993.

CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168455.

Brommer:1993:ICP

[726] Karl D. Brommer, B. E. Larson, M. Needels, and J. D. Joannopoulos. Implementation of the car–Parrinello algorithm for ab initio total energy calculations on a massively parallel computer. Computers in Physics, 7 (3):350–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168464.

Li:1993:PCF

[727] Chihwen “Chris” Li and Chwan-Hwa “John” Wu. A particle-in-cell fluid model for radio frequency glow discharges. Computers in Physics, 7(3): 363–??, May 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168465.

Holmes:1993:WYL

[728] Lewis M. Holmes. Would you like to try some hands-on animation. Computers in Physics, 7(4):379–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823186.

Irving:1993:WPS

[729] Larry Irving, Jr. Why physicists should care about the national information infrastructure. Computers in Physics, 7(4):381–??, July 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823187.

Zeyher:1993:SMR

[730] Allen Zeyher. Simulated materials reveal microfractures. Computers in Physics, 7(4):382–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823188.

Anonymous:1993:Cd

[731] Anonymous. Calendar. Computers in Physics, 7(4):386–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823189.

Singh:1993:ECP

[732] Prabhakar P. Singh and William J. Thompson. Exploring the complex plane: Green’s functions, Hilbert transforms, and analytic continuation. Computers in Physics, 7(4):388–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823190.

Dengler:1993:CVM

[733] Roman Dengler, Karl Luchner, Dean Zollman, and Denis Donnelly. Computer-video method evaluates real motion data in real time for students. Computers in Physics, 7(4):393–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823191.

Lin:1993:EDM

[734] H. Q. Lin, J. E. Gubernatis, Harvey Gould, and Jan Tobochnik. Exact diagonalization methods for quantum systems. Computers in Physics, 7 (4):400–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.

4823192. See [757].

Matey:1993:HCE

[735] James R. Matey and M. J. Lauterbach. How to control errors in analogto-digital conversion. Computers in Physics, 7(4):408–??, July 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823193.

Blair:1993:BRP

[736] J. Morris Blair, A. John Mallinckrodt, and Susan McKay. Book review: Philip R. Bevington and D. Keith Robinson, Data Reduction and Error Analysis for the Physical Sciences. Computers in Physics, 7(4):415–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823194.

Siegel:1993:BRR

[737] Peter B. Siegel, A. John Mallinckrodt, and Susan McKay. Book review: Robert L. Devaney, A First Course in Chaotic Dynamical Systems: Theory and Experiment. Computers in Physics, 7(4):416–??, July 1993. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823195.

Decyk:1993:HWN

[738] Viktor K. Decyk and Paul F. Dubois. How to write (nearly) portable Fortran programs for parallel computers. Computers in Physics, 7(4):

418–425, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823196.

Wolff:1993:MCL

[739] Robert S. Wolff. Multimedia in the classroom and the laboratory. Computers in Physics, 7(4):426–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823197.

Bridges:1993:EEA

[740] David Bridges, Terry J. Deveau, and Margaret H. Gjertsen. Equation editing with Ami Pro, Word, and Wordperfect. Computers in Physics, 7 (4):444–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823198.

Anonymous:1993:NPb

[741] Anonymous. New products. Computers in Physics, 7(4):456–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823199.

Simos:1993:VSP

[742] T. E. Simos. A variable-step procedure for the numerical integration of the one-dimensional Schrödinger equation. Computers in Physics, 7(4): 460–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168466.

Payne:1993:AAS

[743] M. G. Payne and Mark Edwards. Approximate atomic structure calculations using a microcomputer. Computers in Physics, 7(4):465–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168467.

Redondo:1993:VMA

[744] Antonio Redondo and Dipen N. Sinha. Vibrational mode analysis using maximum likelihood and maximum entropy. Computers in Physics, 7(4): 476–??, July 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168468.

Holmes:1993:SGM

[745] Lewis M. Holmes. Some good moves. Computers in Physics, 7(5):483– ??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823200.

Giles:1993:WDC

[746] Roscoe Giles. What does computation have to do with “real” physics? Computers in Physics, 7(5):485–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823201.

Zeyher:1993:CTE

[747] Allen Zeyher. CFD tools expand flow analysis horizons. Computers in Physics, 7(5):486–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823202.

Bajwa:1993:FRP

[748] Param D. Bajwa. First Rahman prize winner focuses on educational reform. Computers in Physics, 7(5):488–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823203.

Stauffer:1993:CDR

[749] Dietrich Stauffer. Computers and dinosaurs on route 66. Computers in Physics, 7(5):489–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823204.

Bajwa:1993:CER

[750] Param D. Bajwa. Companies exhibit resources for computational physicists. Computers in Physics, 7(5):490–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823205.

Borchers:1993:SBH

[751] Robert R. Borchers. SC’93 brings high-performance computing to Portland. Computers in Physics, 7(5):491–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823206.

Lewin:1993:NGA

[752] David I. Lewin. NIST gives America the time of day. Computers in Physics, 7(5):492–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823207.

Anonymous:1993:Ce

[753] Anonymous. Calender. Computers in Physics, 7(5):495–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823209.

Bryson:1993:SCA

[754] Steve Bryson. Symposium considers applications of virtual reality. Computers in Physics, 7(5):495–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823208.

Matey:1993:UMI

[755] James R. Matey. UVR mark II: The history of an automated instrument. Computers in Physics, 7(5):498–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823210.

Ehrlich:1993:TMA

[756] Robert Ehrlich, Maria Dworzecka, William MacDonald, Jaroslaw Tuszynski, and Denis Donnelly. Text materials to accompany simulations for the CUPS project. Computers in Physics, 7(5):508–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823211.

Davidson:1993:MMT

[757] Ernest R. Davidson. Monster matrices: Their eigenvalues and eigenvectors. Computers in Physics, 7(5):519–522, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823212. See [734].

Oran:1993:CTS

[758] Elaine S. Oran and Jay P. Boris. Computing turbulent shear flows — a convenient conspiracy. Computers in Physics, 7(5):523–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823213.

Lam:1993:AWM

[759] Lui Lam, Rocco Pochy, Harvey Gould, and Jan Tobochnik. Active-Walker models: Growth and form in nonequilibrium systems. Computers in Physics, 7(5):534–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823214.

Bourzutschky:1993:BRJ

[760] Marc Bourzutschky, A. John Mallinckrodt, and Susan McKay. Book review: John Wendt, Computational Fluid Dynamics: An Introduction. Computers in Physics, 7(5):542–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823215.

Mallinckrodt:1993:BREa

[761] A. John Mallinckrodt and Susan McKay. Book reviews: Ed Krol, The Whole Internet User’s Guide & Catalog, . . .. Computers in Physics, 7(5):543–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823216.

Mallinckrodt:1993:BRJa

[762] A. John Mallinckrodt and Susan McKay. Book review: Jonathan Kochmer and NorthWestNet, The Internet Passport: NorthWestNet’s Guide to Our World Online (4th ed.). Computers in Physics, 7(5):543– ??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823217.

Mallinckrodt:1993:BRA

[763] A. John Mallinckrodt and Susan McKay. Book review: April Marine and Susan Kirkpatrick and Vivian Neou and Carol Ward, Internet: Getting Started. Computers in Physics, 7(5):543–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823218.

Mallinckrodt:1993:BREb

[764] A. John Mallinckrodt and Susan McKay. Book review: Elizabeth Lane and Craig Summerhill, Internet Primer for Information Professionals: a Basic Guide to Internet Networking Technology. Computers in Physics, 7(5):543–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823219.

Mallinckrodt:1993:BRB

[765] A. John Mallinckrodt and Susan McKay. Book review: Brendan P. Kehoe, Zen and the Art of the Internet: a Beginner’s Guide (2nd ed.). Computers in Physics, 7(5):543–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823220.

Mallinckrodt:1993:BRT

[766] A. John Mallinckrodt and Susan McKay. Book review: Tracy LaQuey and Jeanne C. Ryer, The Internet Companion: a Beginner’s Guide to Global Networking. Computers in Physics, 7(5):543–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823221.

Mallinckrodt:1993:BRRa

[767] A. John Mallinckrodt and Susan McKay. Book review: Roy Tennant and John Ober and Anne G. Lipow, Crossing the Internet Threshold: An Instructional Handbook. Computers in Physics, 7(5):543–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823222.

Dubois:1993:PE

[768] Paul F. Dubois. Perl by example. Computers in Physics, 7(5):545–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823223.

Balick:1993:VIT

[769] Bruce Balick and Margaret H. Gjertsen. Voyager II takes students on joy rides through the Universe. Computers in Physics, 7(5):551–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823224.

Anonymous:1993:LLRc

[770] Anonymous. Listings of the latest releases. Computers in Physics, 7(5): 555–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823225.

Moriarty:1993:OOP

[771] K. J. M. Moriarty, S. Sanielevici, K. Sun, and T. Trappenberg. Objectoriented programming applied to lattice gauge theory. Computers in Physics, 7(5):560–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168469.

Cunningham:1993:CTM

[772] Robert W. Cunningham. Comparison of three methods for determining fit parameter uncertainties for the Marquardt compromise. Computers in Physics, 7(5):570–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168470.

Johansen:1993:EIP

[773] G. A. Johansen, J. E. Nordtvedt, and P. Vold. ElPed: An interactive PC-based learning program for basic electrical circuits. Computers in Physics, 7(5):577–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168471.

Johnson:1993:QSC

[774] M. A. Johnson and L. V. Meisel. Quantifying symmetric chaos. Computers in Physics, 7(5):585–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168472.

Silbar:1993:INI

[775] Richard R. Silbar. An interactive NeXTSTEP interface to a Fortran code for solving coupled differential equations. Computers in Physics, 7(5):592–598, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168473.

Roberts:1993:AFS

[776] D. Aaron Roberts. An algorithm for finding spurious points in turbulent signals. Computers in Physics, 7(5):599–??, September 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168474.

Holmes:1993:AHS

[777] Lewis M. Holmes. Awards honor software developers and computational specialists. Computers in Physics, 7(6):611–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823226.

Redish:1993:CAT

[778] Edward F. Redish. Are computers appropriate for teaching physics? Computers in Physics, 7(6):613–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823227.

Lewin:1993:SSP

[779] David I. Lewin. Supercomputing ’93 showcases parallel computing finalists. Computers in Physics, 7(6):614–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823228.

Bajwa:1993:SDA

[780] Param D. Bajwa. Submission deadline is approaching for PC ’94. Computers in Physics, 7(6):616–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823229.

Duke:1993:CSF

[781] Dennis W. Duke. Computing specialists focus on gigabit networking. Computers in Physics, 7(6):618–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823230.

LeSar:1993:CPN

[782] Richard LeSar. Cover photo needs additional credits. Computers in Physics, 7(6):620–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823231.

Anonymous:1993:Cf

[783] Anonymous. Calendar. Computers in Physics, 7(6):621–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823232.

Mallinckrodt:1993:BRRb

[784] A. John Mallinckrodt and Susan McKay. Book review: Robert G. Parks, Quattro Pro(R) for Scientific and Engineering Spreadsheets. Computers in Physics, 7(6):624–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823233.

Mallinckrodt:1993:BRJb

[785] A. John Mallinckrodt and Susan McKay. Book review: J. M. Haile and Ian Johnston, Molecular Dynamics Simulation: Elementary Methods. Computers in Physics, 7(6):625–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823234.

Mallinckrodt:1993:BRSb

[786] A. John Mallinckrodt and Susan McKay. Book review: Samuel S. M. Wong and Philip Gash, Computational Methods in Physics and Engineering. Computers in Physics, 7(6):625–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823235.

Thompson:1993:NNA

[787] William J. Thompson. Numerous neat algorithms for the Voigt profile function. Computers in Physics, 7(6):627–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823236.

Chen:1993:DBS

[788] Hudong Chen, Harvey Gould, and Jan Tobochnik. Discrete Boltzmann systems and fluid flows. Computers in Physics, 7(6):632–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823237.

Anonymous:1993:HGC

[789] Anonymous. How to get CIP source code over the Internet. Computers in Physics, 7(6):637–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823238.

vanHaren:1993:TDA

[790] P. C. van Haren, F. Wijnoltz, and James R. Matey. The tramp dataacquisition system. Computers in Physics, 7(6):638–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823239.

Donnelly:1993:CFA

[791] Denis Donnelly. CIP’s fourth annual software contest. Computers in Physics, 7(6):647–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823240.

Wong:1993:OON

[792] Michael K. W. Wong, Kent G. Budge, James S. Peery, Allen C. Robinson, and Paul F. Dubois. Object-oriented numerics: a paradigm for numerical object-oriented programming. Computers in Physics, 7(6):655– ??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823241.

Wolff:1993:MVT

[793] Robert S. Wolff. Multiplexed visualization: Terrain rendering and Magellan data. Computers in Physics, 7(6):664–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823242.

Franta:1993:MMT

[794] W. R. Franta, Joel Halpern, and James Hughes. Mix and match technologies boost communication speed in Today’s data-networking environment. Computers in Physics, 7(6):666–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823243.

Tobochnik:1993:FCP

[795] Neal Tobochnik, Jan Tobochnik, and Margaret H. Gjertsen. Fortran compilers for personal computers. Computers in Physics, 7(6):672–677, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823244.

Anonymous:1993:LLRd

[796] Anonymous. Listings of the latest releases. Computers in Physics, 7(6): 678–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823245.

Kaiser:1993:SAC

[797] T. Kaiser and G. Schweiger. Stable algorithm for the computation of Mie coefficients for scattered and transmitted fields of a coated sphere. Computers in Physics, 7(6):682–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168475.

Ding:1993:MCS

[798] Hong Q. Ding. Monte Carlo simulations of quantum systems on massively parallel supercomputers. Computers in Physics, 7(6):687–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168476.

Esaulov:1993:HSM

[799] V. E. Esaulov, R. Bartolo, and F. Penent. Hardware and software for a multiparametric concurrent data acquisition system. Computers in Physics, 7(6):696–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168477.

vanHaren:1993:DDS

[800] P. C. van Haren. The DOM4 database system. Computers in Physics, 7 (6):701–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168478.

Meisel:1993:ELP

[801] L. V. Meisel and P. J. Cote. Exponents in lifetime and power spectral density forms in self-organized critical systems — a Mathematica application. Computers in Physics, 7(6):710–713, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168479.

Wegener:1993:DAC

[802] W. Wegener, P. D’Hooghe, and W. De Boeck. Data acquisition and control system for the neutron time-of-flight spectrometer MIBEMOL. Computers in Physics, 7(6):714–??, November 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168480.

Holmes:1994:NI

[803] Lewis M. Holmes. New in this issue. Computers in Physics, 8(1): 3–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823255.

Duax:1994:CSC

[804] William L. Duax. Crystallography: Send in the computers. Computers in Physics, 8(1):5–??, January/February 1994. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823258.

Lewin:1994:FDR

[805] David I. Lewin. Fortran’s developer receives top engineering award. Computers in Physics, 8(1):6–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823261.

Holmes:1994:SCE

[806] Lewis M. Holmes. Supercomputing conference emphasizes scalable computing and networks. Computers in Physics, 8(1):8–??, January/ February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823268.

Zeyher:1994:PFS

[807] Allen Zeyher. Pattern formation studies combine computation and experiment. Computers in Physics, 8(1):10–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823246.

Bajwa:1994:WCI

[808] Param D. Bajwa. What commercializing the Internet will mean for physicists. Computers in Physics, 8(1):12–??, January/February 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823247.

Gjertsen:1994:ECN

[809] Margaret Gjertsen. Educom ’93 crafts new communities. Computers in Physics, 8(1):14–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823248.

Anonymous:1994:Ca

[810] Anonymous. Calendar. Computers in Physics, 8(1):16–??, January/ February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823249.

Holmes:1994:IIS

[811] Lewis M. Holmes. Introducing… Internet services. Computers in Physics, 8(1):19–??, January/February 1994. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823250.

Ricart:1994:HTF

[812] Glenn Ricart. How to transfer files on the Internet. Computers in Physics, 8(1):20–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823251.

Anonymous:1994:HGC

[813] Anonymous. How to get CIP source code. Computers in Physics, 8 (1):23–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823252.

Judd:1994:AOI

[814] Peggy Judd and Frank Perugini. AIP’s online information service helps physicists stay up to date. Computers in Physics, 8(1):24–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823253.

Boyer:1994:PNM

[815] L. L. Boyer. Predicting new materials. Computers in Physics, 8(1): 27–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823254.

Cohen:1994:EPE

[816] R. E. Cohen. Electrons, phonons, and electron-phonon interactions in high-temperature superconductors. Computers in Physics, 8(1):34–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823256.

Gadre:1994:CRB

[817] Dhananjay V. Gadre, Pramod K. Upadhyay, Vijaya S. Varma, and James R. Matey. Catching the right bus, part 2: Using a parallel printer adapter as an inexpensive interface. Computers in Physics, 8 (1):45–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823257.

Silbar:1994:SSA

[818] R. R. Silbar, H. T. Williams, and William J. Thompson. Symbolic solutions: Angular momentum on a PC. Computers in Physics, 8(1): 52–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823259.

Aguirregabiria:1994:WCE

[819] J. M. Aguirregabiria, A. Hernández, M. Rivas, and Denis Donnelly. Are we careful enough when using computer algebra? Computers in Physics, 8(1):56–61, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823260.

Lai:1994:CC

[820] Ying-Cheng Lai, Harvey Gould, and Jan Tobochnik. Controlling chaos. Computers in Physics, 8(1):62–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823262.

Mallinckrodt:1994:BRR

[821] A. John Mallinckrodt and Susan McKay. Book review: Rubin H. Landau and Paul Fink, Jr. and Geert Wenes, A Scientist’s and Engineer’s Guide to Workstations and Supercomputers. Computers in Physics, 8 (1):68–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823263.

Mallinckrodt:1994:BRF

[822] A. John Mallinckrodt and Susan McKay. Book review: Francis Moon and Paul S. Linsay, Chaotic and Fractal Dynamics: An Introduction for Applied Scientists and Engineers. Computers in Physics, 8(1):69–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823264.

Dubois:1994:MAP

[823] Paul F. Dubois. Making applications programmable. Computers in Physics, 8(1):70–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823265.

Blass:1994:IWI

[824] William E. Blass, Stephen L. Mahan, and Margaret H. Gjertsen. IDL for Windows impresses as a computational system. Computers in Physics, 8(1):74–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823266.

Anonymous:1994:LLRa

[825] Anonymous. Listings of the latest releases. Computers in Physics, 8 (1):77–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823267.

Fu:1994:CMM

[826] Hong Fu, R. Giles, and M. Mansuripur. Coercivity mechanisms in magneto-optical recording media. Computers in Physics, 8(1):80–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168516.

Horwitz:1994:TSR

[827] Paul Horwitz, Edwin F. Taylor, and William Barowy. Teaching special relativity with a computer. Computers in Physics, 8(1):92–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168517.

Blackburn:1994:GSP

[828] James Kent Blackburn. GRPP, a scientific programming language preprocessor designed using lex and yacc. Computers in Physics, 8(1): 98–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168518.

Houlahan:1994:TAS

[829] Padraig Houlahan. Tracking artificial satellites with the Macintosh. Computers in Physics, 8(1):113–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168513.

Marsaglia:1994:SPV

[830] George Marsaglia and Arif Zaman. Some portable very-long-period random number generators. Computers in Physics, 8(1):117–121, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168514.

Lea:1994:CAA

[831] Suzanne M. Lea, Beth A. Thacker, Eunsook Kim, and Kathleen M. Miller. Computer-assisted assessment of student understanding in physics. Computers in Physics, 8(1):122–??, January/February 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168515.

Holmes:1994:GGS

[832] Lewis M. Holmes. On global and not-so-global searches. Computers in Physics, 8(2):131–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823269.

Haigh:1994:IEN

[833] Peter J. Haigh. ICN equals NII here and now — not “if” and “when”. Computers in Physics, 8(2):133–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823270.

Lewin:1994:NFB

[834] David I. Lewin. NSF foresees bright future for its supercomputing centers. Computers in Physics, 8(2):134–136, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823271.

Zeyher:1994:OPL

[835] Allen Zeyher. Optical packages look for global minima. Computers in Physics, 8(2):137–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823272.

Parker:1994:FRS

[836] Geoffrey Parker and Nalini Nadkarni. Forest researchers seek 3-D tools. Computers in Physics, 8(2):140–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823273.

Anonymous:1994:Cb

[837] Anonymous. Calendar. Computers in Physics, 8(2):141–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823274.

Ricart:1994:SLI

[838] Glenn Ricart. Speed limits on the Internet. Computers in Physics, 8(2): 146–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823275.

Holmes:1994:ENP

[839] Lewis M. Holmes. Expanded new-products section goes on line. Computers in Physics, 8(2):150–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823276.

Zollner:1994:SOS

[840] Stefan Zollner and James R. Matey. Selecting an operating system, Part I: OS/2 2.X. Computers in Physics, 8(2):152–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823277.

Thompson:1994:CPA

[841] William J. Thompson. Chebyshev polynomials: After the spelling the rest is easy. Computers in Physics, 8(2):161–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823278.

Abbott:1994:PHM

[842] Hilton Abbott and Denis Donnelly. Physics homework management: Using Hypercard to generate individualized problem sets. Computers in Physics, 8(2):166–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823279.

Potvin:1994:CQF

[843] Jean Potvin, Harvey Gould, and Jan Tobochnik. Computational quantum field theory. Part II: Lattice gauge theory. Computers in Physics, 8 (2):170–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823280.

Matey:1994:BRH

[844] James R. Matey, A. John Mallinckrodt, and Susan McKay. Book review: H. M. Staudenmaier, editor, Physics Experiments Using PCs. Computers in Physics, 8(2):177–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823281.

Matey:1994:BRJ

[845] James R. Matey, A. John Mallinckrodt, and Susan McKay. Book review: John C. Russ, The Image Processing Handbook. Computers in Physics, 8(2):177–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823282.

Silebi:1994:DMT

[846] C. A. Silebi, W. E. Schiesser, and Ras B. Pandey. Dynamic modeling of transport process systems. Computers in Physics, 8(2):178–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823283.

DeVries:1994:FCC

[847] Paul L. DeVries and Robert P. Wolf. A first course in computational physics. Computers in Physics, 8(2):178–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823284.

Amala:1994:PMTa

[848] Paul A. K. Amala, Garry H. Rodrigue, and Paul F. Dubois. Programming models for three-dimensional hydrodynamics on the CM-5 (Part I). Computers in Physics, 8(2):181–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823285.

Backstrom:1994:PSP

[849] Gunnar Backstrom and Margaret H. Gjertsen. PDEase2 solves partial differential equations flexibly using finite elements analysis. Computers in Physics, 8(2):185–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823286.

Shafer:1994:GOO

[850] David Shafer. Global optimization in optical design. Computers in Physics, 8(2):188–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823287.

Anonymous:1994:LLRb

[851] Anonymous. Listings of the latest releases. Computers in Physics, 8(2): 196–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823288.

Rothman:1994:BBB

[852] Daniel H. Rothman and Leo P. Kadanoff. Bubble, bubble, boil, and trouble. Computers in Physics, 8(2):199–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168524.

Hailstone:1994:TBS

[853] R. K. Hailstone. Transputer-based simulations of image recording in silver-halide materials. Computers in Physics, 8(2):205–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168519.

Diercksen:1994:ISO

[854] Geerd H. F. Diercksen and George G. Hall. Intelligent software: The OpenMol program. Computers in Physics, 8(2):215–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168520.

Frichter:1994:QBD

[855] George M. Frichter and J. Piekarewicz. Quark-based description of nuclear matter with simulated annealing. Computers in Physics, 8(2): 223–??, March 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168491.

Holmes:1994:YSS

[856] Lewis M. Holmes. Are you set up for symbolics? Computers in Physics, 8 (3):235–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823289.

Reitmeier:1994:WCH

[857] Glenn A. Reitmeier. Watch out computing here comes HDTV! Computers in Physics, 8(3):237–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823290.

Lewin:1994:MDA

[858] David I. Lewin. Mathematica’s developer asks, “can algorithms replace equations?”. Computers in Physics, 8(3):238–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823291.

Shearin:1994:PCE

[859] Rhonda Shearin. Physics Computing ’94 explores innovations in computational physics. Computers in Physics, 8(3):241–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823292.

Anonymous:1994:Cc

[860] Anonymous. Calendar. Computers in Physics, 8(3):244–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823293.

Ricart:1994:MIB

[861] Glenn Ricart. The Mosaic Internet browser. Computers in Physics, 8(3): 249–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823294.

Holmes:1994:FHP

[862] Lewis M. Holmes. Find hundreds of physics products online. Computers in Physics, 8(3):254–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823295.

Hamlin:1994:HAH

[863] Christopher B. Hamlin and Peggy Judd. How AIP handles electronic submissions of articles. Computers in Physics, 8(3):255–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823296.

Newrock:1994:CRB

[864] Richard Newrock and James R. Matey. Catching the right bus III: IEEE-488.1. Computers in Physics, 8(3):259–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823297.

Thompson:1994:MMC

[865] William J. Thompson. Mathematica and Maple: Has Champaign met its Waterloo? Computers in Physics, 8(3):269–278, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168492.

Ratajack:1994:CST

[866] Mark T. Ratajack, D. O. VanOstenburg, and Denis Donnelly. Computer simulation teaches Fourier transform NMR. Computers in Physics, 8(3): 279–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823298.

Adler:1994:SE

[867] Joan Adler. Series expansions. Computers in Physics, 8(3):287–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168493.

Markel:1994:BRP

[868] Scott Markel, A. John Mallinckrodt, and Susan McKay. Book review: Peter R. Keller and Mary M. Keller, Visual Cues: Practical Data Visualization. Computers in Physics, 8(3):297–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823299.

Mallinckrodt:1994:BRB

[869] A. John Mallinckrodt and Susan McKay. Book review: T. Berners-Lee and Dimitri Dimitroyannis, World Wide Web. Computers in Physics, 8 (3):298–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823300.

Amala:1994:PMTb

[870] Paul A. K. Amala and Garry H. Rodrigue. Programming models for three-dimensional hydrodynamics on the CM-5 (Part II). Computers in Physics, 8(3):300–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823301.

Christensen:1994:RCA

[871] Steven M. Christensen. Resources for computer algebra. Computers in Physics, 8(3):308–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823302.

Blass:1994:ARS

[872] William E. Blass, Wade D. Humbard, Stephen L. Mahan, and Margaret H. Gjertsen. After rough start, Lantastic gives good service in a research laboratory network. Computers in Physics, 8(3):317–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823303.

Anonymous:1994:LLRc

[873] Anonymous. Listings of the latest releases. Computers in Physics, 8(3): 321–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823304.

Hasler:1994:HPI

[874] A. F. Hasler, K. Palaniappan, M. Manyin, and J. Dodge. A high performance Interactive Image SpreadSheet (IISS). Computers in Physics, 8 (3):325–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168495.

Adams:1994:SCE

[875] Barry G. Adams and Gustavo A. Arteca. Symbolic computation of energy perturbation expansions for spherically symmetric hydrogenic systems. Computers in Physics, 8(3):343–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168496.

Onodera:1994:NIR

[876] Y. Onodera. Numerov integration for radial wave equations in cylindrical symmetry. Computers in Physics, 8(3):352–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168497.

Williams:1994:NAC

[877] H. T. Williams and Colin J. Wynne. A new algorithm for computation of SU_{3} Clebsch–Gordon coefficients. Computers in Physics, 8(3):355–??, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168498.

Castellvi:1994:TST

[878] P. Castellví, X. Jaén, and E. Llanta. TTC: Symbolic tensor and exterior calculus. Computers in Physics, 8(3):360–367, May 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168499.

Holmes:1994:JPO

[879] Lewis M. Holmes. Journal publishing: The online challenge. Computers in Physics, 8(4):371–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823305.

Gruber:1994:WCW

[880] Ralf Gruber. What can we learn from Physics Computing ’94? Computers in Physics, 8(4):373–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823306.

Lewin:1994:MFI

[881] David I. Lewin. Meeting future Internet needs. Computers in Physics, 8 (4):374–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823307.

Zeyher:1994:GCM

[882] Allen Zeyher. Gravitational chaos may suggest cosmic structure formation. Computers in Physics, 8(4):377–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823308.

Gruber:1994:PEF

[883] Ralf Gruber. PC ’94 Exhibition features SwissHPC. Computers in Physics, 8(4):379–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823309.

Parisi:1994:RRC

[884] Franklin J. Parisi. A rose is a rose… but a Cray’s not always a Cray. Computers in Physics, 8(4):380–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823310.

Anonymous:1994:Cd

[885] Anonymous. Calendar. Computers in Physics, 8(4):381–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823311.

MacDonald:1994:AIC

[886] William M. MacDonald, Maria Dworzecka, Robert Ehrlich, Jaroslaw Tuszynski, and Glenn Ricart. Authoring on the Internet — the CUPS series. Computers in Physics, 8(4):386–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823312.

Ginsparg:1994:FST

[887] Paul Ginsparg. First steps towards electronic research communication. Computers in Physics, 8(4):390–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823313.

Alexis:1994:NIA

[888] Dominique Alexis. New this issue: Advance program, author guidelines, and 100 more new products. Computers in Physics, 8(4):397–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823314.

Ingoldsby:1994:AAP

[889] Tim Ingoldsby. AIP’s Applied Physics Letters Online: Coming in January. Computers in Physics, 8(4):398–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823315.

Riedel:1994:FEU

[890] K. S. Riedel, A. Sidorenko, and James R. Matey. Function estimation using data-adaptive kernel smoothers — how much smoothing? Computers in Physics, 8(4):402–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823316.

Butcher:1994:RKMa

[891] J. C. Butcher. Runge–Kutta methods in modern computation, Part I: Fundamental concepts. Computers in Physics, 8(4):411–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168500.

Douglas:1994:PPA

[892] Solomon Castillo Douglas, David M. Harrison, Theodore G. Shepherd, and Denis Donnelly. The physical pendulum in an advanced undergraduate course in mechanics. Computers in Physics, 8(4):416–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823317.

Sander:1994:FFC

[893] Evelyn Sander, Leonard M. Sander, and Robert M. Ziff. Fractals and fractal correlations. Computers in Physics, 8(4):420–??, July 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168501.

Christian:1994:BRJ

[894] Wolfgang Christian, G. Andrew Antonelli, A. John Mallinckrodt, and Susan McKay. Book review: James M. Feagin, Quantum Methods with Mathematica(R). Computers in Physics, 8(4):427–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823318.

Ulberg:1994:BRC

[895] Dima E. Ulberg, A. John Mallinckrodt, and Susan McKay. Book review: Charles H. Koelbel, David B. Loveman, Robert S. Schreiber, Guy L. Steele, Jr., and Mary E. Zosel, The High Performance Fortran Handbook. Computers in Physics, 8(4):428–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823319.

Peskin:1994:SCI

[896] Richard L. Peskin, Sandra S. Walther, and Paul F. Dubois. Scene — a computational interface system. Computers in Physics, 8(4):430–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823320.

Blass:1994:ISM

[897] William E. Blass, Stephen L. Mahan, Gordon Chin, and Margaret H. Gjertsen. IDL shines on the Macintosh. Computers in Physics, 8(4): 439–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823321.

Anonymous:1994:LLRd

[898] Anonymous. Listings of the latest releases. Computers in Physics, 8(4): 442–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823322.

Biswas:1994:NES

[899] Tarun Biswas. Nonlinear effects in a stretched string — a numerical computation. Computers in Physics, 8(4):446–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168502.

Casados-Cruz:1994:CPC

[900] Gaspar Casados-Cruz and Arturo Morales-Acevedo. A computer program for the characterization of solar cells. Computers in Physics, 8(4): 451–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168503.

Fairen:1994:NTS

[901] Victor Fairén, Pablo Martin, and José Manuel Ferrándiz. Numerical tracking of small deviations from analytically known periodic orbits. Computers in Physics, 8(4):455–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168504.

Atkinson:1994:TMI

[902] J. Dudley Atkinson and David D. Cleary. A technique for modeling interference patterns from two-beam interferometers. Computers in Physics, 8 (4):462–??, July 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168505.

Houtman:1994:LPE

[903] H. Houtman, F. W. Jones, and C. J. Kost. Laplace and Poisson equation solution by relax3d. Computers in Physics, 8(4):469–??, July 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168506.

Roman:1994:IWM

[904] Thomas A. Roman. The inflating wormhole: a Mathematica animation. Computers in Physics, 8(4):480–??, July 1994. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168507.

Holmes:1994:CM

[905] Lewis M. Holmes. Consider metacomputing. Computers in Physics, 8(5):493–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823323.

Holmes:1994:LRH

[906] Lewis M. Holmes. Let the readers be heard! Computers in Physics, 8(5):495–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823324.

Lewin:1994:NDC

[907] David I. Lewin. NCSA’s director champions broad access to supercomputers. Computers in Physics, 8(5):496–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823325.

Chinnock:1994:HCL

[908] Chris Chinnock. Here come the LCD panels and projectors. Computers in Physics, 8(5):499–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823326.

Anonymous:1994:Ce

[909] Anonymous. Calendar. Computers in Physics, 8(5):504–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823327.

Ricart:1994:FI

[910] Glenn Ricart. Free the Internet! Computers in Physics, 8(5):508–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823328.

Alexis:1994:NIW

[911] Dominique Alexis. New this issue: Web link to CIP online & reader survey report. Computers in Physics, 8(5):511–??, September 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823329.

Butcher:1994:RKMb

[912] J. C. Butcher. Runge–Kutta methods in modern computation, Part II: Implicit methods and related applications. Computers in Physics, 8(5): 512–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168522.

Chaudhury:1994:IPE

[913] S. Raj Chaudhury, Dean Zollman, and Denis Donnelly. Image processing enhances the value of digital video in physics instruction. Computers in Physics, 8(5):518–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823330.

Hu:1994:SMC

[914] Yue Hu. Simulations of the motion of colloidal particles in alternating electric fields. Computers in Physics, 8(5):524–??, September 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168508.

Wenes:1994:BRR

[915] Geert Wenes, A. John Mallinckrodt, and Susan McKay. Book review: Richard E. Crandall, Projects in Scientific Computation. Computers in Physics, 8(5):531–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823331.

Friedman:1994:BRS

[916] Mark Friedman, A. John Mallinckrodt, and Susan McKay. Book review: Steven Strogatz, Nonlinear Dynamics and Chaos: With Applications to Physics, Biology, Chemistry, and Engineering. Computers in Physics, 8(5):532–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823332.

Still:1994:PPC

[917] Charles H. Still and Paul F. Dubois. Portable parallel computing via the MPI1 Message-Passing Standard. Computers in Physics, 8(5):533– 536, 538–539, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823333.

Roskies:1994:MPP

[918] Ralph Roskies. Metacomputing — pipedream or practical reality? Computers in Physics, 8(5):540–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823334.

Anonymous:1994:LLRe

[919] Anonymous. Listings of the latest releases. Computers in Physics, 8(5): 546–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823336.

Renz:1994:IVT

[920] Wolfgang Renz, Rüdiger Westermann, and Wolfgang Krüger. Interactive visualization of three-dimensional cellular automata. Computers in Physics, 8(5):550–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168509.

Kristyan:1994:ESM

[921] Sandor Kristyan. Equipotential surfaces of the model reaction H + H_{2}: Struggling with three dimensions. Computers in Physics, 8(5):556–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168510.

Rijnders:1994:DVP

[922] Frank M. Rijnders, H. J. W. Spoelder, and F. C. A. Groen. Distributed visual programming environment: Possibilities and limitations. Computers in Physics, 8(5):576–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168511.

Chang:1994:CFA

[923] Y. M. Chang, C. Rudowicz, and Y. Y. Yeung. Crystal field analysis of the 3d^{N} ions at low symmetry sites including the “imaginary” terms. Computers in Physics, 8(5):583–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168512.

Musgrave:1994:SIR

[924] Peter Musgrave and Kayll Lake. Scalar invariants of the radiating Kerr–Newman metric: a simple application of GRTensor. Computers in Physics, 8(5):589–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168482.

Raedt:1994:AST

[925] Hans De Raedt and Kristel Michielsen. Algorithm to solve the timedependent Schrödinger equation for a charged particle in an inhomogeneous magnetic field: Application to the Aharonov–Bohm effect. Computers in Physics, 8(5):600–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168483.

Bajema:1994:MBM

[926] J. Bajema, K. Ashktorab, F. D. Becchetti, D. Hotz, and M. Shapiro. A Macintosh^{T M}-based multiparameter pulse height analyzer and multichannel scaler system for advanced teaching laboratories. Computers in Physics, 8(5):608–??, September 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168523.

Holmes:1994:MBP

[927] Lewis M. Holmes. Modeling the blue planet. Computers in Physics, 8 (6):621–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823337.

Petti:1994:WMS

[928] Richard Petti. Why math-software development counts. Computers in Physics, 8(6):623–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823338.

Lewin:1994:ACP

[929] David I. Lewin. Animator chooses physics over movies. Computers in Physics, 8(6):624–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823339.

Stauffer:1994:LTT

[930] Dietrich Stauffer. Landslides towards teraflops? A review of PC ’94. Computers in Physics, 8(6):627–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823340.

Zeyher:1994:WMA

[931] Allen Zeyher. Weather models advance by adding physics. Computers in Physics, 8(6):627–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823341.

Holmes:1994:ELV

[932] Lewis M. Holmes. Exhibition links vendors with research. Computers in Physics, 8(6):628–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823342.

Shearin:1994:GIG

[933] Rhonda Shearin. Government and industry get together on the superhighway. Computers in Physics, 8(6):630–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823343.

Ekstrom:1994:REA

[934] Göran Ekström and Glenn Ricart. Rapid earthquake analysis utilizes the Internet. Computers in Physics, 8(6):632–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823344.

Alexis:1994:NAS

[935] Dominique Alexis. New attractions: Spec sheets for new products and Internet goldmines. Computers in Physics, 8(6):640–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823345.

Matey:1994:GHP

[936] James R. Matey. Good housekeeping Part I: Remaining sane in the face of inevitable disasters. Computers in Physics, 8(6):642–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823346.

Craig:1994:PNA

[937] I. J. D. Craig, A. M. Thompson, and William J. Thompson. Practical numerical algorithms: Why Laplace transforms are difficult to invert numerically. Computers in Physics, 8(6):648–??, November 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823347.

McLaughlin:1994:EFD

[938] Keith L. McLaughlin and Steven M. Day. 3D elastic finite-difference seismic-wave simulations. Computers in Physics, 8(6):656–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823348.

Treinish:1994:VID

[939] Lloyd A. Treinish. Visualizations illuminate disparate data sets in the Earth sciences. Computers in Physics, 8(6):664–??, November 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823349.

Donnelly:1994:CFA

[940] Denis Donnelly. CIP’S fifth annual software contest: The winners. Computers in Physics, 8(6):672–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823350.

Goloskie:1994:QMT

[941] R. Goloskie, J. W. Kramer, and L. R. Ram-Mohan. Quantum-mechanical tunneling and finite elements. Computers in Physics, 8(6):679–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168485.

Coppersmith:1994:BRR

[942] Susan Coppersmith, A. John Mallinckrodt, and Susan McKay. Book review: Robert C. Hilborn, Chaos and Nonlinear Dynamics: An Introduction for Scientists and Engineers. Computers in Physics, 8(6):689– ??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823351.

Haney:1994:CFE

[943] Scott W. Haney. Is C++ fast enough for scientific computing? Computers in Physics, 8(6):690–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168486.

Wilkinson:1994:GTI

[944] Ladye Kathryn Wilkinson, John S. Risley, John E. Gastineau, Paula V. Engelhardt, Scott F. Schultz, and Margaret H. Giertsen. Graphs & tracks impresses as a kinematics teaching tool. Computers in Physics, 8(6):696– ??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823352.

Anonymous:1994:LLRf

[945] Anonymous. Listings of the latest releases. Computers in Physics, 8(6): 701–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823353.

Punzo:1994:HRL

[946] Giovanni Punzo, Federico Massaioli, and Sauro Succi. High-resolution lattice–Boltzmann computing on the IBM SP1 scalable parallel computer. Computers in Physics, 8(6):705–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168487.

Germann:1994:DPT

[947] Timothy C. Germann, Dudley R. Herschbach, and Bruce M. Boghosian. Dimensional perturbation theory on the Connection Machine. Computers in Physics, 8(6):712–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168488.

Akyildiz:1994:PCF

[948] Yilmaz Akyildiz. Parametric curve fitting: An alternative to Lagrange interpolation and splines. Computers in Physics, 8(6):722–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168489.

Cole:1994:NES

[949] James B. Cole. A nearly exact second-order finite-difference timedomain wave propagation algorithm on a coarse grid. Computers in Physics, 8(6):730–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168490.

Yu:1994:SSN

[950] K. W. Yu. Symbolic simulation of nonlinear random resistor networks by Mathematica. Computers in Physics, 8(6):735–??, November 1994. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168521.

Holmes:1995:LAL

[951] Lewis M. Holmes. Learning about learning software. Computers in Physics, 9(1):5–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823364.

Sabelli:1995:OCS

[952] Nora H. Sabelli. For our Children’s sake, take full advantage of technology. Computers in Physics, 9(1):7–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823378.

Lewin:1995:MPT

[953] David I. Lewin. Math professor turns classroom program into successful business. Computers in Physics, 9(1):8–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823379.

Holmes:1995:SSH

[954] Lewis M. Holmes. SC ’94 spotlights high-performance computing. Computers in Physics, 9(1):11–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823354.

Chinnock:1995:TBR

[955] Chris Chinnock. Telemedicine is becoming a reality. Computers in Physics, 9(1):13–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823355.

Shearin:1995:FRL

[956] Rhonda Shearin. Forest researchers look for computer solutions. Computers in Physics, 9(1):16–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823356.

Klimenko:1995:RFC

[957] Stanislav V. Klimenko. Russians form center of computing for physics and technology. Computers in Physics, 9(1):16–??, January 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823357.

Shearin:1995:SDP

[958] Rhonda Shearin. Submission deadline for PC ’95 nears. Computers in Physics, 9(1):17–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823358.

Ricart:1995:NM

[959] Glenn Ricart. A night at the movies. Computers in Physics, 9(1):19– ??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823359.

Alexis:1995:CSI

[960] Dominique Alexis. Calendar section and issue previews go on our Web page. Computers in Physics, 9(1):22–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823360.

Anonymous:1995:NP

[961] Anonymous. NASA publications. Computers in Physics, 9(1):23–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823361.

Jones:1995:CRBa

[962] Eric D. Jones, Bryan L. Preppernau, P. G. Stein, and James R. Matey. Catching the right bus, Part IV: IEEE-488 programming. Computers in Physics, 9(1):24–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823362.

Hindmarsh:1995:ASOa

[963] Alan C. Hindmarsh and Linda R. Petzold. Algorithms and software for ordinary differential equations and differential-algebraic equations, Part I: Euler methods and error estimation. Computers in Physics, 9(1): 34–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168536.

Walker:1995:CMC

[964] P. J. Walker, I. D. Johnston, and Denis Donnelly. Computer model clarifies spontaneous charge redistribution in conductors. Computers in Physics, 9(1):42–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823363.

Novotny:1995:NAO

[965] M. A. Novotny. A new approach to an old algorithm for the simulation of Ising-like systems. Computers in Physics, 9(1):46–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168537.

Catlow:1995:NMM

[966] C. R. A. Catlow, A. M. Stonehas, John Meurig Thomas, and Susan McKay. New methods for modelling processes within solids and at their surfaces. Computers in Physics, 9(1):54–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823365.

Kockler:1995:NMS

[967] Norbert Köckler and Susan McKay. Numerical methods and scientific computing: Using software libraries for problem solving. Computers in Physics, 9(1):54–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823366.

Gutfreund:1995:BCP

[968] H. Gutfreund, G. Toulouse, and Susan McKay. Biology and computation: a physicist’s choice. Computers in Physics, 9(1):54–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823367.

Mallinckrodt:1995:BRAa

[969] A. John Mallinckrodt. Book review: Adam Gaffin, Everybody’s Guide to the Internet. Computers in Physics, 9(1):55–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823368.

Mallinckrodt:1995:BRS

[970] A. John Mallinckrodt. Book review: Stephen Wolfram, Cellular Automata and Complexity. Computers in Physics, 9(1):55–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823369.

McKay:1995:BRR

[971] Susan McKay. Book review: Robert Alan Brown, Machines That Learn: Based on the Principles of Empirical Control. Computers in Physics, 9 (1):55–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823370.

Mallinckrodt:1995:BRJ

[972] A. John Mallinckrodt. Book review: Joseph O’Rourke, Computational Geometry in C. Computers in Physics, 9(1):55–??, January 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823371.

Mallinckrodt:1995:BRAb

[973] A. John Mallinckrodt. Book review: Alejandro L. Garcia, Numerical Methods for Physics. Computers in Physics, 9(1):55–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823372.

McKay:1995:BRF

[974] Susan McKay. Book review: F. B. Jensen, W. A. Kuperman, M. B. Porter, and H. Schmidt, Computational Ocean Acoustics. Computers in Physics, 9(1):55–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823373.

McKay:1995:BRA

[975] Susan McKay. Book review: Allan R. Robinson and Ding Lee, Oceanography and Acoustics: Prediction and Propagation Models. Computers in Physics, 9(1):56–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823374.

McAllister:1995:SCG

[976] David F. McAllister and Susan McKay. Stereo computer graphics and other true 3D technologies. Computers in Physics, 9(1):56–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823375.

Alexander:1995:HCX

[977] P. Alexander, L. F. Gladden, and Paul F. Dubois. How to create an X-Window interface to Gnuplot and Fortran programs using the Tcl/ Tk toolkit. Computers in Physics, 9(1):57–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823376.

Wilkinson:1995:DPC

[978] Ladye K. Wilkinson, Joseph Barnes, Margaret H. Gjertsen, John S. Risley, and Margaret H. Gjertsen. 1995 directory of physics courseware. Computers in Physics, 9(1):65–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823377.

Anonymous:1995:LLRa

[979] Anonymous. Listings of the latest releases. Computers in Physics, 9(1): 81–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823380.

Haider:1995:PMC

[980] Niaz Haider, Souheil A. Khaddaj, Mark R. Wilby, and Dimitri D. Vvedensky. Parallel Monte Carlo simulations of epitaxial growth. Computers in Physics, 9(1):85–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168538.

Phillies:1995:WAI

[981] George D. J. Phillies and Jonathan Stott. Wavelet analysis of Ising model spin dynamics. Computers in Physics, 9(1):97–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168539.

Yeh:1995:PDT

[982] Mei-Ling Yeh, Jeongnim Kim, and Furrukh S. Khan. Parallel decomposition of the tight-binding fictitious Lagrangian algorithm for molecular dynamics simulations of semiconductors. Computers in Physics, 9(1): 108–??, January 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168535.

Holmes:1995:VWN

[983] Lewis M. Holmes. Visualize this: We have a new column! Computers in Physics, 9(2):125–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823381.

Sinervo:1995:RIS

[984] Pekka K. Sinervo. Remember that the Internet is a science resource! Computers in Physics, 9(2):127–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823382.

Lewin:1995:MCE

[985] David I. Lewin. Mathematician compiles enviable record in industrial R&D. Computers in Physics, 9(2):128–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823383.

Holmes:1995:TLA

[986] Lewis M. Holmes. Teachers learn about computers in physics education. Computers in Physics, 9(2):131–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823384.

Shearin:1995:EPT

[987] Rhonda Shearin. Educational programs take advantage of computing techniques. Computers in Physics, 9(2):133–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823385.

Ricart:1995:WN

[988] Glenn Ricart. What’s in a name? Computers in Physics, 9(2):135–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823386.

Alexis:1995:MPI

[989] Dominique Alexis. More product information by Telnet and the World Wide Web. Computers in Physics, 9(2):138–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823387.

Jones:1995:CRBb

[990] Eric D. Jones, Bryan L. Preppernau, and James R. Matey. Catching the right bus V: a beginners’ guide to programming the IEEE-488 bus. Computers in Physics, 9(2):140–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823388.

Hindmarsh:1995:ASOb

[991] Alan C. Hindmarsh and Linda R. Petzold. Algorithms and software for ordinary differential equations and differential-algebraic equations, Part II: Higher-order methods and software packages. Computers in Physics, 9(2):148–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168540.

Winograd:1995:HUC

[992] Joseph M. Winograd, James H. McClellan, and Denis Donnelly. How to use a computer-algebra system for reconstruction of functions from parallel-line projections. Computers in Physics, 9(2):156–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823389.

Grinstein:1995:SMS

[993] G. Grinstein and C. Jayaprakash. Simple models of self-organized criticality. Computers in Physics, 9(2):164–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168541.

Stanley:1995:BRA

[994] H. Eugene Stanley, A. John Mallinckrodt, and Susan McKay. Book review: Alexander Yu. Grosberg and Alexei R. Khokhlov, Statistical Physics of Macromolecules. Computers in Physics, 9(2):171–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823390.

MacKay:1995:BRB

[995] David J. C. MacKay, A. John Mallinckrodt, and Susan McKay. Book review: Benjamin Kuipers, Qualitative Reasoning: Modeling and Simulation with Incomplete Knowledge. Computers in Physics, 9(2):172–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823391.

Brown:1995:CUP

[996] Stewart A. Brown, Paul F. Dubois, and David H. Munro. Creating and using PDB files. Computers in Physics, 9(2):173–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823392.

Bryson:1995:HCW

[997] Steve Bryson. How can we make visualization more useful to the working scientist? Computers in Physics, 9(2):178–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823393.

Wilkinson:1995:BGM

[998] Ladye K. Wilkinson, John E. Gastineau, Margaret H. Gjertsen, Margaret H. Gjertsen, and John S. Risley. A buyers’ guide to microcomputerbased laboratory equipment in physics education. Computers in Physics, 9(2):185–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823394.

Anonymous:1995:LLRb

[999] Anonymous. Listings of the latest releases. Computers in Physics, 9(2): 201–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823395.

Bunge:1995:MCM

[1000] Hans-Peter Bunge and John R. Baumgardner. Mantle convection modeling on parallel virtual machines. Computers in Physics, 9(2):207–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168525.

Haney:1995:FUF

[1001] S. W. Haney, J. P. Freidberg, and C. J. Solomon. A fast, user-friendly code for calculating magnetohydrodynamic equilibria. Computers in Physics, 9(2):216–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168526.

Phillies:1995:MZD

[1002] George D. J. Phillies and Jonathan Stott. Mori–Zwanzig–Daubechies decomposition of Ising-model Monte Carlo dynamics. Computers in Physics, 9(2):225–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168527.

Cole:1995:FDT

[1003] James B. Cole, Rudolph A. Krutar, Susan K. Numrich, and Dennis B. Creamer. Finite-difference time-domain simulations of wave propagation and scattering as a research and educational tool. Computers in Physics, 9(2):235–??, March 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168528.

Holmes:1995:CDP

[1004] Lewis M. Holmes. Be careful with that downloaded program! Computers in Physics, 9(3):245–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823396.

Creutz:1995:HCP

[1005] Michael Creutz. How computational physics is uniting science and revolutionizing society. Computers in Physics, 9(3):247–??, May 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823397.

Lewin:1995:CSA

[1006] David I. Lewin. Computer scientist addresses grand themes. Computers in Physics, 9(3):248–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823398.

Carts-Powell:1995:DDS

[1007] Yvonne Carts-Powell. Dealing with the devil: SATAN is loose on the Internet. Computers in Physics, 9(3):251–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823399.

Ricart:1995:PPP

[1008] Glenn Ricart. Prevention and prudent precautions: Computing security on the Internet. Computers in Physics, 9(3):254–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823400.

Alexis:1995:CWS

[1009] Dominique Alexis. CIP’s Web site now lets you search interactively. Computers in Physics, 9(3):259–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823401.

Anonymous:1995:IGC

[1010] Anonymous. Internet goldmine: Computational science highlights. Computers in Physics, 9(3):260–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823402.

Redman:1995:SOS

[1011] James M. Redman, Richard R. Silbar, and James R. Matey. Selecting an operating system. Part II: Nextstep. Computers in Physics, 9(3): 261–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823403.

Sidhu:1995:EIF

[1012] Satinder S. Sidhu. Elliptic integrals and functions. Computers in Physics, 9(3):268–276, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168529.

Byrne:1995:UDC

[1013] John Byrne, Eric Kincanon, and Denis Donnelly. Undergraduates do chaotic research. Computers in Physics, 9(3):277–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823404.

Bell:1995:NSR

[1014] Stephen C. Bell. A numerical solution of the relativistic Kepler problem. Computers in Physics, 9(3):281–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168530.

Matey:1995:BRR

[1015] James Matey, A. John Mallinckrodt, and Susan McKay. Book review: Ron Ledesma, PC Hardware Configuration Guide for DOS and Solaris. Computers in Physics, 9(3):287–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823405.

Anonymous:1995:PC

[1016] Anonymous. Physics Computing ’95. Computers in Physics, 9(3):289–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823406.

Boehm:1995:DMA

[1017] Hans-J. Boehm and Paul F. Dubois. Dynamic memory allocation and garbage collection. Computers in Physics, 9(3):297–??, May 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823407.

Hesselink:1995:TSO

[1018] Lambertus Hesselink, Thierry Delmarcelle, James L. Helman, and Steve Bryson. Topology of second-order tensor fields. Computers in Physics, 9 (3):304–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823408.

Gastineau:1995:QMS

[1019] John E. Gastineau and Margaret H. Gjertsen. Quantum mechanics simulations excels as an educational tool. Computers in Physics, 9(3):312–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4823409.

Anonymous:1995:LLRc

[1020] Anonymous. Listings of the latest releases. Computers in Physics, 9(3): 314–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823410.

deIzarra:1995:CWF

[1021] C. de Izarra, O. Vallée, J. Picart, and N. Tran Minh. Computation of the Whittaker functions W_{k,μ}(z) with series expansions and Padé approximants. Computers in Physics, 9(3):318–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168531.

Boyd:1995:MPP

[1022] John P. Boyd. Multiple precision pseudospectral computations of the radiation coefficient for weakly nonlocal solitary waves: Fifth-order Korteweg–de Vries equation. Computers in Physics, 9(3):324–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168557.

Castellvi:1995:STC

[1023] P. Castellví, X. Jaén, and E. Llanta. Symbolic tensor calculus using index notation. Computers in Physics, 9(3):335–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168532.

Penna:1995:FCS

[1024] T. J. P. Penna. Fitting curves by simulated annealing. Computers in Physics, 9(3):341–??, May 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168533.

Holmes:1995:WNS

[1025] Lewis M. Holmes. What’s new in scientific publishing? The Web, of course! Computers in Physics, 9(4):349–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823411.

Ricart:1995:SPF

[1026] Glenn Ricart. Scholarly publishing faces cultural challenge. Computers in Physics, 9(4):351–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823412.

Lewin:1995:BAI

[1027] David I. Lewin. Born-again Internet gadfly takes on computational physics. Computers in Physics, 9(4):352–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823413.

Chinnock:1995:SRF

[1028] Chris Chinnock. Speech recognition is finding its way into personal computing. Computers in Physics, 9(4):355–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823414.

Ricart:1995:SIW

[1029] Glenn Ricart. The scholarly information Web. Computers in Physics, 9 (4):360–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823415.

Carts-Powell:1995:SPE

[1030] Yvonne Carts-Powell. Software products enrich Web options for scientists. Computers in Physics, 9(4):364–??, July 1995. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823416.

Judd:1995:SEF

[1031] Peggy Judd and F. Scott Johnson. SGML enables full-text scientific publishing on the Web. Computers in Physics, 9(4):369–370, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823417.

Alexis:1995:CHP

[1032] Dominique Alexis. CIP’s home page gets facelift. Computers in Physics, 9 (4):371–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823418.

Bracewell:1995:CHT

[1033] Ronald N. Bracewell. Computing with the Hartley transform. Computers in Physics, 9(4):373–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168534.

Ravaglia:1995:CBI

[1034] Raymond Ravaglia, J. Acacio de Barros, Patrick Suppes, and Denis Donnelly. Computer-based instruction brings advanced-placement physics to gifted students. Computers in Physics, 9(4):380–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823419.

Kolan:1995:ELN

[1035] Amy Kolan, Barry Cipra, and Bill Titus. Exploring localization in nonperiodic systems. Computers in Physics, 9(4):387–??, July 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168542.

Mallinckrodt:1995:BRB

[1036] A. John Mallinckrodt and Susan McKay. Book review: Bruce Hannon and Matthias Ruth, Dynamic Modeling. Computers in Physics, 9(4): 397–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823420.

Mallinckrodt:1995:BRK

[1037] A. John Mallinckrodt. Book review: Klaus Mainzer, Thinking in Complexity: The Complex Dynamics of Matter, Mind, and Mankind. Computers in Physics, 9(4):398–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823421.

Koniges:1995:PCR

[1038] Alice E. Koniges, Kevin R. Lind, and Paul Dubois. Parallelizing code for real applications on the T3D. Computers in Physics, 9(4):399–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4823422.

Miceli:1995:AAM

[1039] Kristina D. Miceli and Steve Bryson. An automated approach to multidisciplinary visualization. Computers in Physics, 9(4):406–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823423.

Blondin:1995:CAS

[1040] John M. Blondin and Margaret H. Gjertsen. CUPS Astrophysics Simulations will fit well in upper-level courses. Computers in Physics, 9(4): 413–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823424.

Anonymous:1995:LLRd

[1041] Anonymous. Listings of the latest releases. Computers in Physics, 9(4): 416–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823425.

Lyster:1995:ICT

[1042] P. M. Lyster, P. C. Liewer, V. K. Decyk, and R. D. Ferraro. Implementation and characterization of three-dimensional particle-in-cell codes on multiple-instruction-multiple-data massively parallel supercomputers. Computers in Physics, 9(4):420–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168543.

Zheng:1995:SSV

[1043] Yibin Zheng and Peter C. Doerschuk. Symbolic symmetry verification for harmonic functions invariant under polyhedral symmetries. Computers in Physics, 9(4):433–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168544.

Petridis:1995:IPT

[1044] Athanasios N. Petridis and Fred K. Wohn. Implementation of PHENIX trigger algorithms on massively parallel computers. Computers in Physics, 9(4):438–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168545.

Melendez:1995:SQM

[1045] W. Melendez and C. J. Horowitz. Semiclassical quark model of the nuclear matter response. Computers in Physics, 9(4):450–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168546.

Hollerbach:1995:FDA

[1046] Rainer Hollerbach. Fast dynamo action in spherical geometry: Numerical calculations using parallel virtual machines. Computers in Physics, 9(4): 460–??, July 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168547.

Holmes:1995:WMW

[1047] Lewis M. Holmes. Weather makes waves. . .. Computers in Physics, 9(5):469–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823426.

Preppernau:1995:KOS

[1048] Bryan L. Preppernau. Keeping our scientific voice in the computer industry. Computers in Physics, 9(5):471–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823427.

Lewin:1995:POR

[1049] David I. Lewin. Physicist oversees research for software giant. Computers in Physics, 9(5):472–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823428.

Daukantas:1995:SHW

[1050] Patricia Daukantas. SC ’95 highlights Web, applications. Computers in Physics, 9(5):475–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823429.

Williams:1995:SWC

[1051] John Michael Williams. Should we be collecting garbage? Computers in Physics, 9(5):479–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823430.

Alexis:1995:MOC

[1052] Dominique Alexis. More offered from CIP’s home page. Computers in Physics, 9(5):481–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823431.

Anonymous:1995:MOC

[1053] Anonymous. More offered from CIP’s home page. Computers in Physics, 9(5):481–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4825203.

Hunt:1995:DFA

[1054] J. C. R. Hunt. Developments in forecasting the atmospheric environment. Computers in Physics, 9(5):483–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823432.

Kalnay:1995:NWP

[1055] Eugenia Kalnay. Numerical weather prediction. Computers in Physics, 9(5):488–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823433.

Forrest:1995:CRB

[1056] Gary Forrest, Scott McKinney, and James R. Matey. Catching the right bus, Part VI: Some options for real-time image acquisition. Computers in Physics, 9(5):496–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823434.

Vattulainen:1995:MIF

[1057] I. Vattulainen and T. Ala-Nissila. Mission impossible: Find a random pseudorandom number generator. Computers in Physics, 9(5):500–504, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168548.

Caporaloni:1995:SPC

[1058] Marina Caporaloni, Roberto Ambrosini, and Denis Donnelly. Student project checks Leonardo Da Vinci’s map work. Computers in Physics, 9(5):505–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4823435.

Yalcinkaya:1995:CS

[1059] Tolga Yalçinkaya and Ying-Cheng Lai. Chaotic scattering. Computers in Physics, 9(5):511–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168549.

Stoutamire:1995:SRH

[1060] David Stoutamire, Matthew Kennel, and Paul F. Dubois. Sather revisited: a high-performance free alternative to C++. Computers in Physics, 9(5):519–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823436.

Bryson:1995:DGR

[1061] Steve Bryson. The data glut revisited. Computers in Physics, 9(5):525– ??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823437.

Gastineau:1995:TRO

[1062] John E. Gastineau and Margaret H. Gjertsen. Three ray-optics programs support varied instructional uses. Computers in Physics, 9(5):531–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823438.

Anonymous:1995:LLRe

[1063] Anonymous. Listings of the latest releases. Computers in Physics, 9(5): 535–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823439.

Moriarty:1995:OOM

[1064] K. J. M. Moriarty, C. Margio, and T. Trappenberg. Object-oriented Monte Carlo simulations. Computers in Physics, 9(5):540–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168550.

Macdonald:1995:SID

[1065] J. Ross Macdonald. Solution of an “impossible” diffusion–inversion problem. Computers in Physics, 9(5):546–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168551.

Rouhi:1995:SIN

[1066] Ali Rouhi and Jon Wright. Spectral implementation of a new operator splitting method for solving partial differential equations. Computers in Physics, 9(5):554–??, September 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168552.

Holmes:1995:SCW

[1067] Lewis M. Holmes. Software contestants win cash prizes. Computers in Physics, 9(6):569–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823440.

Sipson:1995:CBN

[1068] Roger Sipson and Ronald Thornton. Computers bring new opportunity to science education. Computers in Physics, 9(6):571–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823441.

Lewin:1995:NLE

[1069] David I. Lewin. Nobel Laureate espouses physics, education, and breadth of knowledge. Computers in Physics, 9(6):572–??, November 1995. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4823442.

Daukantas:1995:RSH

[1070] Patricia Daukantas. ‘Reversed shear’ heats up plasma research. Computers in Physics, 9(6):575–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823443.

Schoch:1995:CWH

[1071] Peter K. Schoch. Chemistry workshop helps physics teachers. Computers in Physics, 9(6):579–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823444.

Ricart:1995:UI

[1072] Glenn Ricart. Using up the Internet. Computers in Physics, 9(6):581– ??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823445.

Alexis:1995:WSU

[1073] Dominique Alexis. Web site update. Computers in Physics, 9(6):583– ??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4823446.

Anonymous:1995:IGV

[1074] Anonymous. Internet goldmine: Virtual trips to black holes and neutron stars. Computers in Physics, 9(6):583–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823447.

Scales:1995:SOS

[1075] John A. Scales, Martin L. Smith, Karsten Ballüder, and James R. Matey. Selecting an operating system, Part III: Unix in a laboratory environment. Computers in Physics, 9(6):584–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823448.

Pang:1995:SFD

[1076] Alex Pang and Chueng-Ryong Ji. Symbolic Feynman-diagram calculation. Computers in Physics, 9(6):589–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168553.

Donnelly:1995:CSA

[1077] Denis Donnelly. CIP’s sixth annual software contest: The winners. Computers in Physics, 9(6):594–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823449.

Pang:1995:NMQ

[1078] Tao Pang. A numerical method for quantum tunneling. Computers in Physics, 9(6):602–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168554.

Novotny:1995:BRA

[1079] Mark Novotny, Susan McKay, and Wolfgang Christian. Book review: Al Geist, Adam Beguelin, Jack Dongarra, Weicheng Jiang, Robert Manchek, and Vaidy Sunderam, PVM — Parallel Virtual Machine: a Users’ Guide and Tutorial for Networked Parallel Computing. Computers in Physics, 9(6):607–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4823450.

Munro:1995:UYI

[1080] David H. Munro and Paul F. Dubois. Using the Yorick interpreted language. Computers in Physics, 9(6):609–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823451.

Bryson:1995:EVR

[1081] Steve Bryson. Exploratory visualization I: Real-time interaction. Computers in Physics, 9(6):616–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4823452.

Thompson:1995:MCP

[1082] William J. Thompson. MAC clones: Power computing for frugal physicists. Computers in Physics, 9(6):622–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4823453.

Anonymous:1995:LLRf

[1083] Anonymous. Listings of the latest releases. Computers in Physics, 9(6): 625–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4823454.

Holmes:1995:SNC

[1084] Mark H. Holmes, Yuklun Au, and J. W. Stayman. Symbolic and numerical computing for chemical kinetic reaction schemes. Computers in Physics, 9(6):629–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168555.

Rowe:1995:DWM

[1085] Alistair C. H. Rowe and Paul C. Abbott. Daubechies wavelets and Mathematica. Computers in Physics, 9(6):635–648, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168556.

Zuev:1995:CCL

[1086] V. V. Zuev, A. A. Mitsel’, M. Yu. Kataev, I. V. Ptashnik, and K. M. Firsov. Computer code long path method for long path differentialabsorption predictions using CO_{2} laser lines. Computers in Physics, 9 (6):649–??, November 1995. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168558.

Holmes:1996:LMW

[1087] Lewis M. Holmes. Linux makes waves. Computers in Physics, 10(1): 3–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822352.

Browne:1996:TSO

[1088] J. C. Browne. Is traditional science obsolete as an undergraduate major? Computers in Physics, 10(1):5–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822357.

Lewin:1996:HPH

[1089] David I. Lewin. How a physicist helped to launch a revolution. Computers in Physics, 10(1):6–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822359.

Daukantas:1996:NSP

[1090] Patricia Daukantas. NSF seeks partnerships to replace supercomputer centers. Computers in Physics, 10(1):9–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822361.

Daukantas:1996:SNG

[1091] Patricia Daukantas. Simulating the next great crash-on Earth. Computers in Physics, 10(1):10–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822347.

Ricart:1996:IMC

[1092] Glenn Ricart. The Internet moves out of its corner. Computers in Physics, 10(1):14–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822348.

Alexis:1996:HWY

[1093] Dominique Alexis. Here’s what you can find on our Web page. Computers in Physics, 10(1):16–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822349.

Anonymous:1996:WPL

[1094] Anonymous. The WIS Plasma Laboratory. Computers in Physics, 10(1): 16–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822350.

Balluder:1996:SOS

[1095] Karsten Ballüder, John A. Scales, Claus Schroeter, Martin L. Smith, and James R. Matey. Selecting an operating system, Part IV: Linux. Computers in Physics, 10(1):17–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822351.

Thompson:1996:RAM

[1096] William J. Thompson. Recursive algorithms: a mixed blessing. Computers in Physics, 10(1):25–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168559.

Grayson:1996:UER

[1097] Diane J. Grayson and Denis Donnelly. Using education research to develop waves courseware. Computers in Physics, 10(1):30–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822353.

Regez:1996:HCI

[1098] Nicolas Regez, Wolfgang Breymann, Stefan Weigert, Charles Kaufman, and Gerhard Müller. Hamiltonian chaos IV. Computers in Physics, 10(1): 39–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168560.

Davenport:1996:BRF

[1099] Forest Davenport, Susan McKay, and Wolfgang Christian. Book review: Franz J. Vesely, Computational Physics: An Introduction. Computers in Physics, 10(1):47–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822354.

Enzweiler:1996:BRJ

[1100] Raymond N. Enzweiler, Susan McKay, and Wolfgang Christian. Book reviews: J. F. James, A Student’s Guide to Fourier Transforms with Applications in Physics and Engineering. Computers in Physics, 10(1): 47–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822355.

Templon:1996:EPL

[1101] Jeffrey A. Templon and Paul F. Dubois. Evaluation of PC/Linux systems for use as scientific workstations. Computers in Physics, 10(1): 49–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822356.

Bryson:1996:WNV

[1102] Steve Bryson. What’s new at Visualization ’95. Computers in Physics, 10(1):56–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822358.

Anonymous:1996:LLRa

[1103] Anonymous. Listings of the latest releases. Computers in Physics, 10(1): 64–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822360.

Wang:1996:VAO

[1104] Jing B. Wang, Paul C. Abbott, and Jim F. Williams. Visualizing atomic orbitals. Computers in Physics, 10(1):69–82, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168561.

Fernandez:1996:FAR

[1105] Julio F. Fernández and Juan Rivero. Fast algorithms for random numbers with exponential and normal distributions. Computers in Physics, 10(1):

83–88, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.

168562. See [1527] for a proof of the algorithm proposed here for exponential random numbers.

Bao:1996:USM

[1106] Min-Qi Bao and Anthony F. Starace. Using symbolic manipulation to evaluate Feynman’s path integral for an interaction of the form F (t) · r. Computers in Physics, 10(1):89–??, January 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168563.

Holmes:1996:BI

[1107] Lewis M. Holmes. Beyond InfiniteReality? Computers in Physics, 10(2): 99–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822381.

Thompson:1996:ESW

[1108] Joe F. Thompson. The experts have spoken. Who will listen? Computers in Physics, 10(2):101–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822362.

Lewin:1996:CG

[1109] David I. Lewin. Computing the glueball. Computers in Physics, 10(2): 102–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822363.

Daukantas:1996:TGC

[1110] Patricia Daukantas. Teachers gambol on computers in Reno. Computers in Physics, 10(2):105–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822364.

Ricart:1996:PTS

[1111] Glenn Ricart. Plug-in, tune-in, sleep-in. Computers in Physics, 10(2): 110–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822365.

Anonymous:1996:IIR

[1112] Anonymous. Introducing interactive reader service. Computers in Physics, 10(2):113–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822366.

Anonymous:1996:CM

[1113] Anonymous. Chaos at Maryland. Computers in Physics, 10(2):113–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822367.

Wagner:1996:HIC

[1114] John S. Wagner, Michael W. Trahan, William E. Nelson, Gary C. Tisone, and Bryan L. Preppernau. How intelligent chemical recognition benefits from multivariate analysis and genetic optimization. Computers in Physics, 10(2):114–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822368.

Keister:1996:MQA

[1115] Bradley D. Keister. Multidimensional quadrature algorithms. Computers in Physics, 10(2):119–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168565.

Fischer:1996:MDI

[1116] Susan L. Fischer, Wolfgang Christian, and Denis Donnelly. Molecular dynamics inside ion traps. Computers in Physics, 10(2):123–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822369.

Samuels:1996:PSN

[1117] David C. Samuels, H. G. E. Hentschel, and Alan Fine. Parallel simulations of neuron growth. Computers in Physics, 10(2):129–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168564.

Garcia:1996:BRWa

[1118] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: W. S. Brainerd and C. H. Goldberg and J. C. Adams, Programmer’s Guide to Fortran 90. Computers in Physics, 10(2):135–??, March 1996.

CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822370.

Garcia:1996:BRCa

[1119] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: I. Chivers and J. Sleightholme, Introducing Fortran 90. Computers in Physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822371.

Garcia:1996:BRM

[1120] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: M. Metcalf and J. Reid, Fortran 90 Explained. Computers in Physics, 10(2):135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822372.

Garcia:1996:BRJ

[1121] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: J. Kerrigan, Migrating to Fortran 90. Computers in Physics, 10(2): 135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822373.

Garcia:1996:BRCb

[1122] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: C. Redwine, Upgrading to Fortran 90. Computers in Physics, 10(2): 135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822374.

Garcia:1996:BRWb

[1123] Alejandro Garcia, Susan McKay, and Wolfgang Christian. Book review: W. Gehrke, Fortran 90 Language Guide. Computers in Physics, 10(2): 135–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822375.

Intrator:1996:BRM

[1124] Nathan Intrator, Susan McKay, and Wolfgang Christian. Book review: Mohamad H. Hassoun, Fundamentals of Artificial Neural Networks. Computers in Physics, 10(2):137–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822376.

Cohen:1996:CSN

[1125] Scott D. Cohen, Alan C. Hindmarsh, and Paul F. Dubois. CVODE, a stiff/nonstiff ODE solver in C. Computers in Physics, 10(2):138–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822377.

Bryson:1996:IYT

[1126] Steve Bryson. If you take visualization for granted, it may go away. Computers in Physics, 10(2):144–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822378.

Riedel:1996:ODG

[1127] Michael R. Riedel. Origin 4.0 delivers graphing functionality. Computers in Physics, 10(2):147–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822379.

Anonymous:1996:LLRb

[1128] Anonymous. Listings of the latest releases. Computers in Physics, 10(2): 150–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822380.

Qiu:1996:VAQ

[1129] Shi-Yue Qiu, Corbett R. Rowell, and Marko Vukobrat Jarić. Visualization and analysis of quasicrystal densities. Computers in Physics, 10(2): 154–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168566.

Heddle:1996:HGU

[1130] D. P. Heddle. Hv: a graphical user interface library for scientific and engineering applications. Computers in Physics, 10(2):174–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168567.

Camarda:1996:DEQ

[1131] H. S. Camarda. Determining the eigenvalues of a quaternion matrix with a band structure. Computers in Physics, 10(2):180–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168568.

Hessler:1996:NSC

[1132] Jan P. Hessler, David H. Current, and Paul J. Ogren. A new scheme for calculating weights and describing correlations in nonlinear leastsquares fits. Computers in Physics, 10(2):186–??, March 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168569.

Holmes:1996:CMC

[1133] Lewis M. Holmes. Computer modeling that could make us healthier. Computers in Physics, 10(3):203–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822382.

Rigden:1996:DEL

[1134] John S. Rigden. The disciplines educate & liberate. Computers in Physics, 10(3):205–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822383.

Lewin:1996:BHR

[1135] David I. Lewin. Balancing her roles in engineering and society. Computers in Physics, 10(3):206–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822384.

Daukantas:1996:OAH

[1136] Patricia Daukantas. Olympic athletes hone their performance with the help of computers. Computers in Physics, 10(3):209–??, May 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822385.

Sherwood:1996:CR

[1137] Bruce Sherwood. Clarifying recursion. Computers in Physics, 10(3): 215–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822386.

Boyce:1996:BPR

[1138] Peter B. Boyce. Building a peer-reviewed scientific journal on the Internet. Computers in Physics, 10(3):216–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822387.

Anonymous:1996:CLR

[1139] Anonymous. CIP links readers to featured Web sites. Computers in Physics, 10(3):222–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822388.

Anonymous:1996:ASB

[1140] Anonymous. Astrophysics at Santa Barbara. Computers in Physics, 10 (3):222–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822389.

Taylor:1996:CIV

[1141] Charles A. Taylor, Thomas J. R. Hughes, and Christopher K. Zarins. Computational investigations in vascular disease. Computers in Physics, 10(3):224–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822390.

Matey:1996:SDA

[1142] James R. Matey. Software designs and architectures, Part I: Introduction and user input/output. Computers in Physics, 10(3):233–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822391.

Booten:1996:CLSa

[1143] Albert Booten and Henk van der Vorst. Cracking large-scale eigenvalue problems. Part I. Algorithms. Computers in Physics, 10(3):239–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168574.

Phillies:1996:WNA

[1144] George D. J. Phillies. Wavelets: a new alternative to Fourier transforms. Computers in Physics, 10(3):247–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168573.

Butler:1996:BRR

[1145] Malcolm Butler, Susan R. McKay, and Wolfgang Christian. Book review: Roberta Bigelow, John Philpott, and Joseph Rothberg, Nuclear and Particle Physics Simulations. Computers in Physics, 10(3):255–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822392.

Boekema:1996:BRG

[1146] Carolus Boekema, Susan R. McKay, and Wolfgang Christian. Book review: G. Andrew Antonelli, Wolfgang Christian, Susan Fischer, Robin Giles, and Brian James, Waves and Optics Simulations. Computers in Physics, 10(3):256–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822393.

Snowman:1996:BRD

[1147] Daniel P. Snowman, Susan R. McKay, and Wolfgang Christian. Book review: Douglas Brandt, John Hiller, and Michael Moloney, Modern Physics Simulations. Computers in Physics, 10(3):256–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822394.

Hellman:1996:BRR

[1148] Frances Hellman, Pat Diamond, Susan R. McKay, and Wolfgang Christian. Book review: Robert Ehrlich, Lyle Roelofs, Ronald Stoner, and Jaroslaw Tuszynski, Electricity and Magnetism Simulations. Computers in Physics, 10(3):257–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822395.

Schmittmann:1996:BRH

[1149] Beate Schmittmann, Susan R. McKay, and Wolfgang Christian. Book review: Harvey Gould, Lynna Spomick, and Jan Tobochnik, Thermal and Statistical Physics Simulations. Computers in Physics, 10(3):258–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822396.

Morrow:1996:BRB

[1150] Richard A. Morrow, Susan R. McKay, and Wolfgang Christian. Book review: Bruce Hawkins and Randall S. Jones, Classical Mechanics Simulations. Computers in Physics, 10(3):259–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822397.

Syphers:1996:BRJ

[1151] Dale Syphers, Susan R. McKay, and Wolfgang Christian. Book review: J. R. Hiller, I. D. Johnston, and D. F. Styer, Quantum Mechanics Simulations. Computers in Physics, 10(3):260–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822398.

Levey:1996:BRG

[1152] Christopher Levey, Susan R. McKay, and Wolfgang Christian. Book review: Graham Keeler, Roger Rollins, and Steven Spicklemire, Solid State Physics Simulations. Computers in Physics, 10(3):260–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822399.

Dubois:1996:NP

[1153] Paul F. Dubois, Konrad Hinsen, and James Hugunin. Numerical Python. Computers in Physics, 10(3):262–267, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822400.

Rogowitz:1996:HLV

[1154] Bernice E. Rogowitz, Lloyd A. Treinish, and Steve Bryson. How not to lie with visualization. Computers in Physics, 10(3):268–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822401.

Preppernau:1996:FPP

[1155] Bryan L. Preppernau and Isaac Shokair. Fortran Powerstation 4.0 ports Fortran code smoothly to the PC. Computers in Physics, 10(3):275–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822402.

Anonymous:1996:LLRc

[1156] Anonymous. Listings of the latest releases. Computers in Physics, 10(3): 279–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822403.

Yu:1996:VSE

[1157] K. W. Yu. Variational solution of electrostatic boundary-value problems of nonlinear media. Computers in Physics, 10(3):284–??, May 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168570.

Decyk:1996:OPC

[1158] Viktor K. Decyk, Steve Roy Karmesin, Aeint de Boer, and Paulett C. Liewer. Optimization of particle-in-cell codes on reduced instruction set computer processors. Computers in Physics, 10(3):290–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168571.

Tyler:1996:GFC

[1159] David W. Tyler. A Green’s function for the calculation of density response in rare-gas volumetric thermal gratings. Computers in Physics, 10 (3):299–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168572.

Holmes:1996:WWD

[1160] Lewis M. Holmes. Why we do not live forever. Computers in Physics, 10(4):307–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822404.

Worlton:1996:CPT

[1161] Jack Worlton. Computational physics then and now: a cautionary view. Computers in Physics, 10(4):309–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822405.

Lewin:1996:SHS

[1162] David I. Lewin. Setting his sights on petaflops. Computers in Physics, 10(4):310–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822406.

Daukantas:1996:VPW

[1163] Patricia Daukantas. ‘Virtual posters’ will broaden impact of PC ’96. Computers in Physics, 10(4):312–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822407.

Duax:1996:WCW

[1164] William L. Duax. World of crystallography on the Web. Computers in Physics, 10(4):315–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822408.

Daukantas:1996:SHA

[1165] Patricia Daukantas. Sun helps astronomers’ search for planets. Computers in Physics, 10(4):316–??, July/August 1996. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822409.

Bothun:1996:NPU

[1166] G. D. Bothun and S. D. Kevan. Networked physics in undergraduate instruction. Computers in Physics, 10(4):318–324, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822410.

Huergo:1996:RCH

[1167] Jennifer Huergo. The reorganized CIP home page. Computers in Physics, 10(4):325–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822411.

Anonymous:1996:WLH

[1168] Anonymous. The world lecture hall. Computers in Physics, 10(4):325– ??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822412.

Lister:1996:WS

[1169] Tim Lister and James R. Matey. What is software? Computers in Physics, 10(4):327–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822413.

Booten:1996:CLSb

[1170] Albert Booten and Henk van der Vorst. Cracking large-scale eigenvalue problems, part II: Implementations. Computers in Physics, 10(4):331–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168590.

Rawitscher:1996:VSF

[1171] G. Rawitscher, I. Koltracht, Hong Dai, C. Ribetti, Denis Donnelly, George Rawitscher, Israel Koltracht, Hong Dai, and Christopher Ribetti. The vibrating string: a fertile topic for teaching scientific computing. Computers in Physics, 10(4):335–??, July/August 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822414.

Stauffer:1996:GOM

[1172] Dietrich Stauffer. Getting older — Monte Carlo simulations of biological aging. Computers in Physics, 10(4):341–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168589.

Meredith:1996:BRH

[1173] Dawn C. Meredith and Bill Titus. Book reviews: Harvey Gould and Jan Tobochnik, An Introduction to Computer Simulation Methods: Applications to Physical Systems, 2nd Edition. Richard L. Gaylord and Paul R. Wellin, Mathematica: Explorations in Complex Physical and Biological Systems, Springer-Verlag, New York, 1995; ISBN 0-387-94274-2; 297 pp. plus CD-ROM, cloth, $54.95. Computers in Physics, 10(4):349–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822415.

Titus:1996:BRR

[1174] Bill Titus, Susan R. McKay, and Wolfgang Christian. Book review: Richard J. Gaylord and Paul R. Wellin, Computer Simulations with Mathematica: Explorations in Complex Physical and Biological Systems. Computers in Physics, 10(4):349–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822455.

Makino:1996:CAS

[1175] Junichiro Makino and Makoto Taiji. Computational astrophysics on a special-purpose machine. Computers in Physics, 10(4):352–??, July/ August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822456.

Dubois:1996:EPO

[1176] P. F. Dubois and T.-Y. Yang. Extending Python [object-oriented language]. Computers in Physics, 10(4):359–365, July/August 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822457.

Roy:1996:ACM

[1177] Trina Roy and Steve Bryson. Accurate color mapping through textures. Computers in Physics, 10(4):366–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822458.

Thompson:1996:HMP

[1178] William J. Thompson. How to make a PC talk with a MAC. Computers in Physics, 10(4):371–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822459.

Anonymous:1996:LLRd

[1179] Anonymous. Listing of the latest releases. Computers in Physics, 10 (4):372–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822460.

Bellomonte:1996:CSE

[1180] L. Bellomonte and R. M. Sperandeo-Mineo. Computer simulation of electric conductivity in a solid using classical mechanics. Computers in Physics, 10(4):376–??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168575.

Allodi:1996:RPW

[1181] G. Allodi and R. Coïsson. Reflection and propagation of waves in onedimensional quasi-periodic structures. Computers in Physics, 10(4):385– ??, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168576.

OShea:1996:GMP

[1182] Donald C. O’Shea. Generation of mask patterns for diffractive optical elements using Mathematica^{T M}. Computers in Physics, 10(4):391–399, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168577.

Wang:1996:TED

[1183] J. B. Wang. Time evolution and decay of an excited atom in a weak electric field. Computers in Physics, 10(4):400–406, July/August 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168578.

Holmes:1996:FFE

[1184] Lewis M. Holmes. Fast Fourier explication. Computers in Physics, 10(5):411–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822461.

West:1996:PIR

[1185] Thomas G. West. Playing with images — a return to thinking in pictures. Computers in Physics, 10(5):413–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822462.

Lewin:1996:BIC

[1186] David I. Lewin. Building an imaging collaboratory. Computers in Physics, 10(5):414–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822463.

Sinclair:1996:ECB

[1187] Bruce D. Sinclair. European competition boosts physics software. Computers in Physics, 10(5):417–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822464.

Daukantas:1996:SPC

[1188] Patricia Daukantas. SC ’96 puts computers to work. Computers in Physics, 10(5):418–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822465.

Thiebaut:1996:WTG

[1189] Dominique F. Thièbaut and Glenn Ricart. Web tracking and graphing with Pgperl. Computers in Physics, 10(5):420–??, September 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822466.

Anonymous:1996:IGC

[1190] Anonymous. Internet goldmine: CFD online. Computers in Physics, 10(5):426–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822468.

Huergo:1996:AUC

[1191] Jennifer Huergo. Additions and updates to the CIP home page. Computers in Physics, 10(5):426–??, September 1996. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822467.

Anonymous:1996:PC

[1192] Anonymous. Physics Computing ’96. Computers in Physics, 10(5):427– ??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822469.

Visscher:1996:FFT

[1193] P. B. Visscher. The FFT: Fourier transforming one bit at a time. Computers in Physics, 10(5):438–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168581.

Backstrom:1996:LAW

[1194] Gunnar Backstrom and Denis Donnelly. Learning about wave mechanics by FEA on a PC. Computers in Physics, 10(5):444–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822470.

Andricioaei:1996:FNH

[1195] Ioan Andricioaei and John E. Straub. Finding the needle in the haystack: Algorithms for conformational optimization. Computers in Physics, 10 (5):449–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168582.

Blumberg:1996:BRD

[1196] Robin L. Blumberg, Susan R. McKay, and Wolfgang Christian. Book review: D. C. Rapaport, The Art of Molecular Dynamics Simulation. Computers in Physics, 10(5):456–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822471.

Cook:1996:BRG

[1197] David M. Cook, Susan R. McKay, and Wolfgang Christian. Book review: G. Silverman and H. Silver, Modern Instrumentation: a Computer Approach. Computers in Physics, 10(5):456–??, September 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822472.

Robison:1996:CGF

[1198] Arch D. Robison and Paul F. Dubois. C++ gets faster for scientific computing. Computers in Physics, 10(5):458–??, September 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822473.

Fernandez:1996:VCP

[1199] Victor M. Fernandez, Deborah Silver, Norman J. Zabusky, and Steve Bryson. Visiometrics of complex physical processes: Diagnosing vortexdominated flows. Computers in Physics, 10(5):463–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822474.

Fell:1996:MRI

[1200] Richard Noel Fell. Macsyma 2.1 rebounds with improved capabilities. Computers in Physics, 10(5):471–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822475.

Anonymous:1996:LLRe

[1201] Anonymous. Listings of the latest releases. Computers in Physics, 10(5):474–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822488.

Goloskie:1996:BES

[1202] R. Goloskie, T. Thio, and L. R. Ram-Mohan. Boundary elements and surface plasmons. Computers in Physics, 10(5):477–??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168580.

Weniger:1996:CWF

[1203] Ernst Joachim Weniger. Computation of the Whittaker function of the second kind by summing its divergent asymptotic series with the help of nonlinear sequence transformations. Computers in Physics, 10(5):496– ??, September 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168579.

Holmes:1996:SCA

[1204] Lewis M. Holmes. Software-contest awards. Computers in Physics, 10(6):507–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822489.

Pifer:1996:WDO

[1205] Joe Pifer. Are we doing our students a favor by making them learn Fortran? Computers in Physics, 10(6):509–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822490.

Lewin:1996:SPD

[1206] David I. Lewin. Striking pay dirt in a supercomputer. Computers in Physics, 10(6):510–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822491.

Daukantas:1996:CSE

[1207] Patricia Daukantas. Computer-supported education gets spotlight in College Park. Computers in Physics, 10(6):513–??, November 1996. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822492.

Daukantas:1996:CJS

[1208] Patricia Daukantas. From the CIP journal section. Computers in Physics, 10(6):519–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822493.

Canessa:1996:OSS

[1209] Enrique Canessa, Giorgio Pastore, and Glenn Ricart. One-shot service searches preprint repositories at a mouseclick. Computers in Physics, 10(6):520–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822494.

Huergo:1996:DSC

[1210] Jennifer Huergo. Download software-contest forms. Computers in Physics, 10(6):524–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822495.

Anonymous:1996:SPS

[1211] Anonymous. Space physics at SWRI. Computers in Physics, 10(6):524– ??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822496.

Yoshiyuki:1996:STS

[1212] Usami Yoshiyuki, Kano Yoshiki, and Kitaoka Masatoshi. Solving the traveling-salesman problem by a statistical-physics method. Computers in Physics, 10(6):525–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168588.

Donnelly:1996:CSA

[1213] Denis Donnelly. CIP’s seventh annual educational software contest: The winners. Computers in Physics, 10(6):532–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822497.

Tobochnik:1996:LSB

[1214] Jan Tobochnik and Harvey Gould. Lattice simulations of biological membranes. Computers in Physics, 10(6):542–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168587.

Ytreberg:1996:BRH

[1215] F. Marty Ytreberg. Book review: Herbert Schildt, C++: The Complete Reference, Second Edition. Computers in Physics, 10(6):549–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822498.

Haney:1996:BAP

[1216] Scott W. Haney and Paul F. Dubois. Beating the abstraction penalty in C++ using expression templates. Computers in Physics, 10(6):552– ??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822499.

Pauschenwein:1996:VQM

[1217] J. Pauschenwein and B. Thaller. Visualizing quantum-mechanical wavefunctions in three dimensions with AVS. Computers in Physics, 10(6): 558–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822500.

Thompson:1996:UPS

[1218] William J. Thompson. Uninterruptible power supply delivers. Computers in Physics, 10(6):567–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822501.

Anonymous:1996:LLRf

[1219] Anonymous. Listings of the latest releases. Computers in Physics, 10(6):569–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822502.

Belien:1996:VRA

[1220] A. J. C. Beliën, S. Poedts, H. J. W. Spoelder, R. Leenders, and J. P. Goedbloed. Visualization of resonant absorption in solar coronal loops by simulation of soft X-ray images. Computers in Physics, 10(6):573– ??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168586.

Clark:1996:HMC

[1221] M. M. Clark, L. M. Raff, and H. L. Scott. Hybrid Monte Carlo method for off-lattice simulation of processes involving steps with widely varying rates. Computers in Physics, 10(6):584–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168583.

Dauger:1996:SSF

[1222] Dean E. Dauger. Simulation and study of Fresnel diffraction for arbitrary two-dimensional apertures. Computers in Physics, 10(6):591–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168584.

Bischof:1996:SAM

[1223] Christian H. Bischof, Gordon D. Pusch, and Ralf Knoesel. Sensitivity analysis of the MM5 weather model using automatic differentiation. Computers in Physics, 10(6):605–??, November 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168585.

Holmes:1997:ITN

[1224] Lewis M. Holmes. Introducing two new departments. Computers in Physics, 11(1):3–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822509.

Holmes:1997:CCT

[1225] Lewis M. Holmes. Celebrating CIP’s tenth anniversary year. Computers in Physics, 11(1):5–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4820243.

Lewin:1997:GTK

[1226] David I. Lewin. Growing a tree of knowledge in an object-oriented Eden. Computers in Physics, 11(1):6–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822518.

Carts-Powell:1997:IIN

[1227] Yvonne Carts-Powell. Internet 2 to implement next-generation technology. Computers in Physics, 11(1):8–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822519.

Daukantas:1997:CJSa

[1228] Patricia Daukantas. From the CIP journal section. Computers in Physics, 11(1):16–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822503.

Zollman:1997:IFW

[1229] Kevin Zollman and Dean Zollman. Interactive forms on the Web. Computers in Physics, 11(1):17–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822504.

Huergo:1997:SOE

[1230] Jennifer Huergo. See our 1997 editorial calendar. Computers in Physics, 11(1):21–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822505.

Anonymous:1997:SM

[1231] Anonymous. Science for the millennium. Computers in Physics, 11(1): 21–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822506.

Matey:1997:SDA

[1232] James R. Matey. Software designs and architectures, Part II: Representing observations. Computers in Physics, 11(1):22–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822507.

Gerteisen:1997:DDT

[1233] Edgar Gerteisen, Ralf Gruber, Silvio Merazzi, and Jacques Rappaz. Domain decomposition: Theory and practice. Computers in Physics, 11(1): 26–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822508.

Martinez-Jimenez:1997:IPS

[1234] P. Martínez-Jiménez, E. Casado, J. M. Martínez-Jiménez, M. Cuevas-Rubiño, D. González-Caballero, F. Zafra-López, and Denis Donnelly. Interactive physics simulations appeal to first-year students. Computers in Physics, 11(1):31–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822510.

Karabach:1997:IBA

[1235] Michael Karabach, Gerhard Müller, Harvey Gould, and Jan Tobochnik. Introduction to the Bethe Ansatz. I. Computers in Physics, 11(1):36– ??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822511.

Birnbaum:1997:BRF

[1236] Dave Birnbaum, Susan McKay, and Wolfgang Christian. Book review: Forman S. Acton, Real Computing Made Real: Preventing Errors in Scientific and Engineering Calculations. Computers in Physics, 11(1): 45–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822512.

Sipson:1997:BRD

[1237] Roger Sipson, Susan McKay, and Wolfgang Christian. Book review: David Flanagan, Java in a Nutshell; Patrick Niemeyer and Joshua Peck, Exploring JAVA. Computers in Physics, 11(1):46–??, January 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822513.

Christian:1997:CAL

[1238] Wolfgang Christian. The CIP anniversary library of classics: a call for nominations. Computers in Physics, 11(1):48–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822515.

Anonymous:1997:TAC

[1239] Anonymous. Tenth anniversary CIP educational software directory. Computers in Physics, 11(1):49–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822517.

Dotzel:1997:OHP

[1240] Günter Dotzel, Wojtek Skulski, and Paul F. Dubois. Oberon-2, a high-performance alternative to C++. Computers in Physics, 11(1): 81–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822520.

Coisson:1997:MCC

[1241] R. Coïsson and G. Allodi. Matlab compiler and C math library simplify programming chores. Computers in Physics, 11(1):89–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822521.

Anonymous:1997:LLRa

[1242] Anonymous. Listings of the latest releases. Computers in Physics, 11(1): 91–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822522.

Barach:1997:CSP

[1243] John Paul Barach. Computational study of patterns in simple nonequilibrium systems. Computers in Physics, 11(1):96–??, January 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168599.

Spencer:1997:HSP

[1244] C. D. Spencer and S. R. Paul. Hardware and software for a pulse height analyzer linked to a personal computer. Computers in Physics, 11(1):

101–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168598.

Lampton:1997:DUS

[1245] Michael Lampton. Damping–undamping strategies for the Levenberg– Marquardt nonlinear least-squares method. Computers in Physics, 11(1): 110–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168600.

Baines:1997:NUC

[1246] C. G. Stafford Baines and Michael E. Bacon. Note on using computer simulations in teaching introductory digital electronics courses. Computers in Physics, 11(1):116–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168601.

Holmes:1997:WWF

[1247] Lewis M. Holmes. What will the future bring? Computers in Physics, 11(2):123–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822523.

Levine:1997:WEM

[1248] Judah Levine. Why experimentalists mistrust computer modeling — and why they rely on it. Computers in Physics, 11(2):125–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822524.

Lewin:1997:AFD

[1249] David I. Lewin. Advising fluid dynamicists to be discrete. Computers in Physics, 11(2):126–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822525.

Daukantas:1997:ACR

[1250] Patricia Daukantas. ASCI clocks its ‘Red’ computer at 1 teraflops. Computers in Physics, 11(2):129–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822526.

Anonymous:1997:CJS

[1251] Anonymous. From the CIP journal section. Computers in Physics, 11(2): 132–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822527.

Zollman:1997:PPD

[1252] Kevin Zollman and Dean Zollman. Posting physics documents on the Web. Computers in Physics, 11(2):133–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822528.

Huergo:1997:SCD

[1253] Jennifer Huergo. Software contest deadline approaches. Computers in Physics, 11(2):139–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822529.

Anonymous:1997:CPS

[1254] Anonymous. The center for polymer studies. Computers in Physics, 11 (2):139–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822530.

Matey:1997:MLP

[1255] James R. Matey. Moore’s law and the PC revolution: Actual and anticipated effects on instrumentation and measurement science. Computers in Physics, 11(2):140–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822531.

Rappaz:1997:BEM

[1256] J. Rappaz and M. Swierkosz. Boundary-element method yields external vector potentials in complex industrial applications. Computers in Physics, 11(2):145–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168595.

Boyce:1997:IMM

[1257] William E. Boyce, John B. Brunski, Michael M. Chen, Kevin Craig, James A. Fairweather, Joseph W. Haus, Jr., Joseph Porter, Jr., William L. Siegmann, and Denis Donnelly. Interactive multimedia modules in mathematics, engineering, and science. Computers in Physics, 11 (2):151–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822532.

Gould:1997:EDP

[1258] Harvey Gould, Jan Tobochnik, and Louis Colonna-Romano. Entropydriven phase transitions. Computers in Physics, 11(2):157–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168596.

Styer:1997:BRM

[1259] Daniel Styer, Susan R. McKay, and Wolfgang Christian. Book review: Marko Horbatsch, Quantum Mechanics Using Maple. Computers in Physics, 11(2):165–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822533.

McKay:1997:BRK

[1260] Susan R. McKay. Book review: K. Binder, editor, The Monte Carlo Method in Condensed Matter Physics: Second, Corrected and Updated Edition. Computers in Physics, 11(2):166–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822534.

Dubois:1997:FSP

[1261] Paul F. Dubois. The future of scientific programming. Computers in Physics, 11(2):168–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822535.

Bryson:1997:VVR

[1262] Steve Bryson. Visualization ’96: Volume rendering shifts into high Gear. Computers in Physics, 11(2):174–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822536.

Mathews:1997:TSO

[1263] Kirk A. Mathews. TK Solver 3.0 organizes and solves physics problems. Computers in Physics, 11(2):181–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822537.

Anonymous:1997:LLRb

[1264] Anonymous. Listings of the latest releases. Computers in Physics, 11(2): 185–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822538.

Reader:1997:TAA

[1265] Joseph Reader. Transition arrays in atomic spectroscopy with a commercial spreadsheet program. Computers in Physics, 11(2):190–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168597.

He:1997:SRE

[1266] Song He. Stabilizing the Richardson eigenvector algorithm by controlling chaos. Computers in Physics, 11(2):194–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168602.

Tao:1997:CSA

[1267] Ping-Kee Tao. Confronting students’ alternative conceptions in mechanics with the force and motion microworld. Computers in Physics, 11(2): 199–??, March 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168603.

Holmes:1997:CMW

[1268] Lewis M. Holmes. A computational model that works. Computers in Physics, 11(3):211–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822539.

Kuck:1997:FSG

[1269] David J. Kuck. Facing up to software’s greatest challenge: Practical parallel processing. Computers in Physics, 11(3):213–??, May 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822540.

Lewin:1997:HET

[1270] David I. Lewin. How to be an experimental theorist. Computers in Physics, 11(3):214–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822541.

Daukantas:1997:SDI

[1271] Patricia Daukantas. San Diego, Illinois sites to lead NSF’s new supercomputing partnerships. Computers in Physics, 11(3):216–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822542.

Daukantas:1997:AFN

[1272] Patricia Daukantas. ACM97 forecasts the next 50 years of ‘paleotechnic human’. Computers in Physics, 11(3):217–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822543.

Daukantas:1997:DDB

[1273] Patricia Daukantas. Dome design brings numerical astrophysics down to Earth. Computers in Physics, 11(3):220–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822544.

Daukantas:1997:CJSb

[1274] Patricia Daukantas. From the CIP journal section. Computers in Physics, 11(3):224–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822545.

Zollman:1997:CSS

[1275] Kevin Zollman and Dean Zollman. Creating a simple, searchable database on the Web. Computers in Physics, 11(3):225–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822546.

Huergo:1997:NSO

[1276] Jennifer Huergo. New secure online subscription form. Computers in Physics, 11(3):229–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822547.

Anonymous:1997:SDS

[1277] Anonymous. San Diego Supercomputer Center. Computers in Physics, 11 (3):229–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822548.

Beazley:1997:CDG

[1278] David M. Beazley and Peter S. Lomdahl. Controlling the data glut in large-scale molecular-dynamics simulations. Computers in Physics, 11 (3):230–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822549.

Cook:1997:CLPa

[1279] David M. Cook and Denis Donnelly. Computers in the Lawrence physics curriculum — Part I. Computers in Physics, 11(3):240–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822550.

Frenkel:1997:SPC

[1280] Daan Frenkel and Berend Smit. Simulation of phase coexistence for complex molecules. Computers in Physics, 11(3):246–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168607.

Snowman:1997:BRM

[1281] Daniel P. Snowman, Susan R. McKay, and Wolfgang Christian. Book review: Martin Campbell-Kelly and William Aspray, Computer: A History of the Information Machine. Computers in Physics, 11(3):256–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822551.

Wolpert:1997:BRS

[1282] Seth Wolpert, Susan R. McKay, and Wolfgang Christian. Book review: Stan Franklin, Artificial Minds. Computers in Physics, 11(3):258–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822552.

Grest:1997:BRD

[1283] Gary S. Grest, Susan R. McKay, and Wolfgang Christian. Book review: David J. Kuck, High Performance Computing — Challenges for Future Systems. Computers in Physics, 11(3):259–??, May 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822553.

Harrington:1997:BRN

[1284] Randal Harrington, Susan R. McKay, and Wolfgang Christian. Book review: Nicholas Negroponte, Being Digital. Computers in Physics, 11 (3):261–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822554.

Furnish:1997:DGE

[1285] Geoffrey Furnish and Paul F. Dubois. Disambiguated glommable expression templates. Computers in Physics, 11(3):263–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822555.

Bryson:1997:TDT

[1286] Steve Bryson. Time, data-time, and real-time interactive visualization. Computers in Physics, 11(3):270–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822556.

Burr:1997:IPD

[1287] A. F. Burr. Interactive Physics 3.0 delights Windows users also. Computers in Physics, 11(3):275–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822557.

Anonymous:1997:LLRc

[1288] Anonymous. Listings of the latest releases. Computers in Physics, 11(3): 278–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822558.

Mirasso:1997:VSA

[1289] Claudio R. Mirasso, Martijn Mulder, Hans J. W. Spoelder, and Daan Lenstra. Visualization of the “Sisyphus” attractor. Computers in Physics, 11(3):282–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168604.

Cole:1997:HAS

[1290] James B. Cole. High accuracy solution of Maxwell’s equations using nonstandard finite differences. Computers in Physics, 11(3):287–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168620. See erra-

tum [1409].

John:1997:SSN

[1291] George C. John, J. E. Hasbun, and Vijay A. Singh. Simple scheme for the numerical evaluation of nearly singular integrals. Computers in Physics, 11(3):293–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168605.

Boyd:1997:PAA

[1292] John P. Boyd. Padé approximant algorithm for solving nonlinear ordinary differential equation boundary value problems on an unbounded domain. Computers in Physics, 11(3):299–??, May 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168606.

Holmes:1997:VF

[1293] Lewis M. Holmes. Visualizing the future. Computers in Physics, 11(4): 307–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822559.

Klein:1997:CIP

[1294] Barry M. Klein. Computational and industrial physics at PC ’97: a marriage made in heaven. Computers in Physics, 11(4):309–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822560.

Lewin:1997:GSH

[1295] David I. Lewin. Getting students high on quantum chemistry. Computers in Physics, 11(4):310–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822561.

Daukantas:1997:PMH

[1296] Patricia Daukantas. PC ’97 meeting to highlight variety of applications and methods. Computers in Physics, 11(4):312–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822562.

Daukantas:1997:CJSc

[1297] Patricia Daukantas. From the CIP journal section. Computers in Physics, 11(4):316–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822563.

Rebello:1997:DIW

[1298] N. Sanjay Rebello, Konstantin Sushenko, Kevin Zollman, and Dean Zollman. Designing interactive Web pages using ActiveX. Computers in Physics, 11(4):317–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822564.

Huergo:1997:LR

[1299] Jennifer Huergo. LogoLinks on the rise. Computers in Physics, 11(4): 323–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822565.

Anonymous:1997:PEH

[1300] Anonymous. The Planet Earth home page. Computers in Physics, 11 (4):323–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822566.

Bossy:1997:PNM

[1301] Mireille Bossy, Marco Picasso, and Denis Talay. Probabilistic numerical methods for physical and financial problems. Computers in Physics, 11 (4):325–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168611.

Cook:1997:CLPb

[1302] David M. Cook and Denis Donnelly. Computers in the Lawrence physics curriculum — Part II. Computers in Physics, 11(4):331–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822567.

Gould:1997:BRD

[1303] Harvey Gould and Jan Tobochnik. Book review: D. C. Rapaport, An introduction to interactive molecular-dynamics simulations. Computers in Physics, 11(4):337–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168612.

Christian:1997:CLC

[1304] Wolfgang Christian. The CIP library of classics. Computers in Physics, 11(4):348–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822568.

DeJong:1997:BRN

[1305] Marvin L. De Jong, Susan R. McKay, and Wolfgang Christian. Book review: Nicholas J. Giordano, Computational Physics. Computers in Physics, 11(4):351–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822569.

Tobochnik:1997:BRD

[1306] Jan Tobochnik, Susan R. McKay, and Wolfgang Christian. Book review: Daan Frenkel and Berend Smit, Understanding Molecular Simulation. Computers in Physics, 11(4):351–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822570.

Dubois:1997:BRM

[1307] Paul F. Dubois. Book review: Mark G. Gray and Randy M. Roberts, Object-Based Programming in Fortran 90. Computers in Physics, 11(4): 355–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822571.

Bryson:1997:NTY

[1308] Steve Bryson. The next ten years of visualization. Computers in Physics, 11(4):362–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822572.

Thompson:1997:TFS

[1309] William J. Thompson. Three Fortran systems for desktop computers. Computers in Physics, 11(4):370–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822573.

Anonymous:1997:LLRd

[1310] Anonymous. Listings of the latest releases. Computers in Physics, 11(4): 374–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822574.

Sigurdsson:1997:IEC

[1311] Steinn Sigurdsson, Bohr He, Rami Melhem, and Lars Hernquist. Implementing an efficient collisionless N-body code on the Cray T3D. Computers in Physics, 11(4):378–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168608.

MacLeod:1997:AEE

[1312] Allan J. MacLeod. Accurate and efficient evaluation of the Bose–Einstein functions g_{3/2} and g_{5/2}. Computers in Physics, 11(4):385–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168609.

Villarreal:1997:DTD

[1313] Alberto Villarreal and John A. Scales. Distributed three-dimensional finite-difference modeling of wave propagation in acoustic media. Computers in Physics, 11(4):388–??, July 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168610.

Holmes:1997:FF

[1314] Lewis M. Holmes. Fine fare. Computers in Physics, 11(5):403–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822575.

Kennedy:1997:CNG

[1315] Ken Kennedy. The case for a next generation Internet. Computers in Physics, 11(5):405–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822577.

Lewin:1997:MGW

[1316] David I. Lewin. Mr. GNU’s who. Computers in Physics, 11(5):406–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822578.

Daukantas:1997:SEI

[1317] Patricia Daukantas. SC97 to examine interdisciplinary applications. Computers in Physics, 11(5):408–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822579.

DeHaven:1997:LAC

[1318] William DeHaven, Rami Arieli, Kevin Zollman, and Dean Zollman. “Laser adventure”: Converting print material to a Web site. Computers in Physics, 11(5):411–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822580.

Anonymous:1997:OJS

[1319] Anonymous. Online journal section gets there first. Computers in Physics, 11(5):415–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822581.

Anonymous:1997:CW

[1320] Anonymous. Calculus on the Web. Computers in Physics, 11(5):415–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822582.

Press:1997:NRD

[1321] William H. Press and Saul A. Teukolsky. Numerical recipes: Does this paradigm have a future? Computers in Physics, 11(5):416–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822583.

Singh:1997:CRB

[1322] A. K. Singh, J. Panigrahi, and James R. Matey. Catching the right bus, Part VII: a CAMAC example. Computers in Physics, 11(5):425–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822584.

Vasilyev:1997:SPU

[1323] Oleg V. Vasilyev, David A. Yuen, Samuel Paolucci, Ralf Gruber, and Jacques Rappaz. Solving PDEs using wavelets. Computers in Physics, 11(5):429–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822585.

Christian:1997:DCR

[1324] Wolfgang Christian and Denis Donnelly. Developing a computer-rich physics curriculum at a liberal-arts College. Computers in Physics, 11(5):436–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822586.

deOliveira:1997:CSN

[1325] P. M. C. de Oliveira, Harvey Gould, and Jan Tobochnik. Computer simulations of neural networks. Computers in Physics, 11(5):443–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822587.

Hubisz:1997:BRF

[1326] John L. Hubisz, Susan R. McKay, and Wolfgang Christian. Book review: Ferdi Serim and Melissa Koch, NetLearning: Why Teachers Use the Internet. Computers in Physics, 11(5):451–??, September 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822588.

Hubisz:1997:BRE

[1327] John L. Hubisz, Susan R. McKay, and Wolfgang Christian. Book review: Edward J. Renehan, Jr., Science on the Web: a Connoisseur’s Guide to Over 500 of the Best Most Useful, and Most Fun Science Websites. Computers in Physics, 11(5):451–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822589.

Greene:1997:BRR

[1328] Ronald L. Greene. Book review: Richard H. Enns and George C. McGuire, Nonlinear Physics with Maple for Scientists and Engineers. Computers in Physics, 11(5):451–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822590.

Messina:1997:HPCa

[1329] Paul Messina. High-performance computers: The next generation (Part I). Computers in Physics, 11(5):454–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822591.

Steiner:1997:OOT

[1330] M. M. Steiner, W. Wenzel, J. W. Wilkins, K. G. Wilson, and Paul F. Dubois. Object-oriented techniques in large scientific computing projects: Experience with C++. Computers in Physics, 11(5):467–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822592.

Shen:1997:ULI

[1331] Han-Wei Shen and Steve Bryson. Using line-integral convolution to visualize dense vector fields. Computers in Physics, 11(5):474–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822593.

Anonymous:1997:LLRe

[1332] Anonymous. Listings of the latest releases. Computers in Physics, 11(5):480–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822594.

deRonde:1997:SGW

[1333] J. F. de Ronde, G. D. van Albada, and P. M. A. Sloot. Simulation of gravitational wave detectors. Computers in Physics, 11(5):484–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168591.

Mandrake:1997:LSW

[1334] L. Mandrake and W. Gekelman. Large-scale Windows 95-based dataacquisition system using LabVIEW. Computers in Physics, 11(5):498– ??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168613.

Takach:1997:DCS

[1335] Stephen F. Takach and Kirsten Nelson. Design of the control software for the muon readout system in NA48. Computers in Physics, 11(5):508–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168592.

Gonzalez:1997:GSN

[1336] J. L. M. Quiroz González and D. Thompson. Getting started with Numerov’s method. Computers in Physics, 11(5):514–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168593.

Ghaousia:1997:DPB

[1337] Syedah Fatima Ghaousia, Nasir M. Mirza, and Sikander M. Mirza. Development of PC-based nuclear instruments for gamma-ray spectroscopy using scintillation detectors. Computers in Physics, 11(5):516–??, September 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168594.

Holmes:1997:LF

[1338] Lewis M. Holmes. Looking to the future. Computers in Physics, 11(6): 523–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822595.

Brodsky:1997:TYC

[1339] Marc H. Brodsky. Ten years of computers and physics. Computers in Physics, 11(6):525–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822596.

Lewin:1997:SPH

[1340] David I. Lewin. Still pumping heavy iron. Computers in Physics, 11(6): 526–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822597.

Holmes:1997:PCM

[1341] Lewis M. Holmes and Patricia Daukantas. Physics computing moves out of the Ivory Tower. Computers in Physics, 11(6):528–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822598.

Daukantas:1997:CJSd

[1342] Patricia Daukantas. From the CIP journal section. Computers in Physics, 11(6):535–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822599.

Becker:1997:HLS

[1343] Mark Becker. Hyperimages lead to sample articles. Computers in Physics, 11(6):536–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822600.

Anonymous:1997:LGN

[1344] Anonymous. LASSPTools: Graphical and numerical extensions to Unix. Computers in Physics, 11(6):536–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822601.

Rebello:1997:UDC

[1345] N. Sanjay Rebello, Kevin Zollman, and Dean Zollman. Using demoshield to create interactive demos on the Web. Computers in Physics, 11(6): 537–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822602.

Anonymous:1997:FSC

[1346] Anonymous. The future of scientific computing. Computers in Physics, 11(6):542–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822603.

Borchers:1997:WCW

[1347] Robert R. Borchers. 1987, when the computin’ was easy: Some reflections on the state of the art in high-speed computation. Computers in Physics, 11(6):564–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822604.

Matey:1997:WBE

[1348] James R. Matey. A world without barriers: Editor’s report from Niweek ’97. Computers in Physics, 11(6):570–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822605.

Papageorgiou:1997:FEM

[1349] A. Papageorgiou and J. F. Traub. Faster evaluation of multidimensional integrals. Computers in Physics, 11(6):574–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168616.

Donnelly:1997:CEA

[1350] Denis Donnelly. CIP’s eighth annual educational software contest: The winners. Computers in Physics, 11(6):579–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822606.

Alexander:1997:DSM

[1351] Francis J. Alexander and Alejandro L. Garcia. The direct simulation Monte Carlo method. Computers in Physics, 11(6):588–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168619.

Templon:1997:BRG

[1352] J. A. Templon, W. M. Dennis, Susan R. McKay, and Wolfgang Christian. Book review: Gisela Engeln-Müllges and Frank Uhlig, Numerical Algorithms with Fortran. Numerical Algorithms with C. Computers in Physics, 11(6):594–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822607.

Clendenning:1997:BRG

[1353] Lester M. Clendenning, Patrick T. Tam, Susan R. McKay, and Wolfgang Christian. Book review: Gerd Baumann, Mathematica in Theoretical Physics: Selected Examples from Classical Mechanics to Fractals. Computers in Physics, 11(6):595–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822608.

Daukantas:1997:BRP

[1354] Patricia Daukantas, Susan R. McKay, and Wolfgang Christian. Book review: Peter J. Denning and Robert M. Metcalfe, Beyond Calculation: The Next Fifty Years of Computing. Computers in Physics, 11(6):597– ??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822609.

Messina:1997:HPCb

[1355] Paul Messina. High-performance computers: The next generation (Part II). Computers in Physics, 11(6):598–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822610.

Dubois:1997:JSP

[1356] Paul F. Dubois. Is Java for scientific programming? Computers in Physics, 11(6):611–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822611.

Schmidt:1997:AUE

[1357] Jeffrey R. Schmidt and Steve Bryson. Animation in the Unix environment: a primer. Computers in Physics, 11(6):618–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.4822612.

Shirer:1997:SLL

[1358] Donald L. Shirer. Symbolic I/O leads list of new Mathematica 3.0.1 features. Computers in Physics, 11(6):624–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822613.

Anonymous:1997:SPT

[1359] Anonymous. Signal-processing toolkit. Computers in Physics, 11(6):626– ??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822614.

Danese:1997:PSP

[1360] G. Danese, I. De Lotto, D. Dotti, and F. Leporati. A parallel specialpurpose computer dedicated to the simulation of interacting particle systems. Computers in Physics, 11(6):630–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168617.

Schweizer:1997:DVM

[1361] W. Schweizer and P. Fassbinder. Discrete variable method for nonintegrable quantum systems. Computers in Physics, 11(6):641–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168618.

Schulz:1997:PAM

[1362] Stefan Schulz and Regina de Vivie-Riedle. A practical approach to multivariate interpolation. Computers in Physics, 11(6):647–??, November 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168614.

Hua:1997:MCS

[1363] Xin-Min Hua. Monte Carlo simulation of Comptonization in inhomogeneous media. Computers in Physics, 11(6):660–??, November 1997. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168615.

Holmes:1998:MRA

[1364] Lewis M. Holmes. Moving right along. Computers in Physics, 12(1): 3–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822616.

Truchard:1998:IJP

[1365] James Truchard. Instrumentation joins the PC revolution. Computers in Physics, 12(1):5–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822617.

Lewin:1998:MMS

[1366] David I. Lewin. Making molecular structures crystal clear. Computers in Physics, 12(1):6–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822618.

Forbes:1998:QCS

[1367] Nancy Forbes and Seth Lloyd. Quantum computing: Stepping closer to reality. Computers in Physics, 12(1):8–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822620.

Daukantas:1998:CJSa

[1368] Patricia Daukantas. From the CIP journal section. Computers in Physics, 12(1):15–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822615.

Takats:1998:CDA

[1369] Martha Takats, Kevin Zollman, and Dean Zollman. Creating and displaying animated GIFs for physics instruction. Computers in Physics, 12(1): 16–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168682.

Bothun:1998:VLU

[1370] Gregory Bothun, Sean Russell, and Amy Hulse. Virtual Laboratory (University of Oregon). Computers in Physics, 12(1):20–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168650.

Becker:1998:CNA

[1371] Mark Becker. CIP’s ninth annual educational software contest. Computers in Physics, 12(1):20–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168683.

Descloux:1998:RQI

[1372] J. Descloux, J.-L. Fattebert, F. Gygi, Ralf Gruber, and Jacques Rappaz. Rayleigh Quotient Iteration, an old recipe for solving modern large-scale eigenvalue problems. Computers in Physics, 12(1):22–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168687.

Rebello:1998:SSL

[1373] N. Sanjay Rebello, Chandima Cumaranatunge, Lawrence T. Escalada, Dean A. Zollman, and Denis Donnelly. Simulating the spectra of light sources. Computers in Physics, 12(1):28–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168684.

Ferguson:1998:LRE

[1374] Charles D. Ferguson, W. Klein, John B. Rundle, Harvey Gould, and Jan Tobochnik. Long-range earthquake fault models. Computers in Physics, 12(1):34–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168681.

Luzar:1998:BRH

[1375] Alenka Luzar. Book review: H. D. Holtje and G. Folkers, Molecular Modeling, Basic Principles and Applications. Computers in Physics, 12(1): 41–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168645.

Novak:1998:BRC

[1376] Gregor M. Novak. Book review: Chuck Musciano and Bill Kennedy, HTML: The Definitive Guide, Second Edition. Computers in Physics, 12(1):41–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168646.

Novak:1998:BRD

[1377] Gregor M. Novak. Book review: David Flanagan, Java-Script: The Definitive Guide, Second Edition. Computers in Physics, 12(1):41– ??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168647.

Bennett:1998:BRP

[1378] Charles A. Bennett. Book review: Patrick Tam, A Physicist’s Guide to Mathematica. Computers in Physics, 12(1):41–??, January 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168685.

Bik:1998:CSJ

[1379] Aart J. C. Bik, Fabian Breg, Dennis B. Gannon, and Paul F. Dubois. Compiler support for Java parallelization. Computers in Physics, 12(1): 45–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168686.

Bryson:1998:VBN

[1380] Steve Bryson. Visualization ’97 brings new technical developments. Computers in Physics, 12(1):53–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168648.

Shirer:1998:BAM

[1381] Donald L. Shirer. BBEdit 4.5 adds more HTML tools. Computers in Physics, 12(1):60–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168649.

Anonymous:1998:NPa

[1382] Anonymous. New products. Computers in Physics, 12(1):62–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.4822619.

Hansen:1998:EVP

[1383] G. A. Hansen and C. H. Chang. Efficient visualization of a plasma spray simulation. Computers in Physics, 12(1):65–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168644.

Robbins:1998:UIP

[1384] K. A. Robbins and M. Gorman. Using image processing and computer animation to analyze spatial and temporal dynamics in a pattern-forming system. Computers in Physics, 12(1):73–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168643.

Cole:1998:GNF

[1385] James B. Cole. Generalized nonstandard finite differences and physical applications. Computers in Physics, 12(1):82–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168639.

Nishidate:1998:ACA

[1386] Kazume Nishidate, Mamoru Baba, and Richard J. Gaylord. ASECA: a cellular-automata simulation program for a silicon anisotropic supermicro-etching process in aqueous KOH. Computers in Physics, 12(1): 88–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168640.

Thompson:1998:PVI

[1387] William J. Thompson. Principal-value integrals by a simple and accurate finite-interval method. Computers in Physics, 12(1):94–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168641.

Omelyan:1998:NIM

[1388] Igor P. Omelyan. On the numerical integration of motion for rigid polyatomics: The modified quaternion approach. Computers in Physics, 12 (1):97–??, January 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168642.

Holmes:1998:EF

[1389] Lewis M. Holmes. An educational focus. Computers in Physics, 12(2): 107–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822622.

Foster:1998:CGD

[1390] Ian Foster and Carl Kesselman. Computational grids: On-demand computing in science and engineering. Computers in Physics, 12(2):109–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168654.

Lewin:1998:IRG

[1391] David I. Lewin. An interview with Roscoe Giles: From theory to practice. Computers in Physics, 12(2):110–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168626.

Donnelly:1998:EFW

[1392] Denis Donnelly. Educators focus on the Web in New Orleans. Computers in Physics, 12(2):112–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822623.

Daukantas:1998:UAM

[1393] Patricia Daukantas. Update: Asteroids make waves. Computers in Physics, 12(2):114–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168627.

Daukantas:1998:UAC

[1394] Patricia Daukantas. Update: Astronomers compute. New e-journal surfaces. Computers in Physics, 12(2):115–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168629.

Daukantas:1998:CJSb

[1395] Patricia Daukantas. From the CIP journal section. Computers in Physics, 12(2):116–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822624.

Titus:1998:WBT

[1396] Aaron P. Titus, L. W. Martin, Robert J. Beichner, Kevin Zollman, and Dean Zollman. Web-based testing in physics education: Methods and opportunities. Computers in Physics, 12(2):117–??, March 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168628.

Anonymous:1998:VBC

[1397] Anonymous. The virtual bookshelf: CIP reviews and readers’ opinions. Computers in Physics, 12(2):124–??, March 1998. CODEN CPHYE2.

ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168622.

Anonymous:1998:IGH

[1398] Anonymous. Internet goldmine: Hypatia and the collection of computer science bibliographies. Computers in Physics, 12(2):125–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168636.

Bonomi:1998:PSP

[1399] Ernesto Bonomi, Leesa Brieger, Carlo Nardone, Enrico Pieroni, Ralf Gruber, and Jacques Rappaz. Phase shift plus interpolation; a scheme for high-performance echo-reconstructive imaging. Computers in Physics, 12 (2):126–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168623.

Sung:1998:CEP

[1400] Ruwang Sung, Willard Fadner, Courtney Willis, Kendall Mallory, and Denis Donnelly. Computer-enhanced physics laboratories. Computers in Physics, 12(2):133–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168635.

Giordano:1998:PVS

[1401] Nicholas Giordano, Harvey Gould, and Jan Tobochnik. The physics of vibrating strings. Computers in Physics, 12(2):138–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168621.

Matey:1998:BRE

[1402] James R. Matey, Susan R. McKay, and Wolfgang Christian. Book review: Edward R. Tufte, Visual Explanations: Images and Quantities, Evidence and Narrative. Computers in Physics, 12(2):146–??, March 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168637.

Natarajan:1998:BRM

[1403] Lakshmi Natarajan. Book review: M. Ya Amusia and L. V. Chernysheva, Computation of Atomic Processes: a Handbook for the ATOM Programs. Computers in Physics, 12(2):148–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168638.

Stift:1998:SPA

[1404] Martin J. Stift and Paul F. Dubois. Scientific programming with Ada 95: Object-oriented, parallel, and safe. Computers in Physics, 12(2): 150–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168624.

Armstrong:1998:VEF

[1405] Daryl Armstrong, Ian Llanas, Frank Russo, Jeffrey R. Schmidt, and Steve Bryson (Editor). Visualization of electromagnetic fields using Awk. Computers in Physics, 12(2):159–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168631.

Shirer:1998:FSO

[1406] Donald L. Shirer. Fortran 90 subsets offer simplified environment for firsttime users. Computers in Physics, 12(2):166–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168632.

Anonymous:1998:NSL

[1407] Anonymous. Numerical software library. Computers in Physics, 12(2): 169–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168633.

Elisseev:1998:PTD

[1408] V. V. Elisseev. Parallelization of three-dimensional spectral laser–plasma interaction code using High Performance Fortran. Computers in Physics, 12(2):173–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168630.

Cole:1998:EHA

[1409] J. B. Cole. Erratum: “High accuracy solution of Maxwell’s equations using nonstandard finite differences” [Comput. Phys. 11, 287 (1997)]. Computers in Physics, 12(2):181–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168634. See [1290].

Pruksch:1998:PID

[1410] Matthias Pruksch and Frank Fleischmann. Positive iterative deconvolution with energy conservation. Computers in Physics, 12(2):182–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168651.

Taylor:1998:TFS

[1411] Edwin F. Taylor, Stamatis Vokos, John M. O’Meara, and Nora S. Thornber. Teaching Feynman’s sum-over-paths quantum theory. Computers in Physics, 12(2):190–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168652.

Cordoba-Torres:1998:OST

[1412] Pedro Córdoba-Torres, Fernando Enriquez, and Víctor Fairén. Optimal start of a Taylor integrator by control of local error. Computers in Physics, 12(2):200–??, March 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168653.

Holmes:1998:UIC

[1413] Lewis M. Holmes. Using the Internet to collaborate. Computers in Physics, 12(3):211–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822625.

Sargent:1998:CMS

[1414] Murray Sargent III. Comment: Making the switch from physics to software: a first-order phase transition. Computers in Physics, 12(3):213–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168669.

Ostriker:1998:CU

[1415] Jeremiah Ostriker. Computing the Universe. Computers in Physics, 12 (3):214–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168664.

Holmes:1998:CPC

[1416] Lewis M. Holmes. Condensed phases crack under computational assault: a report on the APS March Meeting. Computers in Physics, 12 (3):216–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822626.

Anonymous:1998:TYAa

[1417] Anonymous. Ten years ago: May/Jun 1988. Computers in Physics, 12 (3):217–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168659.

Anonymous:1998:SPM

[1418] Anonymous. Supercomputing partnerships move ahead; Pittsburgh, Cornell find new projects. Computers in Physics, 12(3):218–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168678.

Daukantas:1998:CJSc

[1419] Patricia Daukantas. From the CIP journal section. Computers in Physics, 12(3):219–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822627.

Casper:1998:REE

[1420] T. A. Casper, B. Howard, R. Hunt, W. H. Meyer, J. M. Moller, B. McHarg, P. Henline, K. Keith, K. Greene, S. Davis, D. Barnes, T. Gibney, P. Roney, and D. Greenwood. Remote experimental environment: Building a collaboratory for fusion scientific research. Computers in Physics, 12(3):220–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168665.

Christian:1998:DWB

[1421] Wolfgang Christian and Aaron Titus. Developing Web-based curricula using Java physlets. Computers in Physics, 12(3):227–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168666.

Becker:1998:CWOa

[1422] Mark Becker. CIP Website: Online version enhances article on computational-physics course. Computers in Physics, 12(3):233–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168662.

Becker:1998:P

[1423] Mark Becker. PhysLINK. Computers in Physics, 12(3):234–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168658.

Paavola:1998:CRB

[1424] Stephen Paavola and James R. Matey. Catching the right bus, Part VIII: Packet switching for very-high-performance data-acquisition and analysis systems. Computers in Physics, 12(3):235–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168668.

Landau:1998:WEU

[1425] Rubin H. Landau, Hans Kowallik, and Manuel J. Páez. Web-enhanced undergraduate course and book for computational physics. Computers in Physics, 12(3):240–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168672.

Gould:1998:BRM

[1426] Harvey Gould and Jan Tobochnik. Book review: M. Di Ventra and C. J. Fall, General solution scheme for second-order differential equations: application to quantum transport. Computers in Physics, 12(3):248–249, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168679.

Zachary:1998:BRA

[1427] Joseph L. Zachary. Book reviews: Atlas for computing mathematical functions, by William J. Thompson. Computers in Physics, 12(3):254–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168670.

Furnish:1998:CFN

[1428] Geoffrey Furnish. Container-free numerical algorithms in C++. Computers in Physics, 12(3):258–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168674.

Shirer:1998:MVT

[1429] Donald L. Shirer. Maple V turns over a new leaf. Computers in Physics, 12(3):267–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168673.

Anonymous:1998:IUN

[1430] Anonymous. Irix and Unicos networking technologies. Computers in Physics, 12(3):269–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168661.

Anonymous:1998:NAP

[1431] Anonymous. Numerical-analysis and plotting software. Computers in Physics, 12(3):269–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168660.

Anonymous:1998:FFS

[1432] Anonymous. Fluid-flow simulation software. Computers in Physics, 12 (3):270–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168663.

Anonymous:1998:DAV

[1433] Anonymous. Data-acquisition and virtual instrumentation software. Computers in Physics, 12(3):270–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168667.

Anonymous:1998:TPD

[1434] Anonymous. Technical-plotting and data-processing software. Computers in Physics, 12(3):270–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168680.

Anonymous:1998:SST

[1435] Anonymous. Simulation software teaches kinematics. Computers in Physics, 12(3):271–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168671.

Anonymous:1998:MRW

[1436] Anonymous. Mathematical reference work on CD-ROM. Computers in Physics, 12(3):271–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168675.

Anonymous:1998:DDA

[1437] Anonymous. DAQ digitizer/averager board. Computers in Physics, 12 (3):271–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168676.

Burgess:1998:OOV

[1438] Mark Burgess, Hårek Haugerud, and Are Strandlie. Object orientation and visualization of physics in two dimensions. Computers in Physics, 12 (3):274–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168655.

Balfagon:1998:TST

[1439] A. Balfagón and X. Jaén. TTC: Symbolic tensor calculus with indices. Computers in Physics, 12(3):286–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168656.

Simos:1998:CAP

[1440] T. E. Simos. Computer-algebra program for constructing exponentially fitted methods for solution of the Schrödinger equation. Computers in Physics, 12(3):290–??, May 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168657.

Holmes:1998:SCY

[1441] Lewis M. Holmes. Scientific computing on your desktop. Computers in Physics, 12(4):299–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822628.

Borcherds:1998:GHW

[1442] MP. H. Borcherds, J. Marro, and J. Nadrchalurray. Granada hosts a world venture in computational science. Computers in Physics, 12(4): 301–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168704.

Anonymous:1998:ISK

[1443] Anonymous. An interview with Sidney Karin. Computers in Physics, 12 (4):302–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168705.

Forbes:1998:DCP

[1444] Nancy A. Forbes and Richard J. Lipton. DNA computing: a possible efficiency boost for specialized problems. Computers in Physics, 12(4):

304–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822629.

Anonymous:1998:CPE

[1445] Anonymous. Collective phenomena, educational software top CCP 1198 program. Computers in Physics, 12(4):306–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168707.

Anonymous:1998:XSD

[1446] Anonymous. XML structures data for the web. Computers in Physics, 12 (4):310–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168708.

Anonymous:1998:TYAb

[1447] Anonymous. Ten years ago: Jul/Aug 1988. Computers in Physics, 12 (4):311–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168709.

Rapkin:1998:AA

[1448] Marwin Rapkin. Ada’s attractions. Computers in Physics, 12(4):312–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822630.

Kunz:1998:BAC

[1449] Michelle Kunz. Bravo on Ada 95 coverage. Computers in Physics, 12(4): 312–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822631.

Daukantas:1998:CJSd

[1450] Patricia Daukantas. From the CIP journal section. Computers in Physics, 12(4):313–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.4822632.

Cary:1998:RJC

[1451] John R. Cary, David A. Alexander, Kevin Zollman, and Dean Zollman. Reusable Java components for physics education. Computers in Physics, 12(4):314–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168714.

Becker:1998:CWOb

[1452] Mark Becker. CIP Website Online version enhances article on computational physics course. Computers in Physics, 12(4):319–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168688.

Becker:1998:SSD

[1453] Mark Becker. The structure, stability, and dynamics of self-gravitating systems. Computers in Physics, 12(4):320–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168712.

Wallin:1998:DLW

[1454] John F. Wallin and Denis Donnelly. Distance learning and the World Wide Web. Computers in Physics, 12(4):322–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168713.

Scott:1998:SLM

[1455] H. Larry Scott, Eric Jakobsson, Shankar Subramaniam, Harvey Gould, and Jan Tobochnik. Simulations of lipid membranes with atomic resolution. Computers in Physics, 12(4):328–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168715.

McKay:1998:BRY

[1456] Susan R. McKay and Wolfgang Christian. Book review: Yaneer Bar-Yam, Dynamics of Complex Systems (Studies in Nonlinearity). Computers in Physics, 12(4):335–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822633.

Newton:1998:BRS

[1457] Roger G. Newton and Wolfgang Christian. Book review: Sadhan K. Adhikari, Variational Principles and the Numerical Solution of Scattering Problems. Computers in Physics, 12(4):336–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822634.

Langer:1998:CFP

[1458] Steven H. Langer and Paul F. Dubois. A comparison of the floatingpoint performance of current computers. Computers in Physics, 12(4): 338–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168693.

Bryson:1998:TDV

[1459] Steve Bryson. Three-dimensional visualization for everyone. Computers in Physics, 12(4):346–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168694.

Shirer:1998:EAO

[1460] Donald L. Shirer. Excel add-ons improve graphing and statistical analyses. Computers in Physics, 12(4):352–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168703.

Anonymous:1998:FC

[1461] Anonymous. Fortran 95 compiler. Computers in Physics, 12(4):355–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.168695.

Anonymous:1998:ACC

[1462] Anonymous. ActiveX custom control for charting real-time data. Computers in Physics, 12(4):355–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168696.

Anonymous:1998:FSB

[1463] Anonymous. Free-surface and bubble-dynamics simulation code. Computers in Physics, 12(4):356–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168697.

Anonymous:1998:QAG

[1464] Anonymous. Quantitative analysis for gamma-ray spectroscopy. Computers in Physics, 12(4):356–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168698.

Anonymous:1998:SDA

[1465] Anonymous. Scientific data analysis on Windows. Computers in Physics, 12(4):356–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168699.

Anonymous:1998:IIA

[1466] Anonymous. IDL 5.1 includes ActiveX support. Computers in Physics, 12(4):356–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168700.

Anonymous:1998:IBV

[1467] Anonymous. Input board for VMEbus computers. Computers in Physics, 12(4):357–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168701.

Anonymous:1998:VTM

[1468] Anonymous. Visualization tool for mathematics. Computers in Physics, 12(4):357–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168702.

Anonymous:1998:SSI

[1469] Anonymous. Single-system-image 128-processor computers. Computers in Physics, 12(4):357–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168725.

Windt:1998:ISM

[1470] David L. Windt. IMD — software for modeling the optical properties of multilayer films. Computers in Physics, 12(4):360–??, July 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168689.

Mitchell:1998:RTP

[1471] Chris Mitchell and Walter Gekelman. Real-time physics datavisualization system using Performer. Computers in Physics, 12(4): 371–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168690.

Karlen:1998:UPC

[1472] Dean Karlen. Using projections and correlations to approximate probability distributions. Computers in Physics, 12(4):380–??, July 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168691.

Ziff:1998:FTS

[1473] Robert M. Ziff. Four-tap shift-register-sequence random-number generators. Computers in Physics, 12(4):385–392, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168692.

Thijsse:1998:PAL

[1474] Barend J. Thijsse, Mark A. Hollanders, and Jan Hendrikse. A practical algorithm for least-squares spline approximation of data containing noise. Computers in Physics, 12(4):393–??, July 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168716.

Holmes:1998:NMW

[1475] Lewis M. Holmes. New magazine will replace computers in physics. Computers in Physics, 12(5):403–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822635.

Duke:1998:MYM

[1476] Dennis Duke. Moore for your money at SC98. Computers in Physics, 12(5):405–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168733.

Lewin:1998:NEC

[1477] David I. Lewin. From nanocomputers to the Earth’s core. Computers in Physics, 12(5):406–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822636.

Daukantas:1998:SCL

[1478] Patricia Daukantas. SC98 Conference looks to the future back in Disneytown. Computers in Physics, 12(5):408–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822637.

Anonymous:1998:TYAc

[1479] Anonymous. Ten years ago: Sep/Oct 1988. Computers in Physics, 12(5):410–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168729.

Neill:1998:RSC

[1480] Greg Neill. Reader studies chaotic-pendulum problem. Computers in Physics, 12(5):411–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822638.

Daukantas:1998:CJSe

[1481] Patricia Daukantas. From the CIP journal section. Computers in Physics, 12(5):412–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822639.

Cumaranatunge:1998:CIM

[1482] Chandima Cumaranatunge. Creating interactive multimedia simulations without programming. Computers in Physics, 12(5):413–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168732.

Anonymous:1998:COI

[1483] Anonymous. Complete online index of CIP. Computers in Physics, 12(5):422–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168730.

Anonymous:1998:PPU

[1484] Anonymous. Parallel processing using Linux. Computers in Physics, 12(5):423–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168731.

Koplik:1998:PFL

[1485] Joel Koplik and Jayanth R. Banavar. Physics of fluids at low Reynolds numbers — a molecular approach. Computers in Physics, 12(5):424–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168734.

Young:1998:EFM

[1486] Richard A. Young and Denis Donnelly. Electric-field mapping revisited. Computers in Physics, 12(5):432–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168735.

Billeter:1998:SLC

[1487] Jeffrey Billeter and Robert Pelcovits. Simulations of liquid crystals. Computers in Physics, 12(5):440–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168726.

Rollins:1998:BRR

[1488] Robert W. Rollins, Susan R. Mackay, and Wolfgang Christian. Book review: Robert H. Silsbee and Jörg Dräger, Simulations for Solid State Physics. Computers in Physics, 12(5):449–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168736.

Snowman:1998:BRJ

[1489] Daniel P. Snowman, Susan R. Mackay, and Wolfgang Christian. Book review: Jerry Glynn and Theodore Gray, The Beginner’s Guide to Mathematica Version 3. Computers in Physics, 12(5):451–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168737.

Reynders:1998:PF

[1490] John V. W. Reynders III, Julian Cummings, and Paul F. Dubois (Editor). The POOMA framework. Computers in Physics, 12(5):453–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168723.

Consularo:1998:MIM

[1491] Luis Augusto Consularo, Luciano da Fontoura Costa, and Donald L. Shirer. MATCOM integrates MATLAB resources into standalone applications. Computers in Physics, 12(5):460–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168722.

Anonymous:1998:LLR

[1492] Anonymous. Listings of the latest releases. Computers in Physics, 12(5):463–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.4822640.

Vigo-Aguiar:1998:HOV

[1493] Jesús Vigo-Aguiar and José M. Ferrándiz. Higher-order variable-step algorithms adapted to the accurate numerical integration of perturbed oscillators. Computers in Physics, 12(5):467–??, September 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168717.

Kidan:1998:NSA

[1494] T. Kidan, Joan Adler, and A. Ron. Numerical simulations of an atomic beam focused by an optical potential. Computers in Physics, 12(5):471– ??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168718.

Fowler:1998:DBS

[1495] Howland A. Fowler, Judith E. Devaney, John G. Hagedorn, and Francis E. Sullivan. Dielectric breakdown in a simplified parallel model. Computers in Physics, 12(5):478–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168738.

Biswas:1998:GTS

[1496] Tarun Biswas. Group theory (symmetries) in a computer game. Computers in Physics, 12(5):488–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168719.

Williams:1998:DRT

[1497] S. A. Williams, M. L. Sawley, and D. Cobut. Distributed run-time visualization and solution steering of parallel flow simulations. Computers in Physics, 12(5):493–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168720.

Trefry:1998:LFN

[1498] M. G. Trefry and L. R. Townley. Laplace flow near an ellipsoidal conductor. Computers in Physics, 12(5):503–??, September 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168721.

Holmes:1998:SF

[1499] Lewis M. Holmes. Segue to the future. Computers in Physics, 12(6):515– ??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822641.

Holmes:1998:NME

[1500] Lewis M. Holmes and George Cybenko. New magazine expands AIP’s Computers in Physics and IEEE Computational Science & Engineering. Computers in Physics, 12(6):516–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.4822642.

Anonymous:1998:AE

[1501] Anonymous. The accidental entrepreneur. Computers in Physics, 12(6): 518–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168753.

Abelson:1998:ACT

[1502] Harold Abelson and Nancy Forbes. Amorphous-computing techniques may lead to intelligent materials. Computers in Physics, 12(6):520– ??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
4822643.

Anonymous:1998:CSA

[1503] Anonymous. Conference spotlights advanced fluid-simulation techniques. Computers in Physics, 12(6):522–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168748.

Anonymous:1998:TCS

[1504] Anonymous. Talking computational shop in the penumbra of the Alhambra. Computers in Physics, 12(6):525–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168749.

Anonymous:1998:TYAd

[1505] Anonymous. Ten year ago: Nov/Dec 1988. Computers in Physics, 12(6):526–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168752.

Sherwood:1998:LE

[1506] Bruce A. Sherwood. Letter to the editor. Computers in Physics, 12(6): 528–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/10.
1063/1.168750.

Daukantas:1998:CJSf

[1507] Patricia Daukantas. From the CIP journal section. Computers in Physics, 12(6):529–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.4822644.

Esteban:1998:UWT

[1508] E. P. Esteban, F. J. Agosto, and Kevin Zollman. Using Web technology in undergraduate research. Computers in Physics, 12(6):530–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL https://aip.scitation.org/doi/10.1063/1.168751.

Anonymous:1998:CWS

[1509] Anonymous. CIP website supports CiSE subscribers. Computers in Physics, 12(6):535–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168755.

Becker:1998:CRP

[1510] Mark Becker. The Center for Research on Parallel Computation. Computers in Physics, 12(6):536–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168761.

Abraham:1998:SLS

[1511] Farid F. Abraham, Jeremy Q. Broughton, Noam Bernstein, and Efthimios Kaxiras. Spanning the length scales in dynamic simulation. Computers in Physics, 12(6):538–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168756.

Matey:1998:NDO

[1512] James R. Matey. New directions and old: a farewell to Computers in Physics. Computers in Physics, 12(6):547–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168757.

Goedecker:1998:SMP

[1513] Stefan Goedecker and Oleg Ivanov. Solution of multiscale partial differential equations using wavelets. Computers in Physics, 12(6):548– ??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-

4208 (electronic). URL https://aip.scitation.org/doi/10.1063/1.
168739.

Karabach:1998:IBA

[1514] Michael Karabach, Kun Hu, and Gerhard Müller. Introduction to the Bethe Ansatz. II. Computers in Physics, 12(6):565–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168740.

Strogatz:1998:BRT

[1515] Steven H. Strogatz. Book review: Thomas P. Weissert, The genesis of simulation in dynamics: Pursuing the Fermi–Pasta–Ulam problem. Computers in Physics, 12(6):574–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168754.

Powell:1998:BRD

[1516] Adam C. Powell IV. Book review: Daniel H. Rothman and Stéphane Zaleski, Lattice-gas cellular automata: Simple models of complex hydrodynamics. Computers in Physics, 12(6):576–??, November 1998. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL

https://aip.scitation.org/doi/10.1063/1.168762.

Lichten:1998:SGC

[1517] William Lichten and Donald L. Shirer. Scientific and graphing calculators. Computers in Physics, 12(6):588–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168760.

Anonymous:1998:NPb

[1518] Anonymous. New products. Computers in Physics, 12(6):591–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (elec-

tronic). URL https://aip.scitation.org/doi/10.1063/1.4822645.

Alda:1998:CFD

[1519] Witold Alda, Witold Dzwinel, Jacek Kitowski, Jacek Mościński, Marek Pogoda, and David A. Yuen. Complex fluid-dynamical phenomena modeled by large-scale molecular-dynamics simulations. Computers in Physics, 12(6):595–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168741.

Martin:1998:UJS

[1520] Jared Martin, Lyle N. Long, and Kevin M. Morooney. Using Java for scientific programming and electromagnetics. Computers in Physics, 12(6):601–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168742.

Gutierrez:1998:MPA

[1521] José Manuel Gutiérrez and Andrés Iglesias. Mathematica package for analysis and control of chaos in nonlinear systems. Computers in Physics, 12(6):608–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.org/

doi/10.1063/1.168743.

Weller:1998:TAC

[1522] H. G. Weller, G. Tabor, H. Jasak, and C. Fureby. A tensorial approach to computational continuum mechanics using object-oriented techniques. Computers in Physics, 12(6):620–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.168744.

Wei:1998:NFS

[1523] Liqiang Wei. New formula for 9–j symbols and their direct calculation. Computers in Physics, 12(6):632–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168745.

Simos:1998:ESO

[1524] T. E. Simos. Explicit sixth-order Bessel and Neumann fitted method for the numerical solution of the Schrödinger equation. Computers in Physics, 12(6):635–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.scitation.

org/doi/10.1063/1.168746.

Kulkarni:1998:NAS

[1525] Jaydeep A. Kulkarni and Antony N. Beris. New approach for simulating chain conformations in dense polymers using fully populated lattice models. Computers in Physics, 12(6):641–??, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https:

//aip.scitation.org/doi/10.1063/1.168747.

Wallace:1996:FPG

[1526] C. S. Wallace. Fast pseudorandom generators for normal and exponential variates. ACM Transactions on Mathematical Software, 22(1):119–127, March 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295

(electronic). URL http://www.acm.org/pubs/citations/journals/
toms/1996-22-1/p119-wallace/. See comments [1528].

Fernandez:1999:AER

[1527] Julio F. Fernández. Algorithm for exponential random numbers. Computer Physics Communications, 121–122:78–82, September/October 1999. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/article/
pii/S0010465599002854.

Brent:2008:SCC

[1528] Richard P. Brent. Some comments on C. S. Wallace’s random number generators. The Computer Journal, 51(5):579–584, February 2008. CO-DEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). See